A Complete Bibliography of Publications in Limnology and Oceanography Letters

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

23 May 2024 Version 1.11

Title word cross-reference

14 [SYM^{+}23]. _{2} [AHL^{+}21, AS24, BHBT23, CSEP^{+}18, FRVD18, HBN^{+}22, LR18, MCS18, RRSL^{+}19, VSB^{+}20, ZPRT23]. _{275−−295} [SYM^{+}23]. _{4} [AS24, CSEP^{+}18, DH21, ZPRT23]. : [LSH^{+}22, OWT^{+}23]. Δ [SYM^{+}23].

10 [Ano22]. 10^{o} [LSH^{+}22]. 14C [MBB^{+}21].

33-yr [CP18].

5 [HFKBW21].

7 [Ano22].

A.M [Ano22]. abiotic [HC20, LHMF20]. absorption [FKG^{+}22]. absorption-based [FKG^{+}22]. abundance [HSCA^{+}23, RBI^{+}23]. abundances [VMF21]. abundant [VCL^{+}20]. abyssal [LK24, MHSD22]. access [HDPEF23]. accessibility [FJS24]. Acclimation [MYS^{+}23, LKB^{+}22]. accounting [MSM^{+}17]. Accumulation [CAO22, PLB^{+}17, STS^{+}17]. accuracy [WLB^{+}20]. acid [HLAW23]. acidification [GES19, MF23, NDRL18, UMDWC^{+}21, ZZD^{+}23]. acids [PKKL24]. Acoustic [LK24]. across [ARG^{+}19, ASV^{+}21, CDC^{+}21, CSW^{+}20, CTC^{+}20, HPS^{+}18, HSCA^{+}23, HSH18, KOLD^{+}21, LSH^{+}22, OAL^{+}19, RBI^{+}23, SBAPS20, SGSdGB17, TFHK18, TSB^{+}22]. act [FBDH18]. actively [SFT^{+}17]. activity [DLT^{+}19]. acute [KHB^{+}23]. adapt [BIF20]. adaptability [BGG^{+}23]. adaptations [DDF^{+}23]. Addition [CRP^{+}18]. address [HDPEF23]. Addressing [MSM^{+}17, CHFF22]. adenosine [HFKBW21]. adenosine-5’-triphosphate [HFKBW21]. Adirondack [WCP23]. adsorb [DBVB20]. advances [FBV^{+}20]. aeolian [RMV^{+}18]. aerobic [CAO22]. affect [SMP19]. affected [CBF17, SMSS19]. afforestation [Kri17]. after [AOPP21]. age [DDH18, Sme21, KHS^{+}21]. aged [BB17]. aggregate [PZC^{+}17]. aggregates [LWP22]. aggregation [GK22]. air [BHBT23]. Al [ZKG^{+}22]. Al-Wajh [ZKG^{+}22]. Alaskan [MCS18]. algae [FKS20, GRLS^{+}23, OJ22, QRF^{+}18]. algal [BCP18, HBL^{+}24, Lit23, RTK^{+}18, XSQ^{+}23]. alkaline [DLT^{+}19]. Alkalinity [PEC^{+}23, CDC^{+}21, SFT^{+}21]. ALOHA [HFKBW21]. along [Ano22, CK18, MFS22, OTP23]. also [RHN^{+}23]. alter [JWSD22]. altered [JEN^{+}20]. alters [MF23, SMC20, UMDWC^{+}21, dCCS21]. Always [Hop22]. Amazon [CSEP^{+}18, STS^{+}17, WSN^{+}18]. Amazonian [CDC^{+}21, STS^{+}17]. American [BSH^{+}20]. amidst [MD23]. amino [PKKL24]. among [CSW^{+}20, HLAW23, RTK^{+}18]. among-lake [RTK^{+}18]. amplified [RWRH16, WSN^{+}18]. amplifies [TSB^{+}22]. Anaerobic [LHL23]. analoga [HGS20]. analysis [BCP18, HAWH19, MPB^{+}23]. annual [KVS^{+}24, PEC^{+}23]. anoxia [HMLC21, RLH^{+}24]. anoxic [BRK^{+}22, ZPRT23]. answers [GD22]. Antarctic [DDF^{+}23, GRLS^{+}23, KHS^{+}21]. Antarctica [CDG^{+}22]. antecedent [AGM^{+}23]. Anthropogenic [RH18, MCGK22]. antioxidant [MB22]. application [THV23]. approach [FKG^{+}22, SdG18]. approaches [MBR23]. Aquatic [BWK^{+}23, ARG^{+}19, BB17, CK18, CAO22, GKR^{+}23, GBH20, HPS^{+}18, JB22, KBB^{+}21, LGGF^{+}23, MPB^{+}23, MIK18, MKKS17, RCA^{+}21, Sor16, TFHK18, THV23, TFF23, VSB^{+}20, WJS18, WWFP20, YWH^{+}17]. archive [AE23, MPB^{+}23]. Arctic [BIF20, Hop22, KVS^{+}24, MCS18, MAE^{+}18, OAH^{+}23, OAL^{+}19, SGB^{+}17, dlVMT^{+}21]. Area [CMR^{+}21, CS22, VH23]. areas [MSSC19]. arid [RMV^{+}18]. arsenic [HMIB22]. arsenic-containing [HMIB22]. article [Clo23]. articles [FJS24]. artificial [HRR24]. ascidian [KCPGE21]. asked [GD22]. ASLO [GD22]. assemblage [WEW23]. assemblages [LCL22]. associated [HBB^{+}16, UMDWC^{+}21]. Atlantic [Bri22, CJ20]. atmospheric [ZTZL21]. attenuation [HFKBW21]. Augmentation [Sme21]. austral [YAML^{+}23]. Australia [CMR^{+}21, CMR^{+}24]. autochthonous [HRR24]. autotrophic [SAE^{+}22]. autotrophy [MAE^{+}18]. availability [FKS20, KCBG^{+}23, KLA^{+}17, MLPM21]. Avenues [GKR^{+}23].

B [LSH^{+}22]. Ba [GYB^{+}23]. Ba/Ca [GYB^{+}23]. bacteria [HBB^{+}16, VCL^{+}20, WGZ^{+}19]. bacterial [BWAJ19, LFS19, RBvdK^{+}18]. bacterivorous [LDWE24]. Baikal [MOW^{+}22]. balance [CHB^{+}22, MAE^{+}18]. bank [WEW23]. Barotropic [CDG^{+}22]. Barrier [CMR^{+}21]. basal [RCA^{+}21]. based [ARK^{+}23, ASV^{+}21, Cae21, FKG^{+}22, Lit23, MPB^{+}23]. basin [SSV^{+}22, CSEP^{+}18]. Bay [CDP^{+}18, VRF^{+}19, MCGK22, WLC20, XSQ^{+}23]. bays [BFC^{+}23]. be [CBBE23, DGS20, RWRH16]. beam [HFKBW21]. becoming [LHMF20]. bed [CEM^{+}22, CCT^{+}20]. behavior [CMS^{+}21, LFH^{+}23, NLC20, RJS^{+}19]. behind [MSSC19]. bend [SH17]. benthic [ARG^{+}19, BMH^{+}23, CCT^{+}20, FOE20, FKS20, HBL^{+}24, HLAW23, MOW^{+}22, PKKL24, RBI^{+}23]. benthos [WBK17]. Bering [CLTK^{+}22]. best [TFF23]. between [CS22, CK18, HDC^{+}17, LLA^{+}22, RWI^{+}21, SKDR21, THV23, YWH^{+}17]. Beyond [TFHK18, WWFP20]. Bifurcate [CQSZ22]. binding [HBB^{+}16]. bioavailable [PKKL24]. biodiversity [AGM^{+}23, BMH^{+}23]. Biogeochemical [MIK18, RWI^{+}21, LHL23, SW23, VRB^{+}23]. bioinvasion [GHHKY^{+}17]. biological [CBF17]. biomass [CLTK^{+}22, FKG^{+}22, KSL^{+}18, RTK^{+}18, SF23]. biomineral [MHSD22]. Biophysical [XSQ^{+}23]. bivalve [BGHC20, NDRL18]. black [BRP^{+}24, CSEP^{+}18, WJS18, YCS^{+}21]. bleaching [CDP^{+}18, MD23]. bloom [Bri22, ZJL^{+}22]. Blooms [RHN^{+}23, BCP18, HBL^{+}24, HBB^{+}16, Lit23, XSQ^{+}23]. blue [CSH^{+}22, MSM^{+}17, WLC20, SFT^{+}21]. bluefin [HCA^{+}23]. bodies [LSER23]. Body [KOLD^{+}21, RBI^{+}23, SH17]. body-size [RBI^{+}23]. boreal [CRP^{+}18, OAH^{+}23, SGSdGB17, dWCJB^{+}18]. both [BMH^{+}23, OWT^{+}23]. Bottom [FOE20]. bottoms [BPL^{+}19]. branconii [DDF^{+}23]. break [SH17]. breathing [VMM^{+}22]. browning [BMH^{+}23, Kri17, OAH^{+}23, PAMB23, SKDR21]. budget [SGB^{+}17]. budgets [STS^{+}17]. burial [BPL^{+}19, CDMH18, HRR24]. burn [BRP^{+}24].

C [HAWH19, LSH^{+}22, SYM^{+}23, SdG18]. C-DOC [SYM^{+}23]. Ca [GYB^{+}23]. calcification [CDP^{+}18]. calcium [MSM^{+}17]. California [BRP^{+}24]. call [TFF23]. Callinectes [WLC20]. Can [STW17, CBBE23, FBDH18, FRVD18, KBB^{+}21, KHS^{+}21, NDRL18, RWRH16, SBJ^{+}23]. Canadian [ASV^{+}21]. Candidate [VCL^{+}20]. Capacity [BIF20, DBG21]. captures [CMS^{+}21]. Carbon [CDC^{+}21, RDFD18, STS^{+}17, WCA^{+}18, AOPP21, BWK^{+}23, BRP^{+}24, BPL^{+}19, BB17, CBF17, CDMH18, CSH^{+}22, DBG21, DBdG^{+}18, DRS18, DDH18, GES19, GRK^{+}17, HFKBW21, HRR24, HSH18, JZKS18, KVS^{+}24, LFS19, MSM^{+}17, MJC18, MYS^{+}23, MIK18, OAL^{+}19, PZC^{+}17, RH18, SLC18, SMT^{+}17, Sme21, SKDR21, SF23, TFHK18, TCP18, VLD^{+}20, VBK^{+}19, VCL^{+}20, WJS18, WCP23, WHW^{+}20, WML^{+}19, dWCJB^{+}18, dCCS21, SFT^{+}21]. carbon-concentrating [VBK^{+}19]. carbonate [MSM^{+}17, YHD23]. career [GD22, HDPEF23]. Caribbean [DCPN20]. Carpentaria [CMR^{+}24]. cascades [McI22]. Cascading [MBB^{+}23]. case [CHFF22, CCT^{+}20, HDC^{+}17, LLA^{+}22]. catalysts [VCL^{+}20]. Catch [KSL^{+}18]. catchment [BWK^{+}23]. catchments [DDH18, OJ22]. caused [McI22]. Causes [SSV^{+}22, BMH^{+}23]. CDOM [SYM^{+}23]. cellular [MB22]. CENA392 [DDF^{+}23]. Centennial [Kri17]. Centennial-long [Kri17]. central [BRP^{+}24, BSSB23]. centric [SNAA16]. century [DAS^{+}23]. CH [AS24, CSEP^{+}18, DH21, ZPRT23]. Chaetoceros [PMP^{+}20]. chain [AAG^{+}20]. challenges [SBJ^{+}23, SDC21]. change [BMH^{+}23, CC21, DBD18, DSA^{+}23, LLA^{+}22, VMM^{+}22, WDT^{+}22, dlVMT^{+}21]. change-related [DSA^{+}23]. Changes [MHSD22, CBF17, HBL^{+}24]. changing [Lit23, MGG^{+}23, SNAA16, YHD23]. characteristics [CSM^{+}21]. Characterization [LEvdM^{+}24, ASV^{+}21]. characterizing [MIK18]. chemical [QPC^{+}16, dCCS21]. chemistry [YHD23]. chemoheterotrophic [WGZ^{+}19]. Chesapeake [MCGK22, XSQ^{+}23]. chimneys [dWCJB^{+}18]. chloride [DHM17, DRKM23, GAFS23, MAR23]. Chlorophyll [OJ22, CP18, LEvdM^{+}24, PE22]. choanoflagellates [MKKS17]. Christi [WLC20]. ciliates [LLA^{+}22]. circulation [CJ20, MCS18]. circumpolar [WRC^{+}18]. citizen [LBB^{+}24]. city [KCBG^{+}23]. Clarifying [TFF23]. clarity [LJ24, RWRH16, TFF23, VH23, WLB^{+}20]. classes [RBI^{+}23]. Classifying [BSSB23, RPM^{+}23]. clear [CSEP^{+}18, KVS^{+}24]. clear-water [KVS^{+}24]. clearance [HC20]. Climate [RWRH16, SMSS19, TSB^{+}22, BPL^{+}19, CC21, DSA^{+}23, HDC^{+}17, HLH23, Lit23, LLA^{+}22, SFW^{+}23, SMS20, VMM^{+}22, WDT^{+}22, YHD23]. Climate-induced [RWRH16]. climatic [OAL^{+}19]. cloud [MPB^{+}23]. cloud-based [MPB^{+}23]. CO [VSB^{+}20, BLZ^{+}22, AHL^{+}21, AS24, BHBT23, CSEP^{+}18, LR18, MCS18, RRSL^{+}19, ZPRT23]. co-limitation [BLZ^{+}22]. coast [LDWE24]. Coastal [BGG^{+}23, BH23, MF23, AS24, BRP^{+}24, MHD21, MCAD23, OTP23, OAH^{+}23, PZB^{+}23, WBK17, WML^{+}19]. coastline [CHLD18]. cobalamin [BMB24]. coccolithophorid [GES19]. cold [BPL^{+}19, RHN^{+}23, SW20]. cold-water [SW20]. Collaboration [LBB^{+}24]. colocated [MOW^{+}22]. Colombia [SSV^{+}22]. colonization [DCPN20]. color [CBBE23, CP18]. column [JBG^{+}21, PZC^{+}17]. commercially [BSH^{+}20]. common [BIF20]. communities [GAFS23, MAR23, MWL^{+}23, MBB^{+}23, OTP23, UMDWC^{+}21, YSN^{+}23, ZZD^{+}23]. community [AFS^{+}23, AGLD23, BWAJ19, CBH^{+}22, CRP^{+}18, FBV^{+}20, KHB^{+}23, MOW^{+}22, SGSdGB17, WML^{+}19, LBB^{+}24]. community-level [AFS^{+}23]. comparisons [YJJ^{+}23]. compensation [LCL22]. complexity [HPS^{+}18]. component [LOS^{+}20, SFT^{+}21, STS^{+}17]. composition [BWAJ19, LRCG20, NDRL18]. compositionally [QRF^{+}18]. concentrating [VBK^{+}19]. concentration [MAR23]. concentrations [BGHC20, HGS20, HMLC21, ICH^{+}23]. conceptual [HDC^{+}17, YJJ^{+}23]. concise [Hot20]. conditions [OAL^{+}19]. confirms [SNAA16]. connected [HJS23]. connections [LGGF^{+}23]. connectivity [KWRC21, RWI^{+}21]. consensus [BHBT23]. Consequences [KSD^{+}19, LHL23, SSV^{+}22, WLL22, YBZ^{+}23]. conservation [MSSC19]. considerations [Hei23]. consistent [HSCA^{+}23]. Constraining [HFKBW21]. constrains [BSYD18, JZKS18]. constraint [Cae21]. constraints [MGG^{+}23, MCGK22]. consumer [RCA^{+}21]. consumption [DH21, FSA21, PCS^{+}20]. containing [HMIB22]. contamination [ARK^{+}23, SW20, WLC20]. content [HLAW23, MHSD22, WGZ^{+}19]. contents [SAE^{+}22]. conterminous [CSM^{+}21, KWRC21, RPM^{+}23, WRT^{+}22]. Continental [JBG^{+}21]. continua [OTP23]. continuum [Ano22, CDC^{+}21, CK18, MFS22]. Contrasting [PAMB23, SAE^{+}22]. contributes [PEC^{+}23]. contribution [KCPGE21, PZB^{+}23, SGB^{+}17]. Contributions [GRK^{+}17, MB22]. control [CDMH18, KML^{+}23, RWI^{+}21, XSQ^{+}23]. controls [LR18, TAC^{+}21, TFHK18]. coordinated [HSCA^{+}23]. Copepod [HAWH19, RGK22]. copepods [AHM21]. Copiotrophic [HBB^{+}16]. Coral [MD23, CTC^{+}20, JWSD22]. corals [QRF^{+}18]. Corpus [WLC20]. corresponding [TFF23]. Corrigendum [Ano22]. cost [Inm22]. coupling [KOLD^{+}21]. covariance [ARG^{+}19, JB22]. cover [CLB^{+}21, ZTZL21]. covered [CDG^{+}22, DBdG^{+}18, MCS18]. crab [HGS20]. crabs [NLC20, WLC20]. creek [CDC^{+}21, RWI^{+}21]. creeks [ZXC^{+}24]. Cross [ARK^{+}23]. Cross-contamination [ARK^{+}23]. Crossing [SFT^{+}17]. cryptic [MGJ^{+}18]. cultural [KSD^{+}19]. Current [CRH^{+}23, AE23, CEM^{+}22, DRS18, BSH^{+}20, DBG21]. current-driven [CEM^{+}22]. cyanobacteria [CLTK^{+}22, HHSL20, LKB^{+}22, OWT^{+}23, CAH^{+}20]. cyanobacterium [DDF^{+}23]. cyanotoxin [SW23]. cycle [FOE20, PZC^{+}17, SW23, TCP18]. cycles [SF23]. cycling [MSM^{+}17, SLC18, SdG18, dWCJB^{+}18]. cyclone [CHLD18]. cylindrus [JB20].

Daily [WRT^{+}22, SFW^{+}23, SF23]. Daphnia [WEW23]. darkening [OAH^{+}23]. data [CMR^{+}21, CMR^{+}24, Clo22, DAS^{+}23, VRB^{+}23, WLB^{+}20, CSM^{+}21]. database [MBB^{+}21, RPM^{+}23]. dataset [MOW^{+}22]. day [DAS^{+}23, SJL23]. debris [BFS20]. decadal [DBRB20, SFW^{+}23]. decadal-scale [DBRB20]. decade [GBH20, WCA^{+}18]. Deciphering [PE22]. decomposition [FT22]. Decoupled [GRLS^{+}23]. Deep [RRKLU^{+}20, CMS^{+}21, LEvdM^{+}24, PEC^{+}23, RDFD18, WRT^{+}22]. deep-sea [CMS^{+}21]. degradation [CK18, SSV^{+}22]. deliver [RBI^{+}23]. denitrification [CHFF22, FOE20, MH19]. Density [PMP^{+}20]. Density-dependent [PMP^{+}20]. dependent [PMP^{+}20]. depletion [AHL^{+}21, LHL23, YBZ^{+}23]. Depth [WH23, CS22, CTC^{+}20, MHSD22]. describe [Bri22]. desert [DAS^{+}23, KSL^{+}18]. despite [AHL^{+}21, FSA21]. determine [NDRL18, VMM^{+}22]. determined [MCGK22]. determines [BRP^{+}24, BOM^{+}23, MYS^{+}23, SMS20]. determining [QPC^{+}16]. diatom [PMP^{+}20, SNAA16]. diatoms [CCT^{+}20, GK22, GRLS^{+}23, KHS^{+}21, RGK22]. Diazotrophy [CRH^{+}23]. Diel [CTC^{+}20]. diet [LRCG20]. differ [HLH23]. differences [HDC^{+}17]. different [ASV^{+}21, KOLD^{+}21]. Differential [MWL^{+}23, CBCD19]. differentially [SMP19]. differing [DBRB20]. differs [LLA^{+}22]. dimensional [AGM^{+}23]. dimethylsulfide [GDP^{+}23]. dioxide [DBdG^{+}18, KVS^{+}24, VLD^{+}20, WCA^{+}18]. Direct [PZB^{+}23, CHLD18, DTSC17]. directions [BSH^{+}20, GBH20]. discharge [DSA^{+}23, GYB^{+}23, LL19, SBJ^{+}23, ZXC^{+}24]. Discovery [VRF^{+}19]. dismutase [CRlLC^{+}22]. dispersal [MKKS17]. dispersants [QPC^{+}16]. Dissolved [FKS20, WJS18, ASV^{+}21, BOM^{+}23, CK18, CAO22, FBDH18, HPS^{+}18, PKKL24, RH18, SKDR21, WCP23, VS18]. distinct [QRF^{+}18]. distribution [RBI^{+}23]. distributions [Bri22, SH17]. disturbance [McI22]. disturbances [AGM^{+}23]. Dityrosine [PCS^{+}20]. Diurnal [DLT^{+}19, BGG^{+}23, HBN^{+}22]. Divergent [ASW^{+}22]. Diverse [HMIB22, SJL23, CTC^{+}20, MYS^{+}23]. diversification [dCCS21]. diversity [CSW^{+}20, HSCA^{+}23, OTP23, UMDWC^{+}21]. DNA [CMS^{+}21]. Do [CCT^{+}20, DBVB20, MSSC19]. DOC [CPW22, SYM^{+}23]. Documenting [MCAD23]. Does [WEW23, LWP22]. DOM [STD^{+}18, VS18]. domains [RRSL^{+}19]. dominance [WRC^{+}18]. dominate [CCT^{+}20]. dominates [KVS^{+}24]. downstream [KSL^{+}18, LL19, TON^{+}20]. Drainage [BMH^{+}23]. Drainage-induced [BMH^{+}23]. drains [MHD21]. drier [OAL^{+}19]. drive [GRLS^{+}23, MAE^{+}18, RRSL^{+}19]. driven [BRK^{+}22, CP18, CEM^{+}22, HLAW23]. driver [HBL^{+}24, NLC20]. drivers [BKM^{+}23, YBZ^{+}23]. drives [HSCA^{+}23, LFS19, MCS18, MHD21, PLB^{+}17]. driving [SMS20]. dual [VMM^{+}22]. dual-breathing [VMM^{+}22]. due [WRC^{+}18]. duration [LHL23, PLB^{+}17]. during [DBdG^{+}18, HBB^{+}16, HMLC21, MHSD22, YAML^{+}23, ZTZL21, ZMB^{+}21]. Dynamic [STD^{+}18, BH23]. dynamical [DDH18]. dynamics [CDMH18, CAH^{+}20, CRP^{+}18, DBdG^{+}18, FBV^{+}20, MFHMD20, RDFD18, RCA^{+}21, SBAPS20, WEW23, WWFP20, ZKG^{+}22].

earliest [VBK^{+}19]. Early [GD22, CBCD19, HDPEF23, ORA^{+}22]. Earthquake [YSN^{+}23]. East [YSN^{+}23, CMR^{+}24, CDG^{+}22, PKKL24]. East/Japan [PKKL24]. eastern [RBvdK^{+}18]. eat [ORA^{+}22]. ebullition [ZTZL21]. ecological [CMS^{+}21, KHS^{+}21, dlVMT^{+}21]. ecology [Hei23, HDSH^{+}22]. ecosystem [ARG^{+}19, BCI17, DTSC17, HMLC21, KSD^{+}19, MH19, OTP23, SFT^{+}21, VSB^{+}20, VCL^{+}20, VRB^{+}23]. ecosystems [CHB^{+}22, GKR^{+}23, MIK18, MCAD23, SAE^{+}22, SBAPS20, TCP18, THV23, TFF23, WJS18, WWFP20]. ecozones [ASV^{+}21]. ectotherms [VMM^{+}22]. eddy [ARG^{+}19, JB22]. editors [GD22]. eDNA [CBH^{+}22]. effect [Kri17, SW20, YHD23, ZZD^{+}23]. effective [MBR23]. Effects [AGLD23, DBG21, GAFS23, HHSL20, HGS20, KCBG^{+}23, AAG^{+}20, BSH^{+}20, BPL^{+}19, CBCD19, CRP^{+}18, CHB^{+}22, FKS20, FT22, JB20, KHB^{+}23, MAR23, MWL^{+}23, MBB^{+}23, OWT^{+}23, CP18]. efficiency [CBF17, ZMB^{+}21]. efflux [LR18]. egg [WEW23]. either [RWRH16]. element [NDRL18]. Elemental [WGZ^{+}19, MB22]. elephant [RJS^{+}19]. Elevated [AOPP21]. elevation [CLB^{+}21, HBL^{+}24]. elusive [LSER23]. Embryos [VMM^{+}22]. emerge [OJ22]. Emergent [KHS^{+}19, VSB^{+}20]. Emerita [HGS20]. emission [VLD^{+}20]. emissions [BRK^{+}22, DBD18, WCA^{+}18]. encounter [LFS19]. end [Hop22]. endosymbiotic [CLTK^{+}22]. energy [KOLD^{+}21]. Engineered [FBDH18]. enhance [SFT^{+}17]. enrichment [CPW22, CHB^{+}22, LEvdM^{+}24, QRF^{+}18, YHD23]. environment [AOPP21, CEM^{+}22, DKZD16, LDWE24, MKKS17]. Environmental [DDF^{+}23, GKR^{+}23, CMS^{+}21, dlVMT^{+}21]. environmentally [HGS20]. environments [AS24, CSEP^{+}18, HPS^{+}18]. Ephemeral [PZC^{+}17]. epifluorescence [BFS20]. epiphytes [JEN^{+}20]. episodic [MHSD22]. equilibration [BHBT23]. erosion [DBG21]. Establishing [LBB^{+}24]. estimate [WCA^{+}18]. estimated [WRT^{+}22]. estimates [DRS18, FKG^{+}22, SGB^{+}17, YGSB22]. Estimating [SF23]. estimation [DTSC17]. estuaries [CQSZ22, YHD23]. estuarine [AOPP21, AAG^{+}20, MF23]. estuary [LDWE24, WHW^{+}20]. estuary-coast [LDWE24]. eukaryote [AGLD23]. eukaryotic [KLA^{+}17, WL19]. eutrophic [CDMH18, CAH^{+}20, DH21, RDFD18, RLH^{+}24, WHW^{+}20]. evaluate [BHBT23]. evasion [KVS^{+}24, RRSL^{+}19]. event [CDP^{+}18]. events [AOPP21, BCI17, MD23, TSB^{+}22]. Evidence [LCL22, RMV^{+}18, WEW23]. evolution [CQSZ22]. evolved [VBK^{+}19]. exceeds [DH21]. exchange [BH23, CSH^{+}22]. exchangeable [SMT^{+}17]. exometabolomics [STD^{+}18]. exopolymers [QPC^{+}16]. expedition [KHS^{+}21]. experiment [FT22, SNAA16]. experiments [CP18, GKR^{+}23, HSCA^{+}23, HMLC21]. explain [RTK^{+}18]. exploit [SFT^{+}17]. Exploring [THV23]. export [BRP^{+}24, CBF17, KCBG^{+}23, LFS19]. exposure [GDEK24, HHSL20]. Extent [BKM^{+}23, BRP^{+}24]. extracted [CBBE23]. Extreme [BCI17, Bri22, HPS^{+}18]. exudates [QRF^{+}18].

face [DBD18]. factors [KML^{+}23, SMS20]. fate [JZKS18, OAL^{+}19, QPC^{+}16, RCCGP20]. fates [SW23]. fatty [HLAW23]. fed [HLH23, WBK17]. Feeding [NLC20, FSA21, JWSD22]. fertilizer [SABR22]. fibers [HGS20]. fields [BH23]. figures [MBR23]. filling [DAS^{+}23]. Finding [MKKS17, YWH^{+}17]. finds [ASV^{+}21]. finfish [BSH^{+}20]. fire [BRP^{+}24, LSER23]. fire-bodies [LSER23]. fish [CC21, CBH^{+}22, GHHKY^{+}17, HOP20, LRCG20, ORA^{+}22, TSB^{+}22]. fishes [CMS^{+}21]. Five [KML^{+}23]. fixation [Ano22, CRlLC^{+}22, FRVD18, MFS22, SGB^{+}17]. fixing [CLTK^{+}22, KSL^{+}18, OWT^{+}23]. fjord [BRK^{+}22]. flagellates [SAE^{+}22]. flats [YSN^{+}23]. floats [SF23]. Flood [McI22]. floodplain [RWI^{+}21, STS^{+}17]. Florida [SABR22, VRF^{+}19]. Flow [dCCS21]. fluctuations [MFHMD20]. fluorescence [STD^{+}18]. Fluorescent [QRF^{+}18, PCS^{+}20]. flushes [CSH^{+}22]. flux [RDFD18]. fluxes [BWK^{+}23, CJ20, MF23, MHSD22, PKKL24, RH18, RRKLU^{+}20, SMC20, SW23, TFHK18]. following [CDP^{+}18, RLH^{+}24, YSN^{+}23]. food [AHM21, AAG^{+}20, BB17, HDSH^{+}22, MOW^{+}22]. food-web [HDSH^{+}22]. footprints [PZC^{+}17]. Foraging [CSC20, LFH^{+}23, RJS^{+}19, SFT^{+}17]. foraminifera [GYB^{+}23]. forcing [BFC^{+}23, ZKG^{+}22]. forests [OAH^{+}23]. formation [DHM17, PCS^{+}20, PMP^{+}20]. forward [SBJ^{+}23]. found [HPS^{+}18]. Founding [SDC21]. fractal [SH17]. Fragilariopsis [JB20]. framework [BHBT23, Lit23, YJJ^{+}23]. frequency [CAH^{+}20, TBB^{+}23, TSB^{+}22]. fresh [SBAPS20]. freshwater [Ano22, AFS^{+}23, AGLD23, ASV^{+}21, BSYD18, BSSB23, CJ20, DBVB20, FT22, FBV^{+}20, GAFS23, KML^{+}23, KCBG^{+}23, LKB^{+}22, MJC18, MFS22, MAR23, MCAD23, MWL^{+}23, MBB^{+}23, OAH^{+}23, RMV^{+}18, SJL23, SBAPS20, ZPRT23]. freshwater-marine [Ano22]. freshwaters [LLA^{+}22, WRC^{+}18]. fringing [CHLD18]. fronts [RJS^{+}19, SFT^{+}17]. Fulweiler [Ano22]. function [OTP23, VSB^{+}20]. future [BSH^{+}20, CRH^{+}23, DBG21, GKR^{+}23, WDT^{+}22].

Ganges [HBN^{+}22]. gas [DBD18]. gastropod [FSA21]. gelatinous [LSER23]. generates [SKDR21]. genome [DDF^{+}23]. geography [SLC18]. geospatial [VRB^{+}23]. giant [HCA^{+}23, SBM^{+}18]. global [DBD18, FOE20, LHMF20, LSER23, MYVD22, SGB^{+}17, Sme21, TCP18, VMF21, YJJ^{+}23, GDP^{+}23]. go [SBJ^{+}23]. goals [TFF23]. gradient [LSH^{+}22, OAL^{+}19]. grandis [KCPGE21]. Grazer [GK22]. Grazer-induced [GK22]. grazers [CP18, CPW22]. grazing [GRLS^{+}23, RGK22]. Great [CMR^{+}21, YSN^{+}23]. greater [HMLC21]. green [GRLS^{+}23]. Greenhouse [DBD18]. gross [AHL^{+}21, PEC^{+}23]. Groundwater [LGGF^{+}23, CDC^{+}21, DSA^{+}23, GYB^{+}23, HLH23, SFW^{+}23, ZXC^{+}24]. groundwater- [HLH23]. groundwater-creek-shelf [CDC^{+}21]. Groundwater-stream [LGGF^{+}23]. growth [AHM21, DBVB20, FKS20, GRLS^{+}23, HFKBW21, JB20, LSH^{+}22, LKB^{+}22, MD23, XSQ^{+}23]. guidance [LLA^{+}22]. Gulf [CMR^{+}24, YGSB22, YHD23, ZJL^{+}22].

habitat [MGG^{+}23, MLPM21]. habitats [ARG^{+}19, CTC^{+}20]. half [DAS^{+}23]. Halotia [DDF^{+}23]. hard [WEW23]. hard-water [WEW23]. hardwater [PEC^{+}23]. harmful [Lit23, XSQ^{+}23]. harvested [BSH^{+}20]. Hawai’i [CDP^{+}18, DSA^{+}23]. Headwater [LL19, BMH^{+}23, HLH23]. headwaters [FBDH18]. heatwave [TBB^{+}23]. help [KBB^{+}21]. hemiboreal [BWK^{+}23]. Herdmania [KCPGE21]. Heritage [CMR^{+}21]. Heroic [KHS^{+}21]. heterogeneity [SYM^{+}23]. heterogeneous [MKKS17]. heterotrophic [BOM^{+}23]. heterotrophy [JZKS18, MAE^{+}18]. Hidden [MGJ^{+}18, Inm22]. Hiding [KCPGE21]. High [BRK^{+}22, CLTK^{+}22, HRR24, CAH^{+}20, HBL^{+}24]. high-elevation [HBL^{+}24]. high-frequency [CAH^{+}20]. higher [RBvdK^{+}18, ZPRT23]. Highest [AHL^{+}21, VBK^{+}19]. highly [BH23]. hinders [Inm22]. historic [WDT^{+}22]. HNLC [GRLS^{+}23]. holobionts [PZB^{+}23]. honor [HDPEF23]. host [FRVD18]. Hot [BPL^{+}19, AS24]. hotspots [BH23]. humic [PCS^{+}20]. humic-like [PCS^{+}20]. hurricane [CBCD19]. hydro [OAL^{+}19]. hydro-climatic [OAL^{+}19]. hydroclimatic [VRB^{+}23]. Hydrologic [JZKS18, LR18]. Hydrology [BRP^{+}24]. hypolimnetic [HMLC21]. Hyporheic [KSL^{+}18]. hypoxia [BKM^{+}23].

Ice [DHM17, KVS^{+}24, LBB^{+}24, PLB^{+}17, CDG^{+}22, CLB^{+}21, DBdG^{+}18, KSD^{+}19, MCS18, ZTZL21]. ice-cover [ZTZL21]. ice-covered [CDG^{+}22, DBdG^{+}18, MCS18]. Ice-melt [KVS^{+}24]. Ichthyochory [GHHKY^{+}17]. ideas [SBAPS20]. identifiers [CSM^{+}21]. identifying [GBH20]. identity [AGM^{+}23]. imaging [SBAPS20]. Impact [LRD23, ZTZL21, CHLD18, MFHMD20, WEW23]. Impacts [HC20, SABR22, CSC20, MCAD23, PAMB23, SJL23, VH23]. impair [DBVB20]. Implications [MPB^{+}23, OAL^{+}19]. importance [LDWE24, MGJ^{+}18, SMS20]. impoundments [DBD18]. improved [FKG^{+}22]. Improving [GBH20]. inadequate [WLL22]. increase [BPL^{+}19, BB21, HAWH19, QRF^{+}18]. Increased [RLH^{+}24, GRLS^{+}23]. increases [DBRB20, HAWH19, RBvdK^{+}18]. Increasing [TBB^{+}23, WLB^{+}20, WRC^{+}18, HBL^{+}24, MAR23, MCAD23]. increasingly [LHMF20, PCS^{+}20]. independent [CBBE23, OAL^{+}19]. Indian [CRH^{+}23, LSH^{+}22, RBvdK^{+}18]. indicator [SNAA16]. indicators [BCP18]. induced [BMH^{+}23, GK22, HBL^{+}24, RWRH16, SMC20]. inequality [KBB^{+}21]. inequities [HDPEF23]. influence [CK18, DKZD16, DTSC17, FOE20, LL19, RCCGP20]. influences [BFC^{+}23, CBH^{+}22, LKB^{+}22, RH18]. information [KBB^{+}21, RBI^{+}23]. ingested [BFS20]. ingestion [CSC20, LRCG20]. inhibit [LJ24]. inhibitory [FT22]. inland [BB17, DRS18, HSH18, KCBG^{+}23, MIK18, SW23, SdG18, TCP18]. inorganic [GES19, RH18]. inputs [DRS18]. Insight [STD^{+}18]. insights [VSB^{+}20]. intake [NLC20]. Integrating [HDSH^{+}22, RBI^{+}23]. integrative [HSH18, SdG18, YJJ^{+}23]. intensifies [CHB^{+}22]. interact [NDRL18]. interacting [FKS20]. interaction [ORA^{+}22, YWH^{+}17]. Interactions [CK18, MGJ^{+}18, OJ22, RCCGP20, XSQ^{+}23]. Interactive [JB20, MAR23]. interannual [Bri22, HJS23]. interface [TFHK18]. intermittence [dCCS21]. Internal [TON^{+}20]. Interpreting [WLL22]. intertidal [CSH^{+}22, DDF^{+}23]. intracellular [KLA^{+}17]. Introduction [MCAD23]. intrusion [MWL^{+}23]. invasion [RLH^{+}24]. invasive [DBRB20]. inverse [WDT^{+}22]. invertebrate [PZB^{+}23, RBI^{+}23]. invertebrates [AGM^{+}23, HLAW23]. investigated [BFS20]. iron [ASW^{+}22, HBB^{+}16, JB20, LEvdM^{+}24, LWP22, TAC^{+}21]. iron-binding [HBB^{+}16]. iron-limited [ASW^{+}22]. irradiance [MGG^{+}23]. isolated [TCP18]. isomeric [HPS^{+}18]. isotopes [HCA^{+}23, OAL^{+}19]. issue [MCAD23].

J.T [Ano22]. Japan [YSN^{+}23, PKKL24]. Jerlov [WH23]. journal [Clo23, Sor16, SDC21].

Kāne’ohe [CDP^{+}18]. kelp [SBM^{+}18]. Key [HDC^{+}17, MFHMD20]. knowledge [BSH^{+}20].

L&O [Sor16, Clo22]. lacustrine [SABR22]. lagoon [ZKG^{+}22]. LAGOS [CSM^{+}21, KWRC21, RPM^{+}23]. LAGOS-US [CSM^{+}21, KWRC21, RPM^{+}23]. Lagrangian [XSQ^{+}23]. Lake [DR23, HSCA^{+}23, KWRC21, LBB^{+}24, SKDR21, BCI17, BWAJ19, CS22, CP18, CAH^{+}20, CDG^{+}22, CLB^{+}21, CRP^{+}18, DH21, DDH18, GRK^{+}17, HHSL20, HLH23, JZKS18, KSD^{+}19, Kri17, LRD23, MSSC19, PEC^{+}23, RDFD18, RLH^{+}24, RTK^{+}18, SNAA16, SLC18, SMS20, VLD^{+}20, VCL^{+}20, WLB^{+}20, WEW23, WDT^{+}22, ZTZL21, ZMB^{+}21, AE23, DAS^{+}23, DRKM23, MOW^{+}22]. lake-catchments [DDH18]. lake-fed [HLH23]. lakes [BPL^{+}19, CSM^{+}21, DBD18, DBdG^{+}18, DHM17, FKS20, GDEK24, HBL^{+}24, HDC^{+}17, ICH^{+}23, KVS^{+}24, LHL23, LL19, LBB^{+}24, MCS18, MAE^{+}18, PLB^{+}17, RPM^{+}23, RWRH16, SMSS19, STS^{+}17, SMP19, STW17, SDHJ23, SGSdGB17, TSB^{+}22, VH23, WLB^{+}20, WCP23, WRT^{+}22, dWCJB^{+}18]. Land [SH17, DSA^{+}23]. land-use [DSA^{+}23]. Landsat [MPB^{+}23]. Landscape [RRSL^{+}19, BWK^{+}23, CSW^{+}20, KOLD^{+}21]. large [DBRB20]. larger [RPM^{+}23]. lasting [PZC^{+}17]. late [CBCD19]. lateral [TFHK18]. latitudes [RBvdK^{+}18]. launching [SDC21]. lay [FJS24]. layers [PZC^{+}17]. leaf [dCCS21]. leakage [MHD21]. learned [LBB^{+}24]. learning [WRT^{+}22]. learnt [WCA^{+}18]. least [PEC^{+}23]. leave [PZC^{+}17]. left [MSSC19]. legacy [AAG^{+}20]. lentic [HSH18]. lessons [WCA^{+}18, LBB^{+}24]. lethal [KHB^{+}23]. Letters [Sor16, Clo22, SDC21]. level [AFS^{+}23, BGG^{+}23, CQSZ22, CST^{+}22, MHD21]. levels [SW20]. leveraging [WLB^{+}20]. life [ORA^{+}22]. ligand [HBB^{+}16]. Light [MFHMD20, ASW^{+}22, BMB24, FKS20, LEvdM^{+}24, LSER23]. light- [ASW^{+}22]. like [PCS^{+}20, RHN^{+}23, RCA^{+}21]. limitation [BSYD18, BOM^{+}23, BLZ^{+}22, FKS20]. limited [AHM21, ASW^{+}22, SMP19]. limiting [SKDR21]. limits [SDHJ23]. Limnol [Ano22]. limnology [DAS^{+}23, LHMF20, SDC21]. line [SFT^{+}17]. linked [HRR24, ICH^{+}23]. Linking [MH19, CLB^{+}21, RCA^{+}21]. lipids [HMIB22]. litter [dCCS21]. littoral [VH23]. live [GHHKY^{+}17]. living [HFKBW21, ZZD^{+}23]. loading [DRKM23, TON^{+}20]. location [CSM^{+}21]. LOCUS [CSM^{+}21]. long [CAH^{+}20, CLB^{+}21, CRP^{+}18, Kri17, LK24, LBB^{+}24, VRB^{+}23, WEW23]. long-read [DDF^{+}23]. long-term [CAH^{+}20, CLB^{+}21, CRP^{+}18, LBB^{+}24, VRB^{+}23, WEW23]. loss [BMH^{+}23, KSD^{+}19, LJ24]. losses [CLB^{+}21, HSCA^{+}23]. lotic [HSH18]. Low [CSW^{+}20, LSH^{+}22, RRKLU^{+}20, SBJ^{+}23, WCA^{+}18]. low-order [WCA^{+}18]. lower [HBN^{+}22, WSN^{+}18]. luminoxyscape [MGG^{+}23].

macroalgae [DSA^{+}23]. Macroalgal [JEN^{+}20]. macrobenthic [YSN^{+}23]. Macrocystis [SBM^{+}18]. macroinvertebrate [MOW^{+}22]. macrophytes [BB21]. MacroSheds [VRB^{+}23]. Magdalena [SSV^{+}22]. Magnitude [SGSdGB17, ZJL^{+}22]. main [NLC20]. Maine [ZJL^{+}22]. maintained [AHL^{+}21]. maintenance [FSA21]. major [AOPP21, CBCD19, Kri17]. management [TFF23]. manganese [LEvdM^{+}24]. Mangrove [SMT^{+}17, NLC20, SFT^{+}21]. mangroves [CDC^{+}21]. many [CBBE23]. Marcarelli [Ano22]. margin [JBG^{+}21]. marginal [SYM^{+}23]. Marine [RBvdK^{+}18, WML^{+}19, AHM21, Ano22, BHBT23, BH23, BOM^{+}23, CBH^{+}22, DSA^{+}23, GHHKY^{+}17, HC20, HDSH^{+}22, HBB^{+}16, KLA^{+}17, LOS^{+}20, LLA^{+}22, MFS22, MYVD22, MGJ^{+}18, RCCGP20, Sme21, SBAPS20, WGZ^{+}19]. marsh [CQSZ22, CST^{+}22, MWL^{+}23, MCGK22, YJJ^{+}23, ZXC^{+}24]. marshes [SJL23]. mass [STD^{+}18, TSB^{+}22]. masses [SYM^{+}23]. mat [KSL^{+}18]. matter [ASV^{+}21, BB17, BWAJ19, BOM^{+}23, CK18, CAO22, FBDH18, FKS20, FT22, HPS^{+}18, LOS^{+}20, PCS^{+}20, VS18, WRC^{+}18]. matters [dWCJB^{+}18]. maximum [CS22, DBRB20, LEvdM^{+}24]. meadow [JB22]. measured [ARG^{+}19, JB22, SF23]. measurements [AE23, CST^{+}22, VSB^{+}20]. measuring [Hei23, SBJ^{+}23]. mechanisms [PMP^{+}20, VBK^{+}19]. mediates [McI22]. meet [Clo22, MSSC19]. megaripple [BH23]. melt [KVS^{+}24]. Mercury [HCA^{+}23]. meridional [CJ20]. mesocosm [HSCA^{+}23]. mesocosms [ZZD^{+}23]. mesoplastics [MYVD22]. mesoscale [RJS^{+}19]. meta [HAWH19]. meta-analysis [HAWH19]. metabolic [BH23, CHB^{+}22, DTSC17]. metabolism [ARG^{+}19, HSH18, LGGF^{+}23, RCA^{+}21]. metacommunity [DKZD16]. metal [SAE^{+}22]. metallic [LWP22]. Methane [VLD^{+}20, BB21, BRK^{+}22, DBdG^{+}18, HMLC21, MF23, PZB^{+}23, WCA^{+}18, YAML^{+}23, ZTZL21, ZMB^{+}21]. Methanogenesis [DH21, BWAJ19]. method [MBB^{+}21]. methylmercury [HCA^{+}23]. metric [GDEK24]. metrics [KWRC21, TFF23]. Mexico [YGSB22, YHD23]. Michigan [DRKM23]. Micro [RCCGP20, MYVD22]. micro- [MYVD22]. Micro-by-micro [RCCGP20]. Microbial [DCPN20, LWP22, ZMB^{+}21, KSL^{+}18, MWL^{+}23, QPC^{+}16, UMDWC^{+}21, WSN^{+}18, WML^{+}19]. microbiome [FRVD18, JEN^{+}20]. microbiomes [MLPM21]. Microcystin [SW23]. micrometazoans [RMV^{+}18]. Micromonas [BIF20]. microorganisms [CK18, RCCGP20]. microparticle [NLC20]. Microplastic [BGHC20, BSH^{+}20, HOP20, LRCG20, WLC20, BFS20, HC20, HGS20, SW20]. Microplastics [GBH20, JWSD22, AAG^{+}20, DCPN20, LWP22, RCCGP20, YWH^{+}17]. microscopy [BFS20]. migration [MCGK22, XSQ^{+}23]. mineralization [KSL^{+}18]. minimum [JBG^{+}21]. Minnesota [VH23]. mismatch [SKDR21, THV23]. missing [YWH^{+}17]. mix [STW17]. mixed [CCT^{+}20]. mixing [BRK^{+}22, LRD23, LJ24]. mixotrophic [XSQ^{+}23]. mixotrophs [LDWE24]. mobilization [WML^{+}19]. model [DDH18, GDP^{+}23, HDC^{+}17, ORA^{+}22, XSQ^{+}23]. modeling [BCP18]. models [Cae21, CST^{+}22]. modifications [DSA^{+}23]. modify [HDC^{+}17, RJS^{+}19]. modulate [FRVD18]. modulating [MFHMD20]. module [CSM^{+}21]. mole [HGS20]. molecular [ASW^{+}22, SBAPS20]. monitoring [LK24]. morphology [BWK^{+}23]. mortality [HGS20, TSB^{+}22]. most [dWCJB^{+}18]. mothers [VMM^{+}22]. Motility [LFS19]. motions [WLL22]. mountain [BRP^{+}24, MH19, SMSS19]. mucus [Inm22]. Mud [KCBG^{+}23]. multiannual [WWFP20]. multifactorial [GKR^{+}23]. multiple [RDFD18, WCP23]. multiscale [OJ22]. mussel [HC20]. my [Clo23].

N [FRVD18, KSL^{+}18, OWT^{+}23]. N-fixing [KSL^{+}18, OWT^{+}23]. N-rich [OWT^{+}23]. naïve [AGLD23]. nanoparticles [CRP^{+}18]. national [WCA^{+}18]. Native [DBRB20, DSA^{+}23]. natural [CRP^{+}18, HC20, LSH^{+}22, LCL22, RPM^{+}23]. near [CEM^{+}22]. near-bed [CEM^{+}22]. nearshore [AS24]. need [Sor16, WWFP20]. neglected [BGG^{+}23]. net [FKG^{+}22, SF23, YGSB22]. network [AHL^{+}21, SdG18]. networks [FBDH18, KWRC21, KHS^{+}19, RRSL^{+}19]. Nevada [MBB^{+}23]. next [GBH20]. Nickel [CRlLC^{+}22]. night [DGS20, Hop22, SJL23]. Nimrod [KHS^{+}21]. nitrate [PLB^{+}17]. Nitrogen [Ano22, MFS22, SMP19, CRlLC^{+}22, CLTK^{+}22, FOE20, FBDH18, KCBG^{+}23, MJC18, MYS^{+}23, SGB^{+}17, SBM^{+}18]. nitrogen-fixing [CLTK^{+}22]. nitrous [JBG^{+}21, KCPGE21, MF23]. No [MSSC19, CRP^{+}18]. non [OWT^{+}23]. non-N-fixing [OWT^{+}23]. Nonlinear [VH23]. nonliving [HFKBW21]. north [CMR^{+}24, PLB^{+}17, TSB^{+}22, BSH^{+}20, BFS20, Bri22, HCA^{+}23, YBZ^{+}23]. north-east [CMR^{+}24]. northern [DBdG^{+}18, FKS20, GRK^{+}17, ICH^{+}23]. Northwest [BLZ^{+}22]. northwestern [YHD23]. Nostoc [KSL^{+}18]. novel [GHHKY^{+}17, VRF^{+}19, XSQ^{+}23]. nucleotides [KLA^{+}17]. Nutrient [BSYD18, BLZ^{+}22, CHB^{+}22, BB21, FKS20, Inm22, QRF^{+}18, SMP19, TAC^{+}21, WLB^{+}20]. nutrient-limited [SMP19]. nutrients [CP18, GAFS23, SKDR21]. Nutritional [BB17].

O [HBN^{+}22, VSB^{+}20]. obscures [DKZD16]. observations [CAH^{+}20, RDFD18, RBI^{+}23, WCA^{+}18]. occurrence [BSH^{+}20]. Ocean [FSA21, MYVD22, NDRL18, UMDWC^{+}21, YAML^{+}23, ZZD^{+}23, BMB24, BIF20, CBF17, Cae21, CBBE23, CMS^{+}21, CHFF22, GES19, HMIB22, LFS19, LSER23, MBB^{+}21, MGG^{+}23, PCS^{+}20, PKKL24, QPC^{+}16, RBI^{+}23, SAE^{+}22, SF23, TAC^{+}21, VMF21, YGSB22, ASW^{+}22, CRH^{+}23, GDP^{+}23, HCA^{+}23, LSH^{+}22, LEvdM^{+}24, OAH^{+}23, RBvdK^{+}18]. oceanic [MYS^{+}23, SMT^{+}17, WLL22]. Oceanogr [Ano22]. Oceanography [SDC21]. oceans [GDEK24, RH18]. off [CDC^{+}21]. offsets [VLD^{+}20]. oil [QPC^{+}16]. oligotrophic [LSH^{+}22, SYM^{+}23]. ongoing [GES19]. ontogeny [NDRL18]. open [CMS^{+}21, HDPEF23, SAE^{+}22, YGSB22, YCS^{+}21]. open-access [HDPEF23]. open-ocean [SAE^{+}22, YGSB22]. order [WCA^{+}18]. ordinances [SABR22]. Ordination [DKZD16]. Oregon [BGHC20]. Organic [BWAJ19, LOS^{+}20, AOPP21, ASV^{+}21, BRP^{+}24, BB17, BB21, BOM^{+}23, CDMH18, CK18, CAO22, FBDH18, FKS20, FT22, GES19, GRK^{+}17, HPS^{+}18, MIK18, PCS^{+}20, QRF^{+}18, SMT^{+}17, SKDR21, VS18, WCP23, WHW^{+}20, WRC^{+}18]. organic-matter [FT22]. organism [JWSD22]. organisms [GBH20, SW20]. organo [LWP22]. organo-metallic [LWP22]. origin [PE22]. outwelling [CDC^{+}21, SMT^{+}17]. overgrazing [DBG21]. overlooked [SFT^{+}21]. overturn [ZMB^{+}21]. overturning [CJ20]. oxidation [YHD23, ZMB^{+}21]. oxide [JBG^{+}21, KCPGE21, MF23]. Oxygen [CDMH18, MGG^{+}23, BH23, FSA21, JBG^{+}21, LHL23, PCS^{+}20]. oxygenation [BRK^{+}22, HMLC21]. oyster [UMDWC^{+}21].

P [LSH^{+}22, OWT^{+}23]. Pacific [HCA^{+}23, BFS20, BLZ^{+}22, HCA^{+}23, HGS20, JBG^{+}21, YBZ^{+}23]. Paired [VSB^{+}20]. partial [LCL22]. particle [CEM^{+}22, RRKLU^{+}20, XSQ^{+}23]. particles [DBVB20, HC20, LFS19, MHSD22, RRKLU^{+}20]. Particulate [GES19, HFKBW21]. passage [GHHKY^{+}17]. patches [MKKS17]. path [SBJ^{+}23]. pathways [KOLD^{+}21]. pattern [LGGF^{+}23]. Patterns [BFS20, BKM^{+}23, CMS^{+}21, HJS23, LK24, RRSL^{+}19]. pCO [HBN^{+}22, LR18]. peer [GD22]. perennially [CDG^{+}22]. performance [LLA^{+}22]. period [DBdG^{+}18, KVS^{+}24]. periphyton [MOW^{+}22, PAMB23]. permafrost [WRC^{+}18]. permeable [CHFF22]. persist [AOPP21]. persistence [MD23]. Persistent [AS24]. perspective [HSH18]. perspectives [CRH^{+}23]. pH [CTC^{+}20, MKKS17]. pH-taxis [MKKS17]. phenological [LK24]. phenology [SJL23]. phosphatase [DLT^{+}19]. phosphate [YBZ^{+}23]. Phosphonate [WL19]. Phosphorus [BOM^{+}23, KLA^{+}17, ICH^{+}23, KCBG^{+}23, MJC18, OJ22, TON^{+}20]. photodegradation [BMB24]. photolysis [GDP^{+}23]. Photomineralization [WHW^{+}20]. photosynthetic [THV23]. phyla [VCL^{+}20]. Physical [ZKG^{+}22, CSM^{+}21]. physiological [ASW^{+}22]. Phytoplankton [MB22, ARK^{+}23, ASW^{+}22, BSYD18, Bri22, BFC^{+}23, CPW22, CRP^{+}18, FKG^{+}22, GRLS^{+}23, HBB^{+}16, Hop22, Inm22, KLA^{+}17, LEvdM^{+}24, LCL22, MYS^{+}23, MHSD22, OTP23, PE22, TAC^{+}21, VBK^{+}19, VMF21, WL19, YCS^{+}21, ZJL^{+}22, ZKG^{+}22]. picoeukaryote [BIF20]. picoeukaryotic [VMF21]. picophytoplankton [LSH^{+}22]. piece [YWH^{+}17]. pipe [MJC18]. Pipes [dWCJB^{+}18]. plain [KCPGE21, MGJ^{+}18]. planktivorous [LRCG20]. plankton [DKZD16, FBV^{+}20, GAFS23, HJS23, MAR23, MGJ^{+}18, MBB^{+}23, MFHMD20, SBAPS20]. planktonic [AGLD23, LLA^{+}22]. plant [DBVB20]. Plastic [DBVB20, CSC20, Sme21, YWH^{+}17, ZZD^{+}23]. plasticity [VBK^{+}19]. Plateau [SJL23]. plateaux [GRLS^{+}23]. players [TCP18]. point [SW20]. points [RWI^{+}21]. polar [Hop22, KSL^{+}18]. pollutant [AAG^{+}20]. pollution [MOW^{+}22, SDHJ23, YWH^{+}17]. polyrhiza [DBVB20]. ponds [HRR24, KHS^{+}21, LJ24, TON^{+}20]. Pools [SW23, OAL^{+}19]. poorly [Ano22, MFS22]. population [FBV^{+}20, KHB^{+}23]. Pore [CSH^{+}22, CSEP^{+}18]. Pore-water [CSH^{+}22, CSEP^{+}18]. position [KOLD^{+}21]. post [BRP^{+}24]. post-fire [BRP^{+}24]. Potential [YBZ^{+}23, CSEP^{+}18, DLT^{+}19, FOE20, HBL^{+}24, MCGK22, VCL^{+}20, YHD23]. potexvirus [VRF^{+}19]. Powell [DAS^{+}23]. precipitation [CP18, ICH^{+}23, dWCJB^{+}18]. precipitation-driven [CP18]. predict [BWAJ19]. predictability [AFS^{+}23]. Predictable [OTP23, BWK^{+}23]. predicted [CC21]. predicting [Lit23]. predictions [WLB^{+}20]. predictive [KHB^{+}23]. predictor [GES19]. predicts [LHL23]. Preliminary [SGB^{+}17]. presence [JEN^{+}20]. present [DAS^{+}23]. press [MWL^{+}23]. pressure [ZTZL21]. prevalence [HOP20]. prey [HCA^{+}23]. primary [AHL^{+}21, Cae21, FKG^{+}22, MBB^{+}21, MFHMD20, PEC^{+}23, SMP19, SF23, YGSB22, YWH^{+}17, Hop22]. priorities [DBG21]. problem [DTSC17]. process [Ano22, MFS22, RRSL^{+}19]. processing [STD^{+}18, dCCS21]. Prochlorococcus [VMF21]. producers [YWH^{+}17]. production [Cae21, CCT^{+}20, FKG^{+}22, GES19, HBB^{+}16, HRR24, Hop22, Inm22, KCPGE21, LKB^{+}22, MCS18, MFHMD20, PEC^{+}23, SFT^{+}21, SMP19, SGSdGB17, YGSB22, ZPRT23]. productive [VLD^{+}20]. productivity [AHL^{+}21, BFC^{+}23, GRLS^{+}23, KHS^{+}19, MBB^{+}21, SF23]. profiles [WH23]. profiling [SF23]. progression [XSQ^{+}23]. progressive [KML^{+}23]. project [LBB^{+}24]. prokaryotes [BOM^{+}23]. protected [MSSC19]. protects [CRlLC^{+}22]. provide [VSB^{+}20]. providing [LLA^{+}22]. proxy [GYB^{+}23]. publication [Clo22, HDPEF23]. publishing [HDPEF23]. pulse [MWL^{+}23]. pulses [AOPP21]. puzzle [YWH^{+}17]. pyrifera [SBM^{+}18]. pyrosomes [LSER23].

quality [AHM21, BOM^{+}23, SABR22]. quantified [GDP^{+}23]. quantities [CBBE23]. quantity [AHM21]. Queensland [CMR^{+}21]. questions [GD22]. quotas [TAC^{+}21]. quotients [THV23].

R.W [Ano22]. radiation [VCL^{+}20]. Radon [CSEP^{+}18]. Radon-traced [CSEP^{+}18]. range [CQSZ22, CC21]. Rapid [BFC^{+}23]. rate [HC20]. rates [AHL^{+}21, BWAJ19, CLTK^{+}22, DTSC17, FSA21, GDP^{+}23, GRLS^{+}23, HFKBW21, HRR24, LSH^{+}22, LGGF^{+}23, WSN^{+}18]. rather [BRP^{+}24, MCGK22]. ratio [HFKBW21]. rationale [ZPRT23]. ratios [LSH^{+}22, MYS^{+}23, ZPRT23]. reactive [PCS^{+}20]. read [DDF^{+}23]. ready [Hop22]. recalcitrant [CAO22, PCS^{+}20]. receding [CSEP^{+}18]. Reconciling [CST^{+}22]. Reconsideration [YCS^{+}21]. record [DAS^{+}23]. Recovery [CDP^{+}18, YSN^{+}23]. Red [ZKG^{+}22]. redox [HBL^{+}24]. reduce [BB21, RGK22]. reduces [AFS^{+}23, CSW^{+}20, FSA21, FOE20]. reduction [LWP22]. Reef [CMR^{+}21, CDP^{+}18, CHLD18, CTC^{+}20, JWSD22, KCPGE21]. reef-scale [CDP^{+}18]. reefs [QRF^{+}18]. reflectance [MPB^{+}23]. regimes [BSSB23, HLH23, KHS^{+}19]. region [MCGK22]. regional [RMV^{+}18, YBZ^{+}23]. regions [RMV^{+}18]. regulate [PMP^{+}20]. regulates [FKS20, KLA^{+}17, ZJL^{+}22]. regulation [SGSdGB17]. related [DSA^{+}23, WML^{+}19]. relationship [CS22]. relationships [OJ22]. Relative [LDWE24, KHB^{+}23, SMS20]. release [BB21, KSL^{+}18, LWP22, PZB^{+}23]. released [BOM^{+}23]. relevant [HGS20]. rely [RRKLU^{+}20]. Remote [ZJL^{+}22]. removal [BHBT23, BB21]. Reproduction [CBH^{+}22, HGS20]. requirement [Clo22]. research [FJS24, GBH20, TFF23]. researchers [GD22]. reservoir [CDMH18, DAS^{+}23, RPM^{+}23, RPM^{+}23]. reservoirs [HDC^{+}17, HMLC21, HOP20]. residency [FT22]. residential [SABR22]. resilience [CPW22]. resolve [KBB^{+}21]. resolved [GDP^{+}23]. resource [MLPM21]. respiration [HAWH19, JB22, MCS18, MH19, TFHK18, WSN^{+}18]. Response [CHLD18, LEvdM^{+}24, BFC^{+}23, CSW^{+}20, HLH23, SMSS19]. responses [ASW^{+}22, CQSZ22, CEM^{+}22, DBRB20, WML^{+}19]. responsible [LFS19]. Results [PKKL24]. resurrection [ARK^{+}23]. retention [KSL^{+}18]. reveal [HCA^{+}23, HMLC21, OAL^{+}19]. revealed [DDF^{+}23]. reveals [FT22, SFT^{+}21]. review [GD22, LSER23]. rewards [SDC21]. rich [OWT^{+}23]. richness [RBvdK^{+}18]. riparian [DTSC17]. rise [BGG^{+}23, CQSZ22, CST^{+}22, MHD21]. Risk [DSA^{+}23, CSC20, DHM17, HHSL20]. risks [ARK^{+}23, SDC21]. River [SSV^{+}22, AHL^{+}21, DBRB20, KSD^{+}19, KHS^{+}19, RRSL^{+}19, HBN^{+}22, WSN^{+}18]. riverine [LHMF20, PE22, RH18]. rivers [BRP^{+}24, BKM^{+}23, BSSB23, GDEK24, LR18, TBB^{+}23, dCCS21]. road [SDHJ23]. robustness [ZMB^{+}21]. role [BGG^{+}23, CHFF22, DLT^{+}19, QPC^{+}16]. root [DTSC17]. roots [DBVB20]. Ross [YAML^{+}23]. rules [Hot20, SH17]. runoff [BWK^{+}23].

S [SYM^{+}23]. salinity [BSSB23, KHB^{+}23, MCAD23, OTP23, OWT^{+}23]. Salinization [LJ24, SFW^{+}23, AGLD23, DR23, HSCA^{+}23, KML^{+}23, KCBG^{+}23, LRD23, MBB^{+}23, WEW23]. salp [BFS20]. salp-ingested [BFS20]. salt [AFS^{+}23, AGLD23, SDHJ23, YJJ^{+}23, ZXC^{+}24]. salt-naïve [AGLD23]. saltmarsh [CSH^{+}22, KOLD^{+}21]. Salty [SBAPS20]. sampling [DGS20, WLL22]. sapidus [WLC20]. SAR11 [WGZ^{+}19]. scale [CDP^{+}18, CTC^{+}20, DBRB20, LFH^{+}23]. scales [HJS23, SFW^{+}23]. science [LBB^{+}24, MPB^{+}23]. sciences [KBB^{+}21, Sor16]. scientific [Clo23, Hot20, SDC21]. scientific-society [Clo23]. Scott [Ano22]. Sea [MHD21, MLPM21, YAML^{+}23, BHBT23, BGG^{+}23, CQSZ22, CMS^{+}21, CSH^{+}22, CST^{+}22, FT22, HJS23, SMC20, SYM^{+}23, CLTK^{+}22, DCPN20, PKKL24, YCS^{+}21, ZKG^{+}22]. Sea-level [MHD21, CQSZ22]. seabirds [CSC20, LFH^{+}23]. seafloor [LK24]. Seagrass [CMR^{+}24, CMR^{+}21, CBCD19, CCT^{+}20, DBG21, JB22, VRF^{+}19]. seals [RJS^{+}19, dlVMT^{+}21]. Seascape [CC21]. Seasonal [ARG^{+}19, CEM^{+}22, HBN^{+}22, Bri22, CBH^{+}22, HHSL20, HJS23]. seasonality [YCS^{+}21]. seasonally [HLAW23]. seawater [MWL^{+}23]. secret [KCPGE21]. Sediment [MCS18, ARK^{+}23, DBG21, MF23, WEW23]. sediment-based [ARK^{+}23]. sedimentary [Sme21]. sediments [BB21, BWAJ19, CSH^{+}22, CHFF22, DH21, GRK^{+}17, JBG^{+}21, MYVD22, PZB^{+}23, ZPRT23]. seiches [CDG^{+}22]. Selective [LFH^{+}23]. semi [BGG^{+}23]. semi-diurnal [BGG^{+}23]. semiarid [WHW^{+}20]. sensitivity [GES19, Hei23, LSH^{+}22, VMM^{+}22]. sensors [PE22, SBAPS20]. sequencing [DDF^{+}23]. series [BCI17, CP18]. services [KSD^{+}19]. setting [JZKS18]. sewage [MOW^{+}22]. Shackleton [KHS^{+}21]. shallow [ARG^{+}19, CEM^{+}22, CCT^{+}20, DHM17]. shape [AGM^{+}23, LGGF^{+}23]. shear [CEM^{+}22]. Shedding [BMB24, LSER23]. shelf [CDC^{+}21, CHFF22, HJS23, SMC20]. shell [NDRL18]. shellfish [BSH^{+}20]. shells [GYB^{+}23, RGK22]. shielding [BPL^{+}19]. shifts [CC21, HBN^{+}22, OTP23]. Should [DGS20, Clo23]. Show [MBR23, DBRB20, Kri17]. Siberia [MOW^{+}22]. Sierra [MBB^{+}23]. sight [KCPGE21, MGJ^{+}18]. signals [HDC^{+}17]. Significant [PKKL24, YAML^{+}23, STS^{+}17, SMT^{+}17]. silicate [ZJL^{+}22]. silver [CRP^{+}18]. similarities [ASV^{+}21]. Simple [Hot20, DDH18]. Single [KHB^{+}23]. Single-species [KHB^{+}23]. sink [DBG21, MYVD22, RDFD18]. siphonophores [HDSH^{+}22]. situ [PE22, YGSB22]. Size [ASV^{+}21, KOLD^{+}21, RBI^{+}23, SH17]. Size-based [ASV^{+}21]. size-distributions [SH17]. slicks [LFH^{+}23]. slope [SYM^{+}23]. slow [DR23]. slows [CC21]. Small [ORA^{+}22, LFH^{+}23, LJ24, RRKLU^{+}20, SNAA16, STW17, VLD^{+}20, VCL^{+}20, VRB^{+}23, WEW23, dWCJB^{+}18]. small-scale [LFH^{+}23]. smaller [ORA^{+}22]. snow [SMSS19]. Snowpack [SMS20, CSW^{+}20]. socialis [PMP^{+}20]. society [Clo23]. soils [ZPRT23]. solitary [KCPGE21]. Some [Hei23]. sorbed [AAG^{+}20]. soundscape [LK24]. source [CSEP^{+}18, JBG^{+}21, RDFD18, SW20, SMT^{+}17, SBM^{+}18, TON^{+}20]. sources [FBDH18, HCA^{+}23]. Southern [YAML^{+}23, ASW^{+}22, LEvdM^{+}24]. southwest [DAS^{+}23]. spanning [LBB^{+}24]. Sparkling [DR23]. spatial [BCP18, CMR^{+}21, CMR^{+}24, LGGF^{+}23, BGHC20]. Spatially [HJS23, LOS^{+}20]. spatiotemporal [SKDR21]. special [MCAD23]. specie [VMM^{+}22]. species [AFS^{+}23, BGHC20, CBCD19, HOP20, KHB^{+}23, MYS^{+}23, ORA^{+}22, RLH^{+}24, SNAA16]. spectral [SYM^{+}23]. spectrally [GDP^{+}23]. spectrometry [STD^{+}18]. spectroscopy [STD^{+}18]. Spirodela [DBVB20]. spore [PMP^{+}20]. spots [AS24]. spring [LRD23, SJL23, ZJL^{+}22]. stability [AGM^{+}23]. Stable [OAL^{+}19, HCA^{+}23]. stage [MKKS17]. stages [KML^{+}23, ORA^{+}22]. state [KML^{+}23, SMP19]. States [DAS^{+}23, TBB^{+}23, WRT^{+}22]. Station [HFKBW21]. statistical [BCP18]. status [RTK^{+}18]. steady [DR23]. stemflow [VS18]. steps [HDPEF23]. Stochastic [CAH^{+}20]. Stoichiometry [MJC18, MB22, WGZ^{+}19]. storage [Sme21]. Storm [SMC20, AOPP21, MHD21]. Storm-induced [SMC20]. stormwater [TON^{+}20]. Strait [CMR^{+}24]. strategies [JWSD22]. strategy [CSC20]. stratification [CMS^{+}21, LRD23, WDT^{+}22]. stratified [PEC^{+}23]. stream [AGM^{+}23, CHB^{+}22, DTSC17, FBDH18, FT22, LGGF^{+}23, McI22, SBJ^{+}23, SFW^{+}23, SW20]. stream-to-sea [FT22]. streams [BKM^{+}23, BSSB23, BMH^{+}23, CSW^{+}20, HLH23, KSL^{+}18, LR18, MH19, TBB^{+}23, WCA^{+}18]. strength [McI22, ORA^{+}22]. stress [CEM^{+}22]. strong [HBB^{+}16, ZZD^{+}23]. structure [DKZD16, MOW^{+}22, SNAA16, UMDWC^{+}21]. studies [ARK^{+}23, LLA^{+}22, VRB^{+}23]. study [CCT^{+}20, LLA^{+}22, TCP18]. Sub [RJS^{+}19, GRLS^{+}23]. sub-Antarctic [GRLS^{+}23]. Sub-mesoscale [RJS^{+}19]. submarine [GYB^{+}23]. submersed [BB21]. submesoscale [SFT^{+}17]. submit [Clo23]. substantially [HMLC21]. subtropical [BLZ^{+}22, YBZ^{+}23]. success [SFT^{+}17]. successional [CBCD19]. suited [TFF23]. Sulfate [YHD23]. sulfide [YHD23]. sulfur [LOS^{+}20]. summaries [FJS24]. summer [ICH^{+}23, SMS20, YAML^{+}23]. sunlight [CK18, GDEK24]. Superior [AE23]. superoxide [CRlLC^{+}22]. supply [BWAJ19, ZJL^{+}22]. support [BB17]. suppressed [RWRH16]. surface [CS22, HMIB22, MPB^{+}23, WRT^{+}22]. survival [FSA21]. suspended [BFS20]. Sustained [ZXC^{+}24]. sustains [KSL^{+}18]. Swedish [WCA^{+}18]. Synchronous [WCP23]. synchrony [HJS23, ICH^{+}23]. syndrome [KML^{+}23]. Synechococcus [VMF21]. Synergistic [BPL^{+}19]. Synthesis [CP18, CMR^{+}24, DBdG^{+}18, DRS18, VRB^{+}23]. Synthesizing [CMR^{+}21]. system [HLAW23]. systems [CAO22, CCT^{+}20, MB22].

Taking [HDPEF23]. Tampa [VRF^{+}19]. targets [MSSC19]. taxis [MKKS17]. Taxonomic [AGM^{+}23, TAC^{+}21]. tell [KHS^{+}21, MBR23]. temperate [AHL^{+}21, BWAJ19, CBH^{+}22, DBRB20, FBV^{+}20, KCPGE21, PLB^{+}17, SFW^{+}23, TSB^{+}22]. temperature [AE23, BWK^{+}23, CBF17, CTC^{+}20, HAWH19, Hei23, ICH^{+}23, JB20, LSH^{+}22, SJL23]. temperatures [DBRB20, WRT^{+}22]. temporal [BGHC20]. term [CAH^{+}20, CLB^{+}21, CRP^{+}18, LBB^{+}24, VRB^{+}23, WEW23]. Terrestrial [DRS18, GRK^{+}17, JZKS18, TFHK18]. terrigenous [WRC^{+}18]. tests [KHB^{+}23]. testudinum [VRF^{+}19]. Thalassia [VRF^{+}19]. thaw [WRC^{+}18]. their [CSM^{+}21, HLH23, LL19, MB22, MFHMD20]. theory [CS22, THV23]. Thermal [HLH23, LKB^{+}22, LLA^{+}22, VMM^{+}22, BSYD18, CSW^{+}20, Hei23, LCL22, SNAA16]. thermokarst [VCL^{+}20]. Thicker [RGK22]. thin [CLB^{+}21]. Thinking [RCA^{+}21]. third [PEC^{+}23]. thousands [WLB^{+}20]. thresholds [BCP18]. throughfall [VS18]. throughout [LFS19]. Tibetan [SJL23]. tidal [CQSZ22, MWL^{+}23, RWI^{+}21, YSN^{+}23, ZXC^{+}24]. tides [BGG^{+}23]. time [BCI17]. timescales [WCP23]. today [KHS^{+}21]. tolerance [AFS^{+}23, BSYD18]. tools [MIK18]. topographic [MCGK22]. topography [CC21, SH17]. tops [BPL^{+}19]. Torres [CMR^{+}24]. Total [SFT^{+}21, OJ22]. toxic [HHSL20]. toxicity [DHM17, KHB^{+}23]. toxin [LKB^{+}22]. toxins [OWT^{+}23]. Trace [SAE^{+}22, NDRL18]. traced [CSEP^{+}18]. tracers [SYM^{+}23, dlVMT^{+}21]. Tracking [OAH^{+}23, XSQ^{+}23]. trait [Lit23]. trait-based [Lit23]. trajectory [KHS^{+}21]. transfer [AAG^{+}20]. transformation [VCL^{+}20]. transformations [SMP19]. transitions [MAE^{+}18]. transport [RMV^{+}18]. trawling [FOE20]. Tree [VS18]. Tree-DOM [VS18]. trends [Kri17, RWRH16, VH23, WWFP20]. Tributary [DRKM23]. Trichodesmium [CRlLC^{+}22, FRVD18]. triphosphate [HFKBW21]. Trophic [AAG^{+}20, KOLD^{+}21, McI22, MFHMD20, RTK^{+}18, SMP19]. tropical [CHLD18, QRF^{+}18]. trout [McI22]. tuna [HCA^{+}23, ORA^{+}22]. Tunas [SFT^{+}17]. turbidity [ZTZL21]. turbulence [SMC20]. turnover [CLTK^{+}22]. twilight [LK24]. two [BGHC20, HOP20, KHB^{+}23, KSL^{+}18, ORA^{+}22]. type [ASV^{+}21]. types [KHB^{+}23, WH23].

U.S. [BSSB23, CSM^{+}21, HOP20, KWRC21, MSSC19, RPM^{+}23]. Ultra [VCL^{+}20]. Ultra-small [VCL^{+}20]. uncertainty [DRS18]. underlying [JBG^{+}21]. undersaturation [YAML^{+}23]. understand [WWFP20]. Understanding [Lit23, CRH^{+}23, DBG21, GBH20, SBAPS20]. understood [Ano22, MFS22]. understudied [LOS^{+}20]. unified [MOW^{+}22]. United [DAS^{+}23, TBB^{+}23, WRT^{+}22]. unresolved [CHFF22]. unveils [LK24]. Upper [SDHJ23]. Upscaling [DBD18]. uptake [Inm22, VLD^{+}20]. upward [ZXC^{+}24]. upwelling [BFC^{+}23]. urban [FBDH18, KCBG^{+}23, LJ24, SW20]. urchin [MLPM21]. Urea [SBM^{+}18]. use [DSA^{+}23, SBAPS20, SH17, TFF23]. using [DDF^{+}23, HFKBW21, PE22, RDFD18, WRT^{+}22]. utilization [WL19].

v1 [KWRC21]. v1.0 [CSM^{+}21]. Validity [MPB^{+}23]. value [Bri22]. Variability [Cae21, MCGK22, BGHC20, Bri22, CTC^{+}20, DLT^{+}19, GKR^{+}23, HLH23, RTK^{+}18]. Variability-based [Cae21]. variable [LOS^{+}20]. Variation [HLAW23, JB22, AFS^{+}23, CBH^{+}22, WCP23]. variations [HBN^{+}22, SMSS19, ZTZL21]. vary [MLPM21]. varying [HJS23, SMP19]. vascular [DBVB20]. vector [GHHKY^{+}17]. Vegetation [MAE^{+}18, DTSC17]. Velocity [WSN^{+}18]. Velocity-amplified [WSN^{+}18]. Vertical [CMS^{+}21, LJ24, SMC20, XSQ^{+}23]. vertically [GDP^{+}23]. via [PCS^{+}20]. virus [FBV^{+}20]. Viscosity [Inm22]. visual [MGG^{+}23]. vs [BGG^{+}23, CBCD19, HC20, MWL^{+}23]. vulnerability [CST^{+}22, VMM^{+}22].

Wajh [ZKG^{+}22]. wake [CBCD19]. warmer [OAL^{+}19, ZPRT23]. Warming [FBV^{+}20, HBL^{+}24, BPL^{+}19, BIF20, CHB^{+}22, FSA21, HHSL20, LJ24, MAR23, PAMB23, RWRH16, SMSS19, SMS20, TSB^{+}22]. Warming-induced [HBL^{+}24]. wastewater [MHD21]. water [CSEP^{+}18, CP18, CSH^{+}22, DDH18, JBG^{+}21, KVS^{+}24, LJ24, OAH^{+}23, PZC^{+}17, RWRH16, SFW^{+}23, SW20, SABR22, SYM^{+}23, SH17, TFF23, VH23, WLB^{+}20, WEW23, WH23]. waterbody [ASV^{+}21]. waters [DRS18, HSH18, MHD21, SW23, SdG18, TON^{+}20, TCP18]. watershed [VRB^{+}23]. watersheds [CSM^{+}21]. wave [CEM^{+}22]. wave- [CEM^{+}22]. web [HDSH^{+}22, MOW^{+}22]. webs [BB17]. well [HJS23, MSSC19]. well-connected [HJS23]. western [BSSB23, CLTK^{+}22, HCA^{+}23, YBZ^{+}23]. wetland [BGG^{+}23, KCBG^{+}23, WML^{+}19]. wetlands [DHM17]. Whole [HMLC21, CP18, SNAA16, SdG18]. Whole-ecosystem [HMLC21]. whole-lake [CP18, SNAA16]. wider [FJS24]. Widespread [AFS^{+}23, ICH^{+}23, SBJ^{+}23]. Wikipedia [KBB^{+}21]. wildfire [BRP^{+}24]. Wind [RTK^{+}18, BFC^{+}23]. Winter [WDT^{+}22, ICH^{+}23, PLB^{+}17]. Wisconsin [DR23]. within [AFS^{+}23, ASV^{+}21, RTK^{+}18]. World [CMR^{+}21, BKM^{+}23]. Worms [BB21]. write [FJS24]. writing [Hot20].

year [LK24]. year-long [LK24]. years [CMR^{+}21, RDFD18]. yields [PCS^{+}20].

zone [JBG^{+}21, WML^{+}19]. zooplankton [AFS^{+}23, DLT^{+}19, DBRB20, DGS20, HSCA^{+}23, RRKLU^{+}20, STW17, SGSdGB17].

References

Athey:2020:TTM

[AAG^{+}20] Samantha N. Athey, Samantha D. Albotra, Cessely A. Gordon, Bonnie Monteleone, Pamela Seaton, Anthony L. Andrady, Alison R. Taylor, and Susanne M. Brander. Trophic transfer of microplastics in an estuarine food chain and the effects of a sorbed legacy pollutant. Limnology and Oceanography Letters, 5(1): 154–162, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Austin:2023:ALS

[AE23] Jay Austin and Cassandra Elmer. An archive of Lake Superior temperature and current measurements, 2005–2020. Limnology and Oceanography Letters, 8(4):595–602, August 2023. CODEN ???? ISSN 2378-2242.

Arnott:2023:WVS

[AFS^{+}23] Shelley E. Arnott, Vincent Fugère, Celia C. Symons, Stephanie J. Melles, Beatrix E. Beisner, Miguel Cañedo-Argüelles, Marie-Pier Hébert, Jennifer A. Brentrup, Amy L. Downing, Derek K. Gray, Danielle Greco, William D. Hintz, Alexandra McClymont, Rick A. Relyea, James A. Rusak, Catherine L. Searle, Louis Astorg, Henry K. Baker, Zeynep Ersoy, Carmen Espinosa, Jaclyn M. Franceschini, Angelina T. Giorgio, Norman Göbeler, Emily Hassal, Mercedes Huynh, Samuel Hylander, Kacie L.

Jonasen, Andrea Kirkwood, Silke Langenheder, Ola Langvall, Hjalmar Laudon, Lovisa Lind, Maria Lundgren, Emma R. Moffett, Lorenzo Proia, Matthew S. Schuler, Jonathan B. Shurin, Christopher F. Steiner, Maren Striebel, Simon Thibodeau, Pablo Urrutia Cordero, Lidia Vendrell-Puigmitja, Gesa A. Weyhenmeyer, and Alison M. Derry. Widespread variation in salt tolerance within freshwater zooplankton species reduces the predictability of community-level salt tolerance. Limnology and Oceanography Letters, 8(1):8–18, February 2023. CODEN ???? ISSN 2378-2242.

Astorg:2023:EFS

[AGLD23] Louis Astorg, Jean-Christophe Gagnon, Cassandre S. Lazar, and Alison M. Derry. Effects of freshwater salinization on a salt-naïve planktonic eukaryote community. Limnology and Oceanography Letters, 8(1):38–47, February 2023. CODEN ???? ISSN 2378-2242.

Allen:2023:TIB

[AGM^{+}23] Daniel C. Allen, Brian A. Gill, Anya Metcalfe, Sophia Bonjour, Scott Starr, Junna Wang, Diana Valentin, and Nancy B. Grimm. Taxonomic identity, biodiversity, and antecedent disturbances shape the dimensional stability of stream invertebrates. Limnology and Oceanography Letters, 8(3):464–472, June 2023. CODEN ???? ISSN 2378-2242.

Aho:2021:HRG

[AHL^{+}21] Kelly S. Aho, Jacob D. Hosen, Laura A. Logozzo, Wade R. McGillis, and Peter A. Raymond. Highest rates of gross primary productivity maintained despite CO_{2} depletion in a temperate river network. Limnology and Oceanography Letters, 6(4):200– 206, August 2021. CODEN LOLIBT. ISSN 2378-2242.

Anderson:2021:GMC

[AHM21] Thomas R. Anderson, Dag O. Hessen, and Daniel J. Mayor. Is the growth of marine copepods limited by food quantity or quality? Limnology and Oceanography Letters, 6(3):127–133, June 2021. CODEN LOLIBT. ISSN 2378-2242.

Anonymous:2022:CMM

[Ano22] Anonymous. Corrigendum for: Marcarelli, A.M., Fulweiler, R.W. and Scott, J.T. 2019. Nitrogen fixation: A poorly understood process along the freshwater-marine continuum. Limnol. Oceanogr. 7(1): 1–10. Limnology and Oceanography Letters, 7 (5):450, October 2022. CODEN ???? ISSN 2378-2242. See [MFS22].

Asmala:2021:EOC

[AOPP21] Eero Asmala, Christopher L. Osburn, Ryan W. Paerl, and Hans W. Paerl. Elevated organic carbon pulses persist in estuarine environment after major storm events. Limnology and Oceanography Letters, 6(1):43–50, February 2021. CODEN LOLIBT. ISSN 2378-2242.

Attard:2019:SEM

[ARG^{+}19] Karl M. Attard, Iván F. Rodil, Ronnie N. Glud, Peter Berg, Joanna Norkko, and Alf Norkko. Seasonal ecosystem metabolism across shallow benthic habitats measured by aquatic eddy covariance. Limnology and Oceanography Letters, 4(3):79– 86, June 2019. CODEN LOLIBT. ISSN 2378-2242.

Andersson:2023:CCR

[ARK^{+}23] Björn Andersson, Karin Rengefors, Olga Kourtchenko, Kerstin Johannesson, Olof Berglund, and Helena L. Filipsson. Crosscontamination risks in sediment-based resurrection studies of phytoplankton. Limnology and Oceanography Letters, 8(2):376– 384, April 2023. CODEN ???? ISSN 2378-2242.

Asmala:2024:PHS

[AS24] Eero Asmala and Matias Scheinin. Persistent hot spots of CO_{2} and CH_{4} in coastal nearshore environments. Limnology and Oceanography Letters, 9(2):119–127, April 2024. CODEN ???? ISSN 2378-2242.

Aukes:2021:SBC

[ASV^{+}21] Pieter J. K. Aukes, Sherry L. Schiff, Jason J. Venkiteswaran, Richard J. Elgood, and John Spoelstra. Size-based characterization of freshwater dissolved organic matter finds similarities within a waterbody type across different Canadian ecozones. Limnology and Oceanography Letters, 6(2):85–95, April 2021. CODEN LOLIBT. ISSN 2378-2242.

Andrew:2022:DPM

[ASW^{+}22] Sarah M. Andrew, Robert F. Strzepek, Spencer M. Whitney, Wah Soon Chow, and Michael J. Ellwood. Divergent physiological and molecular responses of light- and iron-limited Southern Ocean phytoplankton. Limnology and Oceanography Letters, 7 (2):150–158, April 2022. CODEN LOLIBT. ISSN 2378-2242.

Bellamy:2017:NSI

[BB17] Amber R. Bellamy and James E. Bauer. Nutritional support of inland aquatic food webs by aged carbon and organic matter. Limnology and Oceanography Letters, 2(5):131–149, October 2017. CODEN LOLIBT. ISSN 2378-2242.

Benelli:2021:WSM

[BB21] Sara Benelli and Marco Bartoli. Worms and submersed macrophytes reduce methane release and increase nutrient removal in organic sediments. Limnology and Oceanography Letters, 6(6): 329–338, December 2021. CODEN LOLIBT. ISSN 2378-2242.

Batt:2017:EEL

[BCI17] Ryan D. Batt, Stephen R. Carpenter, and Anthony R. Ives. Extreme events in lake ecosystem time series. Limnology and Oceanography Letters, 2(3):63–69, June 2017. CODEN LOLIBT. ISSN 2378-2242.

Buelo:2018:MAS

[BCP18] C. D. Buelo, S. R. Carpenter, and M. L. Pace. A modeling analysis of spatial statistical indicators of thresholds for algal blooms. Limnology and Oceanography Letters, 3(5):384–392, October 2018. CODEN LOLIBT. ISSN 2378-2242.

Broullon:2023:RPR

[BFC^{+}23] Esperanza Broullón, Peter J. S. Franks, Bieito Fernández Castro, Miguel Gilcoto, Antonio Fuentes-Lema, Maria Pérez-Lorenzo, Emilio Fernández, and Beatriz Mouriño-Carballido. Rapid phytoplankton response to wind forcing influences productivity in upwelling bays. Limnology and Oceanography Letters, 8(3):529–537, June 2023. CODEN ???? ISSN 2378-2242.

Brandon:2020:PSS

[BFS20] Jennifer A. Brandon, Alexandra Freibott, and Linsey M. Sala. Patterns of suspended and salp-ingested microplastic debris in the North Pacific investigated with epifluorescence microscopy. Limnology and Oceanography Letters, 5(1):46–53, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Belliard:2023:CWA

[BGG^{+}23] Jean-Philippe Belliard, Olivier Gourgue, Gerard Govers, Matthew L. Kirwan, and Stijn Temmerman. Coastal wetland adaptability to sea level rise: The neglected role of semi-diurnal vs. diurnal tides. Limnology and Oceanography Letters, 8(2): 340–349, April 2023. CODEN ???? ISSN 2378-2242.

Baechler:2020:MCT

[BGHC20] Britta R. Baechler, Elise F. Granek, Matthew V. Hunter, and Kathleen E. Conn. Microplastic concentrations in two Oregon bivalve species: Spatial, temporal, and species variability. Limnology and Oceanography Letters, 5(1):54–65, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Berg:2023:CMM

[BH23] Peter Berg and Markus Huettel. Coastal marine megaripple fields are metabolic hotspots with highly dynamic oxygen exchange. Limnology and Oceanography Letters, 8(6):867–875, December 2023. CODEN ???? ISSN 2378-2242.

Bach:2023:TCF

[BHBT23] Lennart T. Bach, David T. Ho, Philip W. Boyd, and Michael D. Tyka. Toward a consensus framework to evaluate air–sea CO_{2} equilibration for marine CO_{2} removal. Limnology and Oceanography Letters, 8(5):685–691, October 2023. CODEN ???? ISSN 2378-2242.

Benner:2020:CCA

[BIF20] Ina Benner, Andrew J. Irwin, and Zoe V. Finkel. Capacity of the common Arctic picoeukaryote Micromonas to adapt to a warming ocean. Limnology and Oceanography Letters, 5(2): 221–227, April 2020. CODEN LOLIBT. ISSN 2378-2242.

Blaszczak:2023:EPD

[BKM^{+}23] Joanna R. Blaszczak, Lauren E. Koenig, Francine H. Mejia, Lluís Gómez-Gener, Christopher L. Dutton, Alice M. Carter, Nancy B. Grimm, Judson W. Harvey, Ashley M. Helton, and Matthew J. Cohen. Extent, patterns, and drivers of hypoxia in the world’s streams and rivers. Limnology and Oceanography Letters, 8(3):453–463, June 2023. CODEN ???? ISSN 2378-2242.

Browning:2022:NCL

[BLZ^{+}22] Thomas J. Browning, Xin Liu, Ruifeng Zhang, Zuozhu Wen, Jing Liu, Yaqian Zhou, Feipeng Xu, Yihua Cai, Kuanbo Zhou, Zhimian Cao, Yuanli Zhu, Dalin Shi, Eric P. Achterberg, and Minhan Dai. Nutrient co-limitation in the subtropical Northwest Pacific. Limnology and Oceanography Letters, 7(1):52–61, February 2022. CODEN LOLIBT. ISSN 2378-2242.

Bannon:2024:SLC

[BMB24] Catherine C. Bannon, Elizabeth M. Mudge, and Erin M. Bertrand. Shedding light on cobalamin photodegradation in the ocean. Limnology and Oceanography Letters, 9(2):135–144, April 2024. CODEN ???? ISSN 2378-2242.

Brusecke:2023:DIB

[BMH^{+}23] Joanna Brüsecke, Timo Muotka, Kaisa-Leena Huttunen, Sanni Litjo, Wille-Pekka Lepo, and Jussi Jyväsjärvi. Drainage-induced browning causes both loss and change of benthic biodiversity in headwater streams. Limnology and Oceanography Letters, 8(4): 620–627, August 2023. CODEN ???? ISSN 2378-2242.

Bouchachi:2023:PLD

[BOM^{+}23] Nawal Bouchachi, Ingrid Obernosterer, Barbara Marie, Olivier Crispi, and Eva Ortega-Retuerta. Phosphorus limitation determines the quality of dissolved organic matter released by marine heterotrophic prokaryotes. Limnology and Oceanography Letters, 8(2):330–339, April 2023. CODEN ???? ISSN 2378-2242.

Bartosiewicz:2019:HTC

[BPL^{+}19] Maciej Bartosiewicz, Anna Przytulska, Jean-François Lapierre, Isabelle Laurion, Moritz F. Lehmann, and Roxane Maranger. Hot tops, cold bottoms: Synergistic climate warming and shielding effects increase carbon burial in lakes. Limnology and Oceanography Letters, 4(5):132–144, October 2019. CODEN LOLIBT. ISSN 2378-2242.

Britten:2022:EVD

[Bri22] Gregory L. Britten. Extreme value distributions describe interannual variability in the seasonal North Atlantic phytoplankton bloom. Limnology and Oceanography Letters, 7(3):269–276, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Bonaglia:2022:HME

[BRK^{+}22] Stefano Bonaglia, Tobias Rütting, Mikhail Kononets, Anders Stigebrandt, Isaac R. Santos, and Per O. J. Hall. High methane emissions from an anoxic fjord driven by mixing and oxygenation. Limnology and Oceanography Letters, 7(5):392–400, October 2022. CODEN ???? ISSN 2378-2242.

Barton:2024:HRT

[BRP^{+}24] Riley Barton, Christina M. Richardson, Evelyn Pae, Maya S. Montalvo, Michael Redmond, Margaret A. Zimmer, and Sasha Wagner. Hydrology, rather than wildfire burn extent, determines post-fire organic and black carbon export from mountain rivers in central coastal California. Limnology and Oceanography Letters, 9(1):70–80, February 2024. CODEN ???? ISSN 2378-2242.

Baechler:2020:MOE

[BSH^{+}20] Britta R. Baechler, Cheyenne D. Stienbarger, Dorothy A. Horn, Jincy Joseph, Alison R. Taylor, Elise F. Granek, and Susanne M. Brander. Microplastic occurrence and effects in commercially harvested North American finfish and shellfish: Current knowledge and future directions. Limnology and Oceanography Letters, 5(1):113–136, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Bolotin:2023:CFS

[BSSB23] Lauren A. Bolotin, Betsy M. Summers, Philip Savoy, and Joanna R. Blaszczak. Classifying freshwater salinity regimes in central and western U.S. streams and rivers. Limnology and Oceanography Letters, 8(1):103–111, February 2023. CODEN ???? ISSN 2378-2242.

Bestion:2018:NLC

[BSYD18] Elvire Bestion, C-Elisa Schaum, and Gabriel Yvon-Durocher. Nutrient limitation constrains thermal tolerance in freshwater phytoplankton. Limnology and Oceanography Letters, 3(6):436– 443, December 2018. CODEN LOLIBT. ISSN 2378-2242.

Bertolet:2019:OMS

[BWAJ19] Brittni L. Bertolet, William E. West, David W. Armitage, and Stuart E. Jones. Organic matter supply and bacterial community composition predict methanogenesis rates in temperate lake sediments. Limnology and Oceanography Letters, 4(5):164–172, October 2019. CODEN LOLIBT. ISSN 2378-2242.

Balathandayuthabani:2023:ACF

[BWK^{+}23] Sivakiruthika Balathandayuthabani, Marcus B. Wallin, Leif Klemedtsson, Patrick Crill, and David Bastviken. Aquatic carbon fluxes in a hemiboreal catchment are predictable from landscape morphology, temperature, and runoff. Limnology and Oceanography Letters, 8(2):313–322, April 2023. CODEN ???? ISSN 2378-2242.

Cael:2021:VBC

[Cae21] B. B. Cael. Variability-based constraint on ocean primary production models. Limnology and Oceanography Letters, 6(5):262– 269, October 2021. CODEN LOLIBT. ISSN 2378-2242.

Carpenter:2020:SDC

[CAH^{+}20] Stephen R. Carpenter, Babak M. S. Arani, Paul C. Hanson, Marten Scheffer, Emily H. Stanley, and Egbert Van Nes. Stochastic dynamics of Cyanobacteria in long-term highfrequency observations of a eutrophic lake. Limnology and Oceanography Letters, 5(5):331–336, October 2020. CODEN LOLIBT. ISSN 2378-2242.

Cotner:2022:ARD

[CAO22] James B. Cotner, N. J. Anderson, and Christopher Osburn. Accumulation of recalcitrant dissolved organic matter in aerobic aquatic systems. Limnology and Oceanography Letters, 7(5): 401–409, October 2022. CODEN ???? ISSN 2378-2242.

Cael:2023:HMI

[CBBE23] B. B. Cael, Kelsey Bisson, Emmanuel Boss, and Zachary K. Erickson. How many independent quantities can be extracted from ocean color? Limnology and Oceanography Letters, 8(4): 603–610, August 2023. CODEN ???? ISSN 2378-2242.

Congdon:2019:WMH

[CBCD19] Victoria M. Congdon, Christina Bonsell, Meaghan R. Cuddy, and Kenneth H. Dunton. In the wake of a major hurricane: Differential effects on early vs. late successional seagrass species. Limnology and Oceanography Letters, 4(5):155–163, October 2019. CODEN LOLIBT. ISSN 2378-2242.

Cael:2017:HRT

[CBF17] B. B. Cael, Kelsey Bisson, and Michael J. Follows. How have recent temperature changes affected the efficiency of ocean biological carbon export? Limnology and Oceanography Letters, 2 (4):113–118, August 2017. CODEN LOLIBT. ISSN 2378-2242.

Collins:2022:RIS

[CBH^{+}22] Rupert A. Collins, Charles Baillie, Nicholas C. Halliday, Sophie Rainbird, David W. Sims, Stefano Mariani, and Martin J.

Genner. Reproduction influences seasonal eDNA variation in a temperate marine fish community. Limnology and Oceanography Letters, 7(5):443–449, October 2022. CODEN ???? ISSN 2378-2242.

Champion:2021:STS

[CC21] Curtis Champion and Melinda A. Coleman. Seascape topography slows predicted range shifts in fish under climate change. Limnology and Oceanography Letters, 6(3):143–153, June 2021. CODEN LOLIBT. ISSN 2378-2242.

Cox:2020:DDD

[CCT^{+}20] T. Erin Cox, Just Cebrian, Marnie Tabor, Laura West, and Jeffrey W. Krause. Do diatoms dominate benthic production in shallow systems? A case study from a mixed seagrass bed. Limnology and Oceanography Letters, 5(6):425–434, December 2020. CODEN LOLIBT. ISSN 2378-2242.

Cabral:2021:CAO

[CDC^{+}21] Alex Cabral, Thorsten Dittmar, Mitchell Call, Jan Scholten, Carlos E. de Rezende, Nils Asp, Martha Gledhill, Michael Seidel, and Isaac R. Santos. Carbon and alkalinity outwelling across the groundwater-creek-shelf continuum off Amazonian mangroves. Limnology and Oceanography Letters, 6(6):369–378, December 2021. CODEN LOLIBT. ISSN 2378-2242.

Castendyk:2022:BSP

[CDG^{+}22] Devin N. Castendyk, Hilary A. Dugan, Hugh A. Gallagher, Nimish Pujara, Peter T. Doran, John C. Priscu, and W. Berry Lyons. Barotropic seiches in a perennially ice-covered lake, East Antarctica. Limnology and Oceanography Letters, 7(1):26–33, February 2022. CODEN LOLIBT. ISSN 2378-2242.

Carey:2018:ODC

[CDMH18] Cayelan C. Carey, Jonathan P. Doubek, Ryan P. McClure, and Paul C. Hanson. Oxygen dynamics control the burial of organic carbon in a eutrophic reservoir. Limnology and Oceanography Letters, 3(3):293–301, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Courtney:2018:RRS

[CDP^{+}18] T. A. Courtney, E. H. De Carlo, H. N. Page, K. D. Bahr, A. Barro, N. Howins, R. Tabata, G. Terlouw, K. S. Rodgers, and A. J. Andersson. Recovery of reef-scale calcification following a bleaching event in Kāne’ohe Bay, Hawai’i. Limnology and Oceanography Letters, 3(1):1–9, February 2018. CODEN LOLIBT. ISSN 2378-2242.

Chang:2022:SPR

[CEM^{+}22] Grace Chang, Galen Egan, Joseph D. McNeil, Samuel McWilliams, Craig Jones, Frank Spada, Stephen Monismith, and Oliver Fringer. Seasonal particle responses to near-bed shear stress in a shallow, wave- and current-driven environment. Limnology and Oceanography Letters, 7(2):175–183, April 2022. CODEN LOLIBT. ISSN 2378-2242.

Cross:2022:NEI

[CHB^{+}22] Wyatt F. Cross, James M. Hood, Jonathan P. Benstead, Alexander D. Huryn, Jill R. Welter, Gísli M. Gíslason, and Jón S. Ólafsson. Nutrient enrichment intensifies the effects of warming on metabolic balance of stream ecosystems. Limnology and Oceanography Letters, 7(4):332–341, August 2022. CODEN LOLIBT. ISSN 2378-2242.

Chua:2022:CAU

[CHFF22] Emily J. Chua, Markus Huettel, Katja Fennel, and Robinson W. Fulweiler. A case for addressing the unresolved role of permeable shelf sediments in ocean denitrification. Limnology and Oceanography Letters, 7(1):11–25, February 2022. CODEN LOLIBT. ISSN 2378-2242.

Cuttler:2018:RFR

[CHLD18] M. V. W. Cuttler, J. E. Hansen, R. J. Lowe, and E. J. F. Drost. Response of a fringing reef coastline to the direct impact of a tropical cyclone. Limnology and Oceanography Letters, 3(2):31– 38, April 2018. CODEN LOLIBT. ISSN 2378-2242.

Cael:2020:FFA

[CJ20] B. B. Cael and Malte F. Jansen. On freshwater fluxes and the Atlantic meridional overturning circulation. Limnology and Oceanography Letters, 5(2):185–192, April 2020. CODEN LOLIBT. ISSN 2378-2242.

Cory:2018:IBS

[CK18] Rose M. Cory and George W. Kling. Interactions between sunlight and microorganisms influence dissolved organic matter degradation along the aquatic continuum. Limnology and Oceanography Letters, 3(3):102–116, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Christianson:2021:TIL

[CLB^{+}21] Kyle R. Christianson, Kelly A. Loria, Peter D. Blanken, Nel Caine, and Pieter T. J. Johnson. On thin ice: Linking elevation and long-term losses of lake ice cover. Limnology and Oceanography Letters, 6(2):77–84, April 2021. CODEN LOLIBT. ISSN 2378-2242.

Cloern:2022:HMD

[Clo22] James Cloern. How to meet the data publication requirement for L&O Letters. Limnology and Oceanography Letters, 7(3): 189–190, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Cloern:2023:WSS

[Clo23] James Cloern. Why should I submit my article to a scientificsociety journal? Limnology and Oceanography Letters, 8(6): 799–803, December 2023. CODEN ???? ISSN 2378-2242.

Cheung:2022:HBT

[CLTK^{+}22] Shunyan Cheung, Kailin Liu, Kendra A. Turk-Kubo, Jun Nishioka, Koji Suzuki, Michael R. Landry, Jonathan P. Zehr, Szeki Leung, Lixia Deng, and Hongbin Liu. High biomass turnover rates of endosymbiotic nitrogen-fixing cyanobacteria in the western Bering Sea. Limnology and Oceanography Letters, 7(6):501– 509, December 2022. CODEN ???? ISSN 2378-2242.

Carter:2021:SYS

[CMR^{+}21] Alex B. Carter, Skye A. McKenna, Michael A. Rasheed, Catherine Collier, Len McKenzie, Roland Pitcher, and Rob Coles. Synthesizing 35 years of seagrass spatial data from the Great Barrier Reef World Heritage Area, Queensland, Australia. Limnology and Oceanography Letters, 6(4):216–226, August 2021. CODEN LOLIBT. ISSN 2378-2242.

Carter:2024:SSD

[CMR^{+}24] A. Carter, S. McKenna, Ma Rasheed, H. Taylor, C. van de Wetering, K. Chartrand, C. Reason, C. Collier, L. Shepherd, J. Mellors, L. McKenzie, Nc Duke, A. Roelofs, N. Smit, R. Groom, D. Barrett, S. Evans, R. Pitcher, N. Murphy, M. Carlisle, M. David, S. Lui, Torres Strait Indigenous Rangers, and Rg Coles. Seagrass spatial data synthesis from north-east Australia, Torres Strait and Gulf of Carpentaria, 1983 to 2022.

Limnology and Oceanography Letters, 9(1):7–22, February 2024. CODEN ???? ISSN 2378-2242.

Canals:2021:VSE

[CMS^{+}21] Oriol Canals, Iñaki Mendibil, María Santos, Xabier Irigoien, and Naiara Rodríguez-Ezpeleta. Vertical stratification of environmental DNA in the open ocean captures ecological patterns and behavior of deep-sea fishes. Limnology and Oceanography Letters, 6(6):339–347, December 2021. CODEN LOLIBT. ISSN 2378-2242.

Carpenter:2018:SSW

[CP18] Stephen R. Carpenter and Michael L. Pace. Synthesis of a 33-yr series of whole-lake experiments: Effects of nutrients, grazers, and precipitation-driven water color on chlorophyll. Limnology and Oceanography Letters, 3(6):419–427, December 2018. CO-DEN LOLIBT. ISSN 2378-2242.

Carpenter:2022:DGR

[CPW22] Stephen R. Carpenter, Michael L. Pace, and Grace M. Wilkinson. DOC, grazers, and resilience of phytoplankton to enrichment. Limnology and Oceanography Letters, 7(6):466–474, December 2022. CODEN ???? ISSN 2378-2242.

Cai:2022:BRT

[CQSZ22] Xun Cai, Qubin Qin, Jian Shen, and Yinglong J. Zhang. Bifurcate responses of tidal range to sea-level rise in estuaries with marsh evolution. Limnology and Oceanography Letters, 7(3): 210–217, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Chowdhury:2023:DIO

[CRH^{+}23] Subhadeep Chowdhury, Eric Raes, Cora Hörstmann, Ayaz Ahmed, Céline Ridame, Nicolas Metzl, P S Bhavya, Takuya Sato, Takuhei Shiozaki, Sophie Bonnet, Carolin R. Löscher, Arvind Singh, and Mar Benavides. Diazotrophy in the Indian Ocean: Current understanding and future perspectives. Limnology and Oceanography Letters, 8(5):707–722, October 2023. CODEN ???? ISSN 2378-2242.

Chen:2022:NSD

[CRlLC^{+}22] Chih-Chang Chen, Irene B. Rodriguez, Yuh ling Lee Chen, Jonathan P. Zehr, Yun-Ru Chen, Shang-Te Danny Hsu, Shun-Chung Yang, and Tung-Yuan Ho. Nickel superoxide dismutase protects nitrogen fixation in Trichodesmium. Limnology and Oceanography Letters, 7(4):363–371, August 2022. CODEN LOLIBT. ISSN 2378-2242.

Conine:2018:ASN

[CRP^{+}18] Andrea L. Conine, Daniel C. Rearick, Michael J. Paterson, Marguerite A. Xenopoulos, and Paul C. Frost. Addition of silver nanoparticles has no long-term effects on natural phytoplankton community dynamics in a boreal lake. Limnology and Oceanography Letters, 3(4):311–319, August 2018. CODEN LOLIBT. ISSN 2378-2242.

Cael:2022:TRB

[CS22] Brendan B. Cael and David Seekell. A theory for the relationship between lake surface area and maximum depth. Limnology and Oceanography Letters, 7(6):527–533, December 2022. CODEN ???? ISSN 2378-2242.

Caldwell:2020:FSI

[CSC20] Aliya Caldwell, Jennifer Seavey, and Elizabeth Craig. Foraging strategy impacts plastic ingestion risk in seabirds. Limnology and Oceanography Letters, 5(1):163–168, February 2020. CO-DEN LOLIBT. ISSN 2378-2242.

Call:2018:RTP

[CSEP^{+}18] Mitchell Call, Christian J. Sanders, Alex Enrich-Prast, Luciana Sanders, Humberto Marotta, Isaac R. Santos, and Damien T. Maher. Radon-traced pore-water as a potential source of CO_{2} and CH_{4} to receding black and clear water environments in the Amazon Basin. Limnology and Oceanography Letters, 3(5):375– 383, October 2018. CODEN LOLIBT. ISSN 2378-2242.

Chen:2022:PWE

[CSH^{+}22] Xiaogang Chen, Isaac R. Santos, Duofei Hu, Lucheng Zhan, Yan Zhang, Ze Zhao, Shengjie Hu, and Ling Li. Pore-water exchange flushes blue carbon from intertidal saltmarsh sediments into the sea. Limnology and Oceanography Letters, 7(4):312–320, August 2022. CODEN LOLIBT. ISSN 2378-2242.

Cheruvelil:2021:LUL

[CSM^{+}21] Kendra Spence Cheruvelil, Patricia A. Soranno, Ian M. Mc-Cullough, Katherine E. Webster, Lauren K. Rodriguez, and Nicole J. Smith. LAGOS-US LOCUS v1.0: Data module of location, identifiers, and physical characteristics of lakes and their watersheds in the conterminous U.S. Limnology and Oceanography Letters, 6(5):270–292, October 2021. CODEN LOLIBT. ISSN 2378-2242.

Coleman:2022:RMM

[CST^{+}22] Daniel J. Coleman, Mark Schuerch, Stijn Temmerman, Glenn Guntenspergen, Christopher G. Smith, and Matthew L. Kirwan. Reconciling models and measurements of marsh vulnerability to sea level rise. Limnology and Oceanography Letters, 7(2):140– 149, April 2022. CODEN LOLIBT. ISSN 2378-2242.

Cline:2020:LSR

[CSW^{+}20] Timothy J. Cline, Daniel E. Schindler, Timothy E. Walsworth, David W. French, and Peter J. Lisi. Low snowpack reduces thermal response diversity among streams across a landscape. Limnology and Oceanography Letters, 5(3):254–263, June 2020. CODEN LOLIBT. ISSN 2378-2242.

Cyronak:2020:DTP

[CTC^{+}20] Tyler Cyronak, Yuichiro Takeshita, Travis A. Courtney, Eric H. DeCarlo, Bradley D. Eyre, David I. Kline, Todd Martz, Heather Page, Nichole N. Price, Jennifer Smith, Laura Stoltenberg, Martin Tresguerres, and Andreas J. Andersson. Diel temperature and pH variability scale with depth across diverse coral reef habitats. Limnology and Oceanography Letters, 5(2):193–203, April 2020. CODEN LOLIBT. ISSN 2378-2242.

Deemer:2023:HCR

[DAS^{+}23] Bridget R. Deemer, Caitlin M. Andrews, Kristin E. Strock, Nicholas Voichick, James Hensleigh, John R. Beaver, and Robert Radtke. Over half a century record of limnology data from Lake Powell, desert southwest United States: From reservoir filling to present day (1964–2021). Limnology and Oceanography Letters, 8(4):580–594, August 2023. CODEN ???? ISSN 2378-2242.

DelSontro:2018:GGE

[DBD18] Tonya DelSontro, Jake J. Beaulieu, and John A. Downing. Greenhouse gas emissions from lakes and impoundments: Upscaling in the face of global change. Limnology and Oceanography Letters, 3(3):64–75, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Denfeld:2018:SCD

[DBdG^{+}18] Blaize A. Denfeld, Helen M. Baulch, Paul A. del Giorgio, Stephanie E. Hampton, and Jan Karlsson. A synthesis of carbon dioxide and methane dynamics during the ice-covered period of northern lakes. Limnology and Oceanography Letters, 3(3):117– 131, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Dahl:2021:ESO

[DBG21] Martin Dahl, Mats Björk, and Martin Gullström. Effects of seagrass overgrazing on sediment erosion and carbon sink capacity: Current understanding and future priorities. Limnology and Oceanography Letters, 6(6):309–319, December 2021. CO-DEN LOLIBT. ISSN 2378-2242.

Dexter:2020:NIZ

[DBRB20] Eric Dexter, Stephen M. Bollens, and Gretchen Rollwagen-Bollens. Native and invasive zooplankton show differing responses to decadal-scale increases in maximum temperatures in a large temperate river. Limnology and Oceanography Letters, 5(6):403–409, December 2020. CODEN LOLIBT. ISSN 2378-2242.

Dovidat:2020:PPA

[DBVB20] Lena C. Dovidat, Bregje W. Brinkmann, Martina G. Vijver, and Thijs Bosker. Plastic particles adsorb to the roots of freshwater vascular plant Spirodela polyrhiza but do not impair growth. Limnology and Oceanography Letters, 5(1):37–45, February 2020. CODEN LOLIBT. ISSN 2378-2242.

delCampo:2021:FIA

[dCCS21] Rubén del Campo, Roland Corti, and Gabriel Singer. Flow intermittence alters carbon processing in rivers through chemical diversification of leaf litter. Limnology and Oceanography Letters, 6(5):232–242, October 2021. CODEN LOLIBT. ISSN 2378-2242.

Dudek:2020:MCM

[DCPN20] Kassandra L. Dudek, Bianca N. Cruz, Beth Polidoro, and Susanne Neuer. Microbial colonization of microplastics in the Caribbean Sea. Limnology and Oceanography Letters, 5(1):5– 17, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Dextro:2023:EAI

[DDF^{+}23] Rafael B. Dextro, Endrews Delbaje, Paloma N. N. Freitas, Vanessa Geraldes, Ernani Pinto, Paul F. Long, and Marli F. Fiore. Environmental adaptations by the intertidal Antarctic cyanobacterium Halotia branconii CENA392 as revealed using long-read genome sequencing. Limnology and Oceanography Letters, 8(5):770–780, October 2023. CODEN ???? ISSN 2378-2242.

Duffy:2018:AWC

[DDH18] Christopher J. Duffy, Hilary A. Dugan, and Paul C. Hanson. The age of water and carbon in lake-catchments: a simple dynamical model. Limnology and Oceanography Letters, 3(3):236– 245, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Doubek:2020:SWS

[DGS20] Jonathan P. Doubek, Sadye K. Goldfarb, and Jason D. Stockwell. Should we be sampling zooplankton at night? Limnology and Oceanography Letters, 5(4):313–321, August 2020. CODEN LOLIBT. ISSN 2378-2242.

DAmbrosio:2021:MEC

[DH21] Sofia L. D’Ambrosio and John A. Harrison. Methanogenesis exceeds CH_{4} consumption in eutrophic lake sediments. Limnology and Oceanography Letters, 6(4):173–181, August 2021. CODEN LOLIBT. ISSN 2378-2242.

Dugan:2017:IFR

[DHM17] Hilary A. Dugan, Greta Helmueller, and John J. Magnuson. Ice formation and the risk of chloride toxicity in shallow wetlands and lakes. Limnology and Oceanography Letters, 2(5):150–158, October 2017. CODEN LOLIBT. ISSN 2378-2242.

Dallas:2016:OOI

[DKZD16] Tad A. Dallas, Andrew M. Kramer, Marcus Zokan, and John M. Drake. Ordination obscures the influence of environment on plankton metacommunity structure. Limnology and Oceanography Letters, 1(1):54–61, December 2016. CODEN LOLIBT. ISSN 2378-2242.

Davis:2019:DVA

[DLT^{+}19] C. Davis, M. C. Lohan, R. Tuerena, E. Cerdan-Garcia, E. M. S. Woodward, A. Tagliabue, and C. Mahaffey. Diurnal variability in alkaline phosphatase activity and the potential role of zooplankton. Limnology and Oceanography Letters, 4(3):71–78, June 2019. CODEN LOLIBT. ISSN 2378-2242.

delaVega:2021:AST

[dlVMT^{+}21] Camille de la Vega, Claire Mahaffey, Robyn E. Tuerena, David J. Yurkowski, Steven H. Ferguson, Garry B. Stenson, Erling S. Nordøy, Tore Haug, Martin Biuw, Sophie Smout, Jo Hopkins, Alessandro Tagliabue, and Rachel M. Jeffreys. Arctic seals as tracers of environmental and ecological change. Limnology and Oceanography Letters, 6(1):24–32, February 2021. CODEN LOLIBT. ISSN 2378-2242.

Dugan:2023:SSS

[DR23] Hilary A. Dugan and Linnea A. Rock. The slow and steady salinization of Sparkling Lake, Wisconsin. Limnology and Oceanography Letters, 8(1):74–82, February 2023. CODEN ???? ISSN 2378-2242.

Dugan:2023:TCL

[DRKM23] Hilary A. Dugan, Linnea A. Rock, Anthony D. Kendall, and Robert J. Mooney. Tributary chloride loading into Lake Michigan. Limnology and Oceanography Letters, 8(1):83–92, February 2023. CODEN ???? ISSN 2378-2242.

Drake:2018:TCI

[DRS18] Travis W. Drake, Peter A. Raymond, and Robert G. M. Spencer. Terrestrial carbon inputs to inland waters: a current synthesis of estimates and uncertainty. Limnology and Oceanography Letters, 3(3):132–142, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Dulai:2023:RNM

[DSA^{+}23] Henrietta Dulai, Celia M. Smith, Daniel W. Amato, Veronica Gibson, and Leah L. Bremer. Risk to native marine macroalgae from land-use and climate change-related modifications to groundwater discharge in Hawai’i. Limnology and Oceanography Letters, 8(1):141–153, February 2023. CODEN ???? ISSN 2378-2242.

Dodds:2017:RPD

[DTSC17] Walter K. Dodds, Flavia Tromboni, Wesley Aparecido Saltarelli, and Davi Gasparini Fernandes Cunha. The root of the problem: Direct influence of riparian vegetation on estimation of stream ecosystem metabolic rates. Limnology and Oceanography Letters, 2(1):9–17, February 2017. CODEN LOLIBT. ISSN 2378-2242.

deWit:2018:PCC

[dWCJB^{+}18] Heleen A. de Wit, Raoul-Marie Couture, Leah Jackson-Blake, Martyn N. Futter, Salar Valinia, Kari Austnes, Jose-Luis Guerrero, and Yan Lin. Pipes or chimneys? For carbon cycling in small boreal lakes, precipitation matters most. Limnology and Oceanography Letters, 3(3):275–284, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Fork:2018:EHC

[FBDH18] Megan L. Fork, Joanna R. Blaszczak, Joseph M. Delesantro, and James B. Heffernan. Engineered headwaters can act as sources of dissolved organic matter and nitrogen to urban stream networks. Limnology and Oceanography Letters, 3(3):215–224, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Frenken:2020:WAV

[FBV^{+}20] Thijs Frenken, Corina P. D. Brussaard, Mandy Velthuis, Ralf Aben, Garabet Kazanjian, Sabine Hilt, Sarian Kosten, Edwin T. H. M. Peeters, Lisette N. de Senerpont Domis, Susanne Stephan, Ellen van Donk, and Dedmer B. Van de Waal. Warming advances virus population dynamics in a temperate freshwater plankton community. Limnology and Oceanography Letters, 5 (4):295–304, August 2020. CODEN LOLIBT. ISSN 2378-2242.

Falkenberg:2024:HWL

[FJS24] Laura J. Falkenberg, Patrick W. S. Joyce, and Patricia A. Soranno. How to write lay summaries of research articles for wider accessibility. Limnology and Oceanography Letters, 9(2):93–98, April 2024. CODEN ???? ISSN 2378-2242.

Fox:2022:ABA

[FKG^{+}22] James Fox, Sasha J. Kramer, Jason R. Graff, Michael J. Behrenfeld, Emmanuel Boss, Gavin Tilstone, and Kimberly H. Halsey. An absorption-based approach to improved estimates of phytoplankton biomass and net primary production. Limnology and Oceanography Letters, 7(5):419–426, October 2022. CO-DEN ???? ISSN 2378-2242.

Fork:2020:DOM

[FKS20] Megan L. Fork, Jan Karlsson, and Ryan A. Sponseller. Dissolved organic matter regulates nutrient limitation and growth of benthic algae in northern lakes through interacting effects on nutrient and light availability. Limnology and Oceanography Letters, 5(6):417–424, December 2020. CODEN LOLIBT. ISSN 2378-2242.

Ferguson:2020:BTR

[FOE20] Angus J. P. Ferguson, Joanne Oakes, and Bradley D. Eyre. Bottom trawling reduces benthic denitrification and has the potential to influence the global nitrogen cycle. Limnology and Oceanography Letters, 5(3):237–245, June 2020. CODEN LOLIBT. ISSN 2378-2242.

Frischkorn:2018:TMC

[FRVD18] Kyle R. Frischkorn, Mónica Rouco, Benjamin A. S. Van Mooy, and Sonya T. Dyhrman. The Trichodesmium microbiome can modulate host N_{2} fixation. Limnology and Oceanography Letters, 3(6):401–408, December 2018. CODEN LOLIBT. ISSN 2378-2242.

Falkenberg:2021:OWR

[FSA21] Laura J. Falkenberg, Dina-Leigh Simons, and Kathryn M. Anderson. Ocean warming reduces gastropod survival despite maintenance of feeding and oxygen consumption rates. Limnology and Oceanography Letters, 6(4):165–172, August 2021. CODEN LOLIBT. ISSN 2378-2242.

Frainer:2022:SSE

[FT22] André Frainer and Scott D. Tiegs. A stream-to-sea experiment reveals inhibitory effects of freshwater residency on organicmatter decomposition in the sea. Limnology and Oceanography Letters, 7(3):202–209, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Greco:2023:ECN

[GAFS23] Danielle A. Greco, Shelley E. Arnott, Isabelle B. Fournier, and Brandon S. Schamp. Effects of chloride and nutrients on freshwater plankton communities. Limnology and Oceanography Letters, 8(1):48–55, February 2023. CODEN ???? ISSN 2378-2242.

Granek:2020:MAO

[GBH20] E. F. Granek, S. M. Brander, and E. B. Holland. Microplastics in aquatic organisms: Improving understanding and identifying research directions for the next decade. Limnology and Oceanography Letters, 5(1):1–4, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Gradoville:2022:ECR

[GD22] Mary R. Gradoville and Bridget R. Deemer. Early career researchers have questions about peer review — we asked the ASLO editors for answers. Limnology and Oceanography Letters, 7(3):185–188, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Gardner:2024:NMS

[GDEK24] John R. Gardner, Martin W. Doyle, Scott H. Ensign, and David M. Kahler. A new metric for sunlight exposure in rivers, lakes, and oceans. Limnology and Oceanography Letters, 9(2): 128–134, April 2024. CODEN ???? ISSN 2378-2242.

Gali:2023:GOD

[GDP^{+}23] Martí Galí, Emmanuel Devred, Gonzalo L. Pérez, David J. Kieber, and Rafel Simó. Global Ocean dimethylsulfide photolysis rates quantified with a spectrally and vertically resolved model. Limnology and Oceanography Letters, 8(5):760–769, October 2023. CODEN ???? ISSN 2378-2242.

Gafar:2019:PIO

[GES19] N. A. Gafar, B. D. Eyre, and K. G. Schulz. Particulate inorganic to organic carbon production as a predictor for coccolithophorid sensitivity to ongoing ocean acidification. Limnology and Oceanography Letters, 4(3):62–70, June 2019. CODEN LOLIBT. ISSN 2378-2242.

Guy-Haim:2017:NMB

[GHHKY^{+}17] Tamar Guy-Haim, Orit Hyams-Kaphzan, Erez Yeruham, Ahuva Almogi-Labin, and James T. Carlton. A novel marine bioinvasion vector: Ichthyochory, live passage through fish. Limnology and Oceanography Letters, 2(3):81–90, June 2017. CODEN LOLIBT. ISSN 2378-2242.

Gronning:2022:GIA

[GK22] Josephine Grønning and Thomas Kiørboe. Grazer-induced aggregation in diatoms. Limnology and Oceanography Letters, 7 (6):492–500, December 2022. CODEN ???? ISSN 2378-2242.

Gerhard:2023:EVA

[GKR^{+}23] Miriam Gerhard, Apostolos-Manuel Koussoroplis, Michael Raatz, Christian Pansch, Samuel B. Fey, Jahangir Vajedsamiei, Maria Calderó-Pascual, David Cunillera-Montcusí, Noël P. D. Juvigny-Khenafou, Francesco Polazzo, Patrick K. Thomas, Celia C. Symons, Meryem Beklioğlu, Stella A. Berger, Rosa M. Chefaoui, Kemal Ali Ger, Silke Langenheder, Jens C. Nejstgaard, Robert Ptacnik, and Maren Striebel. Environmental variability in aquatic ecosystems: Avenues for future multifactorial experiments. Limnology and Oceanography Letters, 8(2): 247–266, April 2023. CODEN ???? ISSN 2378-2242.

Gudasz:2017:CTO

[GRK^{+}17] Cristian Gudasz, Marc Ruppenthal, Karsten Kalbitz, Chiara Cerli, Sabine Fiedler, Yvonne Oelmann, August Andersson, and Jan Karlsson. Contributions of terrestrial organic carbon to northern lake sediments. Limnology and Oceanography Letters, 2(6):218–227, December 2017. CODEN LOLIBT. ISSN 2378-2242.

Gutierrez-Rodriguez:2023:DGG

[GRLS^{+}23] Andres Gutiérrez-Rodríguez, Mikel Latasa, Karl Safi, Matthew H. Pinkerton, and Scott D. Nodder. Decoupled growth and grazing rates of diatoms and green algae drive increased phytoplankton productivity on HNLC sub-Antarctic plateaux. Limnology and Oceanography Letters, 8(6):896–905, December 2023. CODEN ???? ISSN 2378-2242.

Guo:2023:BCF

[GYB^{+}23] Xiaoyi Guo, Huamao Yuan, William Burnett, Han Zhang, Ergang Lian, Haowei Xu, Zhigang Yu, and Bochao Xu. Ba/ Ca in foraminifera shells as a proxy of submarine groundwater discharge. Limnology and Oceanography Letters, 8(3):490–497, June 2023. CODEN ???? ISSN 2378-2242.

Heine:2019:CRI

[HAWH19] Kyle B. Heine, Ash Abebe, Alan E. Wilson, and Wendy R. Hood. Copepod respiration increases by 7% per ^{o}C increase in temperature: a meta-analysis. Limnology and Oceanography Letters, 4(3):53–61, June 2019. CODEN LOLIBT. ISSN 2378-2242.

Hogle:2016:CMB

[HBB^{+}16] Shane L. Hogle, Randelle M. Bundy, Jessica M. Blanton, Eric E. Allen, and Katherine A. Barbeau. Copiotrophic marine bacteria are associated with strong iron-binding ligand production during phytoplankton blooms. Limnology and Oceanography Letters, 1 (1):36–43, December 2016. CODEN LOLIBT. ISSN 2378-2242.

Hampton:2024:WIC

[HBL^{+}24] Stephanie E. Hampton, Jill S. Baron, Robert Ladwig, Ryan P. McClure, Michael F. Meyer, Isabella A. Oleksy, and Anna Shampain. Warming-induced changes in benthic redox as a potential driver of increasing benthic algal blooms in high-elevation lakes. Limnology and Oceanography Letters, 9(1):1–6, February 2024. CODEN ???? ISSN 2378-2242.

Haque:2022:SSD

[HBN^{+}22] Md. Morshedul Haque, Most Shirina Begum, Omme K. Nayna, Shafi M. Tareq, and Ji-Hyung Park. Seasonal shifts in diurnal variations of pCO_{2} and O_{2} in the lower Ganges River. Limnology and Oceanography Letters, 7(3):191–201, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Harris:2020:IMV

[HC20] Lyda S. T. Harris and Emily Carrington. Impacts of microplastic vs. natural abiotic particles on the clearance rate of a marine mussel. Limnology and Oceanography Letters, 5(1):66–73, February 2020. CODEN LOLIBT. ISSN 2378-2242.

He:2023:MSI

[HCA^{+}23] Xiaoshuai He, Yi-Sheng Chen, Shin-Jing Ang, Jen-Chieh Shiao, Chun-Mao Tseng, and John R. Reinfelder. Mercury stable isotopes reveal sources of methylmercury and prey in giant Pacific bluefin tuna from the western North Pacific Ocean. Limnology and Oceanography Letters, 8(3):481–489, June 2023. CODEN ???? ISSN 2378-2242.

Hayes:2017:KDB

[HDC^{+}17] Nicole M. Hayes, Bridget R. Deemer, Jessica R. Corman, N. Roxanna Razavi, and Kristin E. Strock. Key differences between lakes and reservoirs modify climate signals: a case for a new conceptual model. Limnology and Oceanography Letters, 2 (2):47–62, April 2017. CODEN LOLIBT. ISSN 2378-2242.

Hotaling:2023:TSA

[HDPEF23] Scott Hotaling, Bridget R. Deemer, Kelsey Poulson-Ellestad, and Laura J. Falkenberg. Taking steps to address inequities in open-access publishing through an early career publication honor. Limnology and Oceanography Letters, 8(3):385–387, June 2023. CODEN ???? ISSN 2378-2242.

Hetherington:2022:ISM

[HDSH^{+}22] Elizabeth D. Hetherington, Alejandro Damian-Serrano, Steven H. D. Haddock, Casey W. Dunn, and C. Anela Choy. Integrating siphonophores into marine food-web ecology. Limnology and Oceanography Letters, 7(2):81–95, April 2022. CODEN LOLIBT. ISSN 2378-2242.

Heine:2023:SCM

[Hei23] Kyle B. Heine. Some considerations of measuring temperature sensitivity in thermal ecology. Limnology and Oceanography Letters, 8(4):539–545, August 2023. CODEN ???? ISSN 2378-2242.

Henderikx-Freitas:2021:CGR

[HFKBW21] Fernanda Henderikx-Freitas, David M. Karl, Karin M. Björkman, and Angelicque E. White. Constraining growth rates and the ratio of living to nonliving particulate carbon using beam attenuation and adenosine-5’-triphosphate at Station ALOHA. Limnology and Oceanography Letters, 6(5):243–252, October 2021. CODEN LOLIBT. ISSN 2378-2242.

Horn:2020:EER

[HGS20] Dorothy A. Horn, Elise F. Granek, and Clare L. Steele. Effects of environmentally relevant concentrations of microplastic fibers on Pacific mole crab (Emerita analoga) mortality and reproduction. Limnology and Oceanography Letters, 5(1):74–83, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Hayes:2020:ELW

[HHSL20] Nicole M. Hayes, Heather A. Haig, Gavin L. Simpson, and Peter R. Leavitt. Effects of lake warming on the seasonal risk of toxic cyanobacteria exposure. Limnology and Oceanography Letters, 5(6):393–402, December 2020. CODEN LOLIBT. ISSN 2378-2242.

Honda:2023:SVP

[HJS23] Isabel A. Honda, Rubao Ji, and Andrew R. Solow. Spatially varying plankton synchrony patterns at seasonal and interannual scales in a well-connected shelf sea. Limnology and Oceanography Letters, 8(6):906–915, December 2023. CODEN ???? ISSN 2378-2242.

Hedberg:2023:VFA

[HLAW23] Per Hedberg, Danny C. P. Lau, Séréna Albert, and Monika Winder. Variation in fatty acid content among benthic invertebrates in a seasonally driven system. Limnology and Oceanography Letters, 8(5):751–759, October 2023. CODEN ???? ISSN 2378-2242.

Hudson:2023:TRG

[HLH23] Danielle T. Hudson, Jason A. Leach, and Daniel Houle. Thermal regimes of groundwater- and lake-fed headwater streams differ in their response to climate variability. Limnology and Oceanography Letters, 8(6):885–895, December 2023. CODEN ???? ISSN 2378-2242.

Heal:2022:DAC

[HMIB22] Katherine R. Heal, Ashley E. Maloney, Anitra E. Ingalls, and Randelle M. Bundy. Diverse arsenic-containing lipids in the surface ocean. Limnology and Oceanography Letters, 7(1):43– 51, February 2022. CODEN LOLIBT. ISSN 2378-2242.

Hounshell:2021:WEO

[HMLC21] Alexandria G. Hounshell, Ryan P. McClure, Mary E. Lofton, and Cayelan C. Carey. Whole-ecosystem oxygenation experiments reveal substantially greater hypolimnetic methane concentrations in reservoirs during anoxia. Limnology and Oceanography Letters, 6(1):33–42, February 2021. CODEN LOLIBT. ISSN 2378-2242.

Hurt:2020:MPT

[HOP20] Raven Hurt, Catherine M. O’Reilly, and William L. Perry. Microplastic prevalence in two fish species in two U.S. reservoirs. Limnology and Oceanography Letters, 5(1):147–153, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Hoppe:2022:ARP

[Hop22] Clara J. M. Hoppe. Always ready? Primary production of Arctic phytoplankton at the end of the polar night. Limnology and Oceanography Letters, 7(2):167–174, April 2022. CODEN LOLIBT. ISSN 2378-2242.

Hotaling:2020:SRC

[Hot20] Scott Hotaling. Simple rules for concise scientific writing. Limnology and Oceanography Letters, 5(6):379–383, December 2020. CODEN LOLIBT. ISSN 2378-2242.

Hawkes:2018:EIC

[HPS^{+}18] Jeffrey A. Hawkes, Claudia Patriarca, Per J. R. Sjöberg, Lars J. Tranvik, and Jonas Bergquist. Extreme isomeric complexity of dissolved organic matter found across aquatic environments. Limnology and Oceanography Letters, 3(2):21–30, April 2018. CODEN LOLIBT. ISSN 2378-2242.

Holgerson:2024:HRC

[HRR24] Meredith A. Holgerson, Nicholas E. Ray, and Chelsea Russ. High rates of carbon burial linked to autochthonous production in artificial ponds. Limnology and Oceanography Letters, 9(1):43– 51, February 2024. CODEN ???? ISSN 2378-2242.

Hebert:2023:LSD

[HSCA^{+}23] Marie-Pier Hébert, Celia C. Symons, Miguel Cañedo-Argüelles, Shelley E. Arnott, Alison M. Derry, Vincent Fugère, William D. Hintz, Stephanie J. Melles, Louis Astorg, Henry K. Baker, Jennifer A. Brentrup, Amy L. Downing, Zeynep Ersoy, Carmen Espinosa, Jaclyn M. Franceschini, Angelina T. Giorgio, Norman Göbeler, Derek K. Gray, Danielle Greco, Emily Hassal, Mercedes Huynh, Samuel Hylander, Kacie L. Jonasen, Andrea Kirkwood, Silke Langenheder, Ola Langvall, Hjalmar Laudon, Lovisa Lind, Maria Lundgren, Alexandra McClymont, Lorenzo Proia, Rick A. Relyea, James A. Rusak, Matthew S. Schuler, Catherine L. Searle, Jonathan B. Shurin, Christopher F. Steiner, Maren Striebel, Simon Thibodeau, Pablo Urrutia Cordero, Lidia Vendrell-Puigmitja, Gesa A. Weyhenmeyer, and Beatrix E. Beisner. Lake salinization drives consistent losses of zooplankton abundance and diversity across coordinated mesocosm experiments. Limnology and Oceanography Letters, 8(1):19–29, February 2023. CODEN ???? ISSN 2378-2242.

Hotchkiss:2018:TMI

[HSH18] E. R. Hotchkiss, S. Sadro, and P. C. Hanson. Toward a more integrative perspective on carbon metabolism across lentic and lotic inland waters. Limnology and Oceanography Letters, 3(3): 57–63, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Isles:2023:WSP

[ICH^{+}23] Peter D. F. Isles, Irena F. Creed, Dag O. Hessen, Pirkko Kortelainen, Michael Paterson, Francesco Pomati, James A. Rusak, Jussi Vuorenmaa, and Ann-Kristin Bergström. Widespread synchrony in phosphorus concentrations in northern lakes linked to winter temperature and summer precipitation. Limnology and Oceanography Letters, 8(4):639–648, August 2023. CODEN ???? ISSN 2378-2242.

Inman:2022:HCM

[Inm22] Bryce G. Inman. A hidden cost of mucus production by phytoplankton: Viscosity hinders nutrient uptake. Limnology and Oceanography Letters, 7(3):261–268, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Jabre:2020:IEI

[JB20] Loay Jabre and Erin M. Bertrand. Interactive effects of iron and temperature on the growth of Fragilariopsis cylindrus. Limnology and Oceanography Letters, 5(5):363–370, October 2020. CODEN LOLIBT. ISSN 2378-2242.

Juska:2022:VSM

[JB22] Ieva Juska and Peter Berg. Variation in seagrass meadow respiration measured by aquatic eddy covariance. Limnology and Oceanography Letters, 7(5):410–418, October 2022. CODEN ???? ISSN 2378-2242.

Jameson:2021:CMS

[JBG^{+}21] Brett D. Jameson, Peter Berg, Damian S. Grundle, Catherine J. Stevens, and S. Kim Juniper. Continental margin sediments underlying the NE Pacific oxygen minimum zone are a source of nitrous oxide to the water column. Limnology and Oceanography Letters, 6(2):68–76, April 2021. CODEN LOLIBT. ISSN 2378-2242.

James:2020:MMA

[JEN^{+}20] Anna K. James, Chance J. English, Nicholas J. Nidzieko, Craig A. Carlson, and Elizabeth G. Wilbanks. Macroalgal microbiome altered in the presence of epiphytes. Limnology and Oceanography Letters, 5(5):354–362, October 2020. CODEN LOLIBT. ISSN 2378-2242.

Joppien:2022:MAF

[JWSD22] Marlena Joppien, Hildegard Westphal, Marleen Stuhr, and Steve S. Doo. Microplastics alter feeding strategies of a coral reef organism. Limnology and Oceanography Letters, 7(2):131–139, April 2022. CODEN LOLIBT. ISSN 2378-2242.

Jones:2018:HSC

[JZKS18] Stuart E. Jones, Jacob A. Zwart, Patrick T. Kelly, and Christopher T. Solomon. Hydrologic setting constrains lake heterotrophy and terrestrial carbon fate. Limnology and Oceanography Letters, 3(3):256–264, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Kincaid:2021:WCH

[KBB^{+}21] Dustin W. Kincaid, Whitney S. Beck, Jessica E. Brandt, Margaret Mars Brisbin, Kaitlin J. Farrell, Kelly L. Hondula, Erin I. Larson, and Arial J. Shogren. Wikipedia can help resolve information inequality in the aquatic sciences. Limnology and Oceanography Letters, 6(1):18–23, February 2021. CODEN LOLIBT. ISSN 2378-2242.

Kinsman-Costello:2023:MCE

[KCBG^{+}23] Lauren Kinsman-Costello, Eban Bean, Audrey Goeckner, Jeffrey W. Matthews, Michael O’Driscoll, Monica M. Palta, Ariane L. Peralta, Alexander J. Reisinger, Gabriela J. Reyes, Ashley R. Smyth, and Marie Stofan. Mud in the city: Effects of freshwater salinization on inland urban wetland nitrogen and phosphorus availability and export. Limnology and Oceanography Letters, 8(1):112–130, February 2023. CODEN ???? ISSN 2378-2242.

Kelly:2021:HPS

[KCPGE21] Katie M. Kelly, Natasha L. Carlson-Perret, Thomas D. Glaze, and Dirk V. Erler. Hiding in plain sight: the secret contribution of the solitary ascidian Herdmania grandis to temperate reef nitrous oxide production. Limnology and Oceanography Letters, 6(4):192–199, August 2021. CODEN LOLIBT. ISSN 2378-2242.

Kefford:2023:SSA

[KHB^{+}23] Ben J. Kefford, Ross V. Hyne, Andrew J. Brooks, Jonathan P. Bray, Mark Shenton, Kasey Hills, and Susan J. Nichols. Singlespecies acute lethal toxicity tests are not predictive of relative population and community effects of two salinity types. Limnology and Oceanography Letters, 8(1):181–189, February 2023. CODEN ???? ISSN 2378-2242.

Koenig:2019:EPR

[KHS^{+}19] Lauren E. Koenig, Ashley M. Helton, Philip Savoy, Enrico Bertuzzo, James B. Heffernan, Robert O. Hall, Jr., and Emily S. Bernhardt. Emergent productivity regimes of river networks. Limnology and Oceanography Letters, 4(5):173–181, October 2019. CODEN LOLIBT. ISSN 2378-2242.

Kohler:2021:HAT

[KHS^{+}21] Tyler J. Kohler, Adrian Howkins, Eric R. Sokol, Katerina Kopalová, Aneliya Cox, Joshua P. Darling, Michael N. Gooseff, and Diane M. McKnight. From the Heroic Age to today: What diatoms from Shackleton’s Nimrod expedition can tell us about the ecological trajectory of Antarctic ponds. Limnology and Oceanography Letters, 6(6):379–387, December 2021. CO-DEN LOLIBT. ISSN 2378-2242.

Kujawinski:2017:PAR

[KLA^{+}17] Elizabeth B. Kujawinski, Krista Longnecker, Harriet Alexander, Sonya T. Dyhrman, Cara L. Fiore, Sheean T. Haley, and Winifred M. Johnson. Phosphorus availability regulates intracellular nucleotides in marine eukaryotic phytoplankton. Limnology and Oceanography Letters, 2(4):119–129, August 2017. CODEN LOLIBT. ISSN 2378-2242.

Kaushal:2023:FSF

[KML^{+}23] Sujay S. Kaushal, Paul M. Mayer, Gene E. Likens, Jenna E. Reimer, Carly M. Maas, Megan A. Rippy, Stanley B. Grant, Ian Hart, Ryan M. Utz, Ruth R. Shatkay, Barret M. Wessel, Christine E. Maietta, Michael L. Pace, Shuiwang Duan, Walter L. Boger, Alexis M. Yaculak, Joseph G. Galella, Kelsey L. Wood, Carol J. Morel, William Nguyen, Shane Elizabeth C. Querubin, Rebecca A. Sukert, Anna Lowien, Alyssa Wellman Houde, Anaïs Roussel, Andrew J. Houston, Ari Cacopardo, Cristy Ho, Haley Talbot-Wendlandt, Jacob M. Widmer, Jairus Slagle, James A. Bader, Jeng Hann Chong, Jenna Wollney, Jordan Kim, Lauren Shepherd, Matthew T. Wilfong, Megan Houlihan, Nathan Sedghi, Rebecca Butcher, Sona Chaudhary, and William D. Becker. Five state factors control progressive stages of freshwater salinization syndrome. Limnology and Oceanography Letters, 8(1):190–211, February 2023. CODEN ???? ISSN 2378-2242.

Keppeler:2021:BST

[KOLD^{+}21] Friedrich W. Keppeler, Jill A. Olin, Paola C. López-Duarte, Michael J. Polito, Linda M. Hooper-Bùi, Sabrina S. Taylor, Nancy N. Rabalais, F. Joel Fodrie, Brian J. Roberts, R. Eugene Turner, Charles W. Martin, and Olaf P. Jensen. Body size, trophic position, and the coupling of different energy pathways across a saltmarsh landscape. Limnology and Oceanography Letters, 6(6):360–368, December 2021. CODEN LOLIBT. ISSN 2378-2242.

Kritzberg:2017:CLT

[Kri17] Emma S. Kritzberg. Centennial-long trends of lake browning show major effect of afforestation. Limnology and Oceanography Letters, 2(4):105–112, August 2017. CODEN LOLIBT. ISSN 2378-2242.

Knoll:2019:CLR

[KSD^{+}19] Lesley B. Knoll, Sapna Sharma, Blaize A. Denfeld, Giovanna Flaim, Yukari Hori, John J. Magnuson, Dietmar Straile, and Gesa A. Weyhenmeyer. Consequences of lake and river ice loss on cultural ecosystem services. Limnology and Oceanography Letters, 4(5):119–131, October 2019. CODEN LOLIBT. ISSN 2378-2242.

Kohler:2018:CRH

[KSL^{+}18] Tyler J. Kohler, Lee F. Stanish, Daniel Liptzin, John E. Barrett, and Diane M. McKnight. Catch and release: Hyporheic retention and mineralization of N-fixing Nostoc sustains downstream microbial mat biomass in two polar desert streams. Limnology and Oceanography Letters, 3(4):357–364, August 2018. CODEN LOLIBT. ISSN 2378-2242.

Karlsson:2024:IMP

[KVS^{+}24] J. Karlsson, H. A. Verheijen, D. A. Seekell, D. Vachon, and M. Klaus. Ice-melt period dominates annual carbon dioxide evasion from clear-water Arctic lakes. Limnology and Oceanography Letters, 9(2):112–118, April 2024. CODEN ???? ISSN 2378-2242.

King:2021:LNC

[KWRC21] Katelyn B. S. King, Qi Wang, Lauren K. Rodriguez, and Kendra S. Cheruvelil. Lake networks and connectivity metrics for the conterminous U.S. (LAGOS-US Networks v1). Limnology and Oceanography Letters, 6(5):293–307, October 2021. CO-DEN LOLIBT. ISSN 2378-2242.

Lopez:2024:ELT

[LBB^{+}24] Lianna S. Lopez, Aman Basu, Kevin Blagrave, Gerald Bove, Kenton Stewart, Dawn Bazely, and Sapna Sharma. Establishing a long-term citizen science project? Lessons learned from the Community Lake Ice Collaboration spanning over 30 yr and 1000 lakes. Limnology and Oceanography Letters, 9(2):99–111, April 2024. CODEN ???? ISSN 2378-2242.

Liu:2022:EPT

[LCL22] Kailin Liu, Bingzhang Chen, and Hongbin Liu. Evidence of partial thermal compensation in natural phytoplankton assemblages. Limnology and Oceanography Letters, 7(2):122–130, April 2022. CODEN LOLIBT. ISSN 2378-2242.

Li:2024:RIB

[LDWE24] Qian Li, Kaiyi Dong, Ying Wang, and Kyle F. Edwards. Relative importance of bacterivorous mixotrophs in an estuary-coast environment. Limnology and Oceanography Letters, 9(1):81–91, February 2024. CODEN ???? ISSN 2378-2242.

Latour:2024:CSO

[LEvdM^{+}24] Pauline Latour, Sam Eggins, Pier van der Merwe, Lennart T. Bach, Philip W. Boyd, Michael J. Ellwood, Andrew R. Bowie, Kathrin Wuttig, and Robert F. Strzepek. Characterization of a Southern Ocean deep chlorophyll maximum: Response of phytoplankton to light, iron, and manganese enrichment. Limnology and Oceanography Letters, 9(2):145–154, April 2024. CODEN ???? ISSN 2378-2242.

Lieber:2023:SFB

[LFH^{+}23] Lilian Lieber, Cynthia Füchtencordsjürgen, Rebecca L. Hilder, Paula J. Revering, Ina Siekmann, Roland Langrock, and W. Alex M. Nimmo-Smith. Selective foraging behavior of seabirds in small-scale slicks. Limnology and Oceanography Letters, 8(2):286–294, April 2023. CODEN ???? ISSN 2378-2242.

Lambert:2019:MDB

[LFS19] Bennett S. Lambert, Vicente I. Fernandez, and Roman Stocker. Motility drives bacterial encounter with particles responsible for carbon export throughout the ocean. Limnology and Oceanography Letters, 4(5):113–118, October 2019. CODEN LOLIBT. ISSN 2378-2242.

Lupon:2023:GSC

[LGGF^{+}23] Anna Lupon, Lluís Gómez-Gener, Megan L. Fork, Hjalmar Laudon, Eugènia Martí, William Lidberg, and Ryan A. Sponseller. Groundwater-stream connections shape the spatial pattern and rates of aquatic metabolism. Limnology and Oceanography Letters, 8(2):350–358, April 2023. CODEN ???? ISSN 2378-2242.

LaBrie:2023:ADP

[LHL23] Richard LaBrie, Michael Hupfer, and Maximilian P. Lau. Anaerobic duration predicts biogeochemical consequences of oxygen depletion in lakes. Limnology and Oceanography Letters, 8(4): 666–674, August 2023. CODEN ???? ISSN 2378-2242.

Lapierre:2020:LBI

[LHMF20] Jean-François Lapierre, Adam J. Heathcote, Philippe Maisonneuve, and Christopher T. Filstrup. Is limnology becoming increasingly abiotic, riverine, and global? Limnology and Oceanography Letters, 5(2):204–211, April 2020. CODEN LOLIBT. ISSN 2378-2242.

Litchman:2023:UPH

[Lit23] Elena Litchman. Understanding and predicting harmful algal blooms in a changing climate: a trait-based framework. Limnology and Oceanography Letters, 8(2):229–246, April 2023. CO-DEN ???? ISSN 2378-2242.

Loewen:2024:SWL

[LJ24] Charlie J. G. Loewen and Donald A. Jackson. Salinization, warming, and loss of water clarity inhibit vertical mixing of small urban ponds. Limnology and Oceanography Letters, 9(2): 155–164, April 2024. CODEN ???? ISSN 2378-2242.

Lin:2024:ATY

[LK24] Tzu-Hao Lin and Shinsuke Kawagucci. Acoustic twilight: a year-long seafloor monitoring unveils phenological patterns in the abyssal soundscape. Limnology and Oceanography Letters, 9(1):23–32, February 2024. CODEN ???? ISSN 2378-2242.

Layden:2022:TAI

[LKB^{+}22] Tamara J. Layden, Colin T. Kremer, Delaney L. Brubaker, Maeve A. Kolk, Jessica V. Trout-Haney, David A. Vasseur, and Samuel B. Fey. Thermal acclimation influences the growth and toxin production of freshwater cyanobacteria. Limnology and Oceanography Letters, 7(1):34–42, February 2022. CODEN LOLIBT. ISSN 2378-2242.

Leach:2019:HLT

[LL19] Jason A. Leach and Hjalmar Laudon. Headwater lakes and their influence on downstream discharge. Limnology and Oceanography Letters, 4(4):105–112, August 2019. CODEN LOLIBT. ISSN 2378-2242.

Lukic:2022:TPP

[LLA^{+}22] Dunja Lukić, Romana Limberger, Sabine Agatha, David J. S. Montagnes, and Thomas Weisse. Thermal performance of planktonic ciliates differs between marine and freshwaters: a case study providing guidance for climate change studies. Limnology and Oceanography Letters, 7(6):520–526, December 2022. CO-DEN ???? ISSN 2378-2242.

Longnecker:2020:OSS

[LOS^{+}20] Krista Longnecker, Lisa Oswald, Melissa C. Kido Soule, Gregory A. Cutter, and Elizabeth B. Kujawinski. Organic sulfur: a spatially variable and understudied component of marine organic matter. Limnology and Oceanography Letters, 5(4):305– 312, August 2020. CODEN LOLIBT. ISSN 2378-2242.

Liu:2018:HCP

[LR18] Shaoda Liu and Peter A. Raymond. Hydrologic controls on pCO_{2} and CO_{2} efflux in US streams and rivers. Limnology and Oceanography Letters, 3(6):428–435, December 2018. CODEN LOLIBT. ISSN 2378-2242.

Lopes:2020:MID

[LRCG20] Clara Lopes, Joana Raimundo, Miguel Caetano, and Susana Garrido. Microplastic ingestion and diet composition of planktivorous fish. Limnology and Oceanography Letters, 5(1):103– 112, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Ladwig:2023:ISL

[LRD23] Robert Ladwig, Linnea A. Rock, and Hilary A. Dugan. Impact of salinization on lake stratification and spring mixing. Limnology and Oceanography Letters, 8(1):93–102, February 2023. CODEN ???? ISSN 2378-2242.

Lilly:2023:GRP

[LSER23] Laura E. Lilly, Iain M. Suthers, Jason D. Everett, and Anthony J. Richardson. A global review of pyrosomes: Shedding light on the ocean’s elusive gelatinous “fire-bodies”. Limnology and Oceanography Letters, 8(6):812–829, December 2023. CO-DEN ???? ISSN 2378-2242.

Landry:2022:LTS

[LSH^{+}22] Michael R. Landry, Karen E. Selph, Raleigh R. Hood, Claire H. Davies, and Lynnath E. Beckley. Low temperature sensitivity of picophytoplankton P : B ratios and growth rates across a natural 10^{o}C temperature gradient in the oligotrophic Indian Ocean. Limnology and Oceanography Letters, 7(2):112–121, April 2022. CODEN LOLIBT. ISSN 2378-2242.

Leiser:2022:MIR

[LWP22] Rico Leiser and Katrin Wendt-Potthoff. Microbial iron reduction does not release microplastics from organo-metallic aggregates. Limnology and Oceanography Letters, 7(3):244–250, June 2022. CODEN LOLIBT. ISSN 2378-2242.

McGowan:2018:VTD

[MAE^{+}18] Suzanne McGowan, N. John Anderson, Mary E. Edwards, Emma Hopla, Viv Jones, Pete G. Langdon, Antonia Law, Nadia Solovieva, Simon Turner, Maarten van Hardenbroek, Erika J. Whiteford, and Emma Wiik. Vegetation transitions drive the autotrophy–heterotrophy balance in Arctic lakes. Limnology and Oceanography Letters, 3(3):246–255, June 2018. CODEN LOLIBT. ISSN 2378-2242.

McClymont:2023:IEI

[MAR23] Alexandra McClymont, Shelley E. Arnott, and James A. Rusak. Interactive effects of increasing chloride concentration and warming on freshwater plankton communities. Limnology and Oceanography Letters, 8(1):56–64, February 2023. CODEN ???? ISSN 2378-2242.

McCain:2022:PAS

[MB22] J. Scott P. McCain and Erin M. Bertrand. Phytoplankton antioxidant systems and their contributions to cellular elemental stoichiometry. Limnology and Oceanography Letters, 7(2):96– 111, April 2022. CODEN LOLIBT. ISSN 2378-2242.

Marra:2021:DOP

[MBB^{+}21] J. F. Marra, R. T. Barber, E. Barber, R. R. Bidigare, W. S. Chamberlin, R. Goericke, B. R. Hargreaves, M. Hiscock, R. Iturriaga, Z. I. Johnson, D. A. Kiefer, C. Kinkade, C. Knudson, V. Lance, C. Langdon, Z.-P. Lee, M. J. Perry, W. O. Smith, R. Vaillancourt, and L. Zoffoli. A database of ocean primary productivity from the 14C method. Limnology and Oceanography Letters, 6(2):107–111, April 2021. CODEN LOLIBT. ISSN 2378-2242.

Moffett:2023:CEF

[MBB^{+}23] Emma R. Moffett, Henry K. Baker, Christine C. Bonadonna, Jonathan B. Shurin, and Celia C. Symons. Cascading effects of freshwater salinization on plankton communities in the Sierra Nevada. Limnology and Oceanography Letters, 8(1):30– 37, February 2023. CODEN ???? ISSN 2378-2242.

Midway:2023:STA

[MBR23] Stephen R. Midway, Jennifer R. Brum, and Matthew Robertson. Show and tell: approaches for effective figures. Limnology and Oceanography Letters, 8(2):213–219, April 2023. CODEN ???? ISSN 2378-2242.

Melles:2023:DII

[MCAD23] Stephanie J. Melles, Miguel Cañedo-Argüelles, and Alison M. Derry. Documenting the impacts of increasing salinity in freshwater and coastal ecosystems: Introduction to the special issue. Limnology and Oceanography Letters, 8(1):1–7, February 2023. CODEN ???? ISSN 2378-2242.

Molino:2022:VMM

[MCGK22] Grace D. Molino, Joel A. Carr, Neil K. Ganju, and Matthew L. Kirwan. Variability in marsh migration potential determined by topographic rather than anthropogenic constraints in the Chesapeake Bay region. Limnology and Oceanography Letters, 7(4): 321–331, August 2022. CODEN LOLIBT. ISSN 2378-2242.

McIntosh:2022:FDM

[McI22] Angus R. McIntosh. Flood disturbance mediates the strength of stream trophic cascades caused by trout. Limnology and Oceanography Letters, 7(3):218–226, June 2022. CODEN LOLIBT. ISSN 2378-2242.

MacIntyre:2018:SRD

[MCS18] Sally MacIntyre, Alicia Cortés, and Steven Sadro. Sediment respiration drives circulation and production of CO_{2} in ice-covered Alaskan Arctic lakes. Limnology and Oceanography Letters, 3 (3):302–310, June 2018. CODEN LOLIBT. ISSN 2378-2242. Special Issue — Carbon cycling in inland waters: Progress and perspectives:.

Mantanona:2023:CGP

[MD23] Hanna C. Mantanona and Thomas M. DeCarlo. Coral growth persistence amidst bleaching events. Limnology and Oceanography Letters, 8(5):734–741, October 2023. CODEN ???? ISSN 2378-2242.

Mazur:2023:CAA

[MF23] Claudia I. Mazur and Robinson W. Fulweiler. Coastal acidification alters estuarine sediment nitrous oxide and methane fluxes. Limnology and Oceanography Letters, 8(5):723–733, October 2023. CODEN ???? ISSN 2378-2242.

Morison:2020:LFK

[MFHMD20] Françoise Morison, Gayantonia Franzè, Elizabeth Harvey, and Susanne Menden-Deuer. Light fluctuations are key in modulating plankton trophic dynamics and their impact on primary production. Limnology and Oceanography Letters, 5(5):346–353, October 2020. CODEN LOLIBT. ISSN 2378-2242.

Marcarelli:2022:NFP

[MFS22] Amy M. Marcarelli, Robinson W. Fulweiler, and J. Thad Scott. Nitrogen fixation: a poorly understood process along the freshwater–marine continuum. Limnology and Oceanography Letters, 7(1):1–10, February 2022. CODEN LOLIBT. ISSN 2378-2242. See corrigendum [Ano22].

McCormick:2023:OIC

[MGG^{+}23] Lillian R. McCormick, Shailja Gangrade, Jessica C. Garwood, Nicholas W. Oesch, and Lisa A. Levin. Oxygen and irradiance constraints on visual habitat in a changing ocean: The luminoxyscape. Limnology and Oceanography Letters, 8(2):220–228, April 2023. CODEN ???? ISSN 2378-2242.

Millette:2018:HPS

[MGJ^{+}18] Nicole C. Millette, Julia Grosse, Winifred M. Johnson, Michelle J. Jungbluth, and Elizabeth A. Suter. Hidden in plain sight: the importance of cryptic interactions in marine plankton. Limnology and Oceanography Letters, 3(4):341–356, August 2018. CODEN LOLIBT. ISSN 2378-2242.

Madinger:2019:LDE

[MH19] Hilary L. Madinger and Robert O. Hall, Jr. Linking denitrification with ecosystem respiration in mountain streams. Limnology and Oceanography Letters, 4(5):145–154, October 2019. CODEN LOLIBT. ISSN 2378-2242.

McKenzie:2021:SLR

[MHD21] Tristan McKenzie, Shellie Habel, and Henrietta Dulai. Sealevel rise drives wastewater leakage to coastal waters and storm drains. Limnology and Oceanography Letters, 6(3):154–163, June 2021. CODEN LOLIBT. ISSN 2378-2242.

Michaud:2022:CPB

[MHSD22] Cynthia A. Michaud, Christine L. Huffard, Kenneth L. Smith, Jr., and Colleen A. Durkin. Changes in phytoplankton and biomineral content of particles during episodic fluxes to abyssal depth. Limnology and Oceanography Letters, 7(4):342–353, August 2022. CODEN LOLIBT. ISSN 2378-2242.

McCallister:2018:BTC

[MIK18] S. L. McCallister, N. F. Ishikawa, and D. N. Kothawala. Biogeochemical tools for characterizing organic carbon in inland aquatic ecosystems. Limnology and Oceanography Letters, 3(6): 444–457, December 2018. CODEN LOLIBT. ISSN 2378-2242.

Maranger:2018:SCN

[MJC18] Roxane Maranger, Stuart E. Jones, and James B. Cotner. Stoichiometry of carbon, nitrogen, and phosphorus through the freshwater pipe. Limnology and Oceanography Letters, 3(3):89– 101, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Mino:2017:FPH

[MKKS17] Gastón L. Miño, M. A. R. Koehl, Nicole King, and Roman Stocker. Finding patches in a heterogeneous aquatic environment: pH-taxis by the dispersal stage of choanoflagellates. Limnology and Oceanography Letters, 2(2):37–46, April 2017. CO-DEN LOLIBT. ISSN 2378-2242.

Miller:2021:SUM

[MLPM21] Paige M. Miller, Thomas Lamy, Henry M. Page, and Robert J. Miller. Sea urchin microbiomes vary with habitat and resource availability. Limnology and Oceanography Letters, 6(3):119–126, June 2021. CODEN LOLIBT. ISSN 2378-2242.

Meyer:2022:UDC

[MOW^{+}22] Michael F. Meyer, Ted Ozersky, Kara H. Woo, Kirill Shchapov, Aaron W. E. Galloway, Julie B. Schram, Daniel D. Snow, Maxim A. Timofeyev, Dmitry Yu. Karnaukhov, Matthew R. Brousil, and Stephanie E. Hampton. A unified dataset of colocated sewage pollution, periphyton, and benthic macroinvertebrate community and food web structure from Lake Baikal (Siberia). Limnology and Oceanography Letters, 7(1):62–79, February 2022. CODEN LOLIBT. ISSN 2378-2242.

Maciel:2023:VLS

[MPB^{+}23] Daniel Andrade Maciel, Nima Pahlevan, Claudio Clemente Faria Barbosa, Evlyn Márcia Leão de Moraes de Novo, Rejane Souza Paulino, Vitor Souza Martins, Eric Vermote, and Christopher J. Crawford. Validity of the Landsat surface reflectance archive for aquatic science: Implications for cloud-based analysis. Limnology and Oceanography Letters, 8(6):850–858, December 2023. CODEN ???? ISSN 2378-2242.

Macreadie:2017:ACC

[MSM^{+}17] Peter I. Macreadie, Oscar Serrano, Damien T. Maher, Carlos M. Duarte, and John Beardall. Addressing calcium carbonate cycling in blue carbon accounting. Limnology and Oceanography Letters, 2(6):195–201, December 2017. CODEN LOLIBT. ISSN 2378-2242.

McCullough:2019:NLL

[MSSC19] Ian M. McCullough, Nicholas K. Skaff, Patricia A. Soranno, and Kendra Spence Cheruvelil. No lake left behind: How well do U.S. protected areas meet lake conservation targets? Limnology and Oceanography Letters, 4(6):183–192, December 2019. CODEN LOLIBT. ISSN 2378-2242.

Mobilian:2023:DEP

[MWL^{+}23] Courtney Mobilian, Nathan I. Wisnoski, Jay T. Lennon, Merryl Alber, Sarah Widney, and Christopher B. Craft. Differential effects of press vs. pulse seawater intrusion on microbial communities of a tidal freshwater marsh. Limnology and Oceanography Letters, 8(1):154–161, February 2023. CODEN ???? ISSN 2378-2242.

Masuda:2023:ADP

[MYS^{+}23] Yoshio Masuda, Yasuhiro Yamanaka, Sherwood Lan Smith, Takafumi Hirata, Hideyuki Nakano, Akira Oka, Hiroshi Sumata, and Maki Noguchi Aita. Acclimation by diverse phytoplankton species determines oceanic carbon to nitrogen ratios. Limnology and Oceanography Letters, 8(3):519–528, June 2023. CODEN ???? ISSN 2378-2242.

Martin:2022:OSG

[MYVD22] Cecilia Martin, Charlotte A. Young, Letizia Valluzzi, and Carlos M. Duarte. Ocean sediments as the global sink for marine micro- and mesoplastics. Limnology and Oceanography Letters, 7(3):235–243, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Norrie:2018:OAC

[NDRL18] C. R. Norrie, B. J. Dunphy, N. L. C. Ragg, and C. J. Lundquist. Ocean acidification can interact with ontogeny to determine the trace element composition of bivalve shell. Limnology and Oceanography Letters, 3(5):393–400, October 2018. CODEN LOLIBT. ISSN 2378-2242.

Not:2020:FBM

[NLC20] Christelle Not, Cheuk Yan Iris Lui, and Stefano Cannicci. Feeding behavior is the main driver for microparticle intake in mangrove crabs. Limnology and Oceanography Letters, 5(1):84–91, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Opdal:2023:TFB

[OAH^{+}23] Anders Frugård Opdal, Tom Andersen, Dag O. Hessen, Christian Lindemann, and Dag L. Aksnes. Tracking freshwater browning and coastal water darkening from boreal forests to the Arctic Ocean. Limnology and Oceanography Letters, 8(4): 611–619, August 2023. CODEN ???? ISSN 2378-2242.

Osburn:2019:SIR

[OAL^{+}19] Christopher L. Osburn, N. John Anderson, Melanie J. Leng, Christopher D. Barry, and Erika J. Whiteford. Stable isotopes reveal independent carbon pools across an Arctic hydro-climatic gradient: Implications for the fate of carbon in warmer and drier conditions. Limnology and Oceanography Letters, 4(6):205–213, December 2019. CODEN LOLIBT. ISSN 2378-2242.

Olson:2022:CTP

[OJ22] Carly R. Olson and Stuart E. Jones. Chlorophyll–total phosphorus relationships emerge from multiscale interactions from algae to catchments. Limnology and Oceanography Letters, 7 (6):483–491, December 2022. CODEN ???? ISSN 2378-2242.

Ottmann:2022:SFE

[ORA^{+}22] Daniel Ottmann, Patricia Reglero, Francisco Alemany, Diego Alvarez-Berastegui, Melissa Martín, and Øyvind Fiksen. Small fish eat smaller fish: a model of interaction strength in early life stages of two tuna species. Limnology and Oceanography Letters, 7(3):227–234, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Olli:2023:PSD

[OTP23] Kalle Olli, Timo Tamminen, and Robert Ptacnik. Predictable shifts in diversity and ecosystem function in phytoplankton communities along coastal salinity continua. Limnology and Oceanography Letters, 8(1):173–180, February 2023. CODEN ???? ISSN 2378-2242.

Osburn:2023:ESP

[OWT^{+}23] Felicia S. Osburn, Nicole D. Wagner, Raegyn B. Taylor, C. Kevin Chambliss, Bryan W. Brooks, and J. Thad Scott. The effects of salinity and N : P on N-rich toxins by both an N-fixing and non-N-fixing cyanobacteria. Limnology and Oceanography Letters, 8(1):162–172, February 2023. CODEN ???? ISSN 2378-2242.

Puts:2023:CIW

[PAMB23] Isolde Callisto Puts, Jenny Ask, Maria Myrstener, and Ann-Kristin Bergström. Contrasting impacts of warming and browning on periphyton. Limnology and Oceanography Letters, 8(4): 628–638, August 2023. CODEN ???? ISSN 2378-2242.

Paerl:2020:DFR

[PCS^{+}20] Ryan W. Paerl, Iliana M. Claudio, Michael R. Shields, Thomas S. Bianchi, and Christopher L. Osburn. Dityrosine formation via reactive oxygen consumption yields increasingly recalcitrant humic-like fluorescent organic matter in the ocean. Limnology and Oceanography Letters, 5(5):337–345, October 2020. CODEN LOLIBT. ISSN 2378-2242.

Peipoch:2022:DOR

[PE22] Marc Peipoch and Scott H. Ensign. Deciphering the origin of riverine phytoplankton using in situ chlorophyll sensors. Limnology and Oceanography Letters, 7(2):159–166, April 2022. CO-DEN LOLIBT. ISSN 2378-2242.

Perolo:2023:ACL

[PEC^{+}23] Pascal Perolo, Nicolas Escoffier, Hannah E. Chmiel, Gaël Many, Damien Bouffard, and Marie-Elodie Perga. Alkalinity contributes at least a third of annual gross primary production in a deep stratified hardwater lake. Limnology and Oceanography Letters, 8(2):359–367, April 2023. CODEN ???? ISSN 2378-2242.

Park:2024:SBF

[PKKL24] Hyekyung Park, Guebuem Kim, Sung-Han Kim, and Jae Seong Lee. Significant benthic fluxes of bioavailable dissolved amino acids to the ocean: Results from the East/Japan Sea. Limnology and Oceanography Letters, 9(1):52–61, February 2024. CODEN ???? ISSN 2378-2242.

Powers:2017:IDD

[PLB^{+}17] S. M. Powers, S. G. Labou, H. M. Baulch, R. J. Hunt, N. R. Lottig, S. E. Hampton, and E. H. Stanley. Ice duration drives winter nitrate accumulation in north temperate lakes. Limnology and Oceanography Letters, 2(5):177–186, October 2017. CODEN LOLIBT. ISSN 2378-2242.

Pelusi:2020:DDM

[PMP^{+}20] Angela Pelusi, Francesca Margiotta, Augusto Passarelli, Maria Immacolata Ferrante, Maurizio Ribera d’Alcalà, and Marina Montresor. Density-dependent mechanisms regulate spore formation in the diatom Chaetoceros socialis. Limnology and Oceanography Letters, 5(5):371–378, October 2020. CODEN LOLIBT. ISSN 2378-2242.

Politi:2023:DCI

[PZB^{+}23] Tobia Politi, Mindaugas Zilius, Marco Bartoli, Ulisse Cardini, Ugo Marzocchi, and Stefano Bonaglia. Direct contribution of invertebrate holobionts to methane release from coastal sediments. Limnology and Oceanography Letters, 8(6):876–884, December 2023. CODEN ???? ISSN 2378-2242.

Prairie:2017:EAL

[PZC^{+}17] Jennifer C. Prairie, Kai Ziervogel, Roberto Camassa, Richard M. McLaughlin, Brian L. White, Zackary I. Johnson, and Carol Arnosti. Ephemeral aggregate layers in the water column leave lasting footprints in the carbon cycle. Limnology and Oceanography Letters, 2(6):202–209, December 2017. CODEN LOLIBT. ISSN 2378-2242.

Quigg:2016:RME

[QPC^{+}16] Antonietta Quigg, Uta Passow, Wei-Chun Chin, Chen Xu, Shawn Doyle, Laura Bretherton, Manoj Kamalanathan, Alicia K. Williams, Jason B. Sylvan, Zoe V. Finkel, Anthony H. Knap, Kathleen A. Schwehr, Saijin Zhang, Luni Sun, Terry L. Wade, Wassim Obeid, Patrick G. Hatcher, and Peter H. Santschi. The role of microbial exopolymers in determining the fate of oil and chemical dispersants in the ocean. Limnology and Oceanography Letters, 1(1):3–26, December 2016. CODEN LOLIBT. ISSN 2378-2242.

Quinlan:2018:FOE

[QRF^{+}18] Zachary A. Quinlan, Kristina Remple, Michael D. Fox, Nyssa J. Silbiger, Thomas A. Oliver, Hollie M. Putnam, Linda Wegley Kelly, Craig A. Carlson, Megan J. Donahue, and Craig E. Nelson. Fluorescent organic exudates of corals and algae in tropical reefs are compositionally distinct and increase with nutrient enrichment. Limnology and Oceanography Letters, 3(4):331–340, August 2018. CODEN LOLIBT. ISSN 2378-2242.

Ruhl:2023:IOO

[RBI^{+}23] Henry A. Ruhl, Brian J. Bett, Jeroen Ingels, Adrian Martin, Andrew R. Gates, Andrew Yool, Noëlie M. A. Benoist, Ward Appeltans, Kerry L. Howell, and Roberto Danovaro. Integrating ocean observations across body-size classes to deliver benthic invertebrate abundance and distribution information. Limnology and Oceanography Letters, 8(5):692–706, October 2023. CODEN ???? ISSN 2378-2242.

Raes:2018:MBR

[RBvdK^{+}18] Eric J. Raes, Levente Bodrossy, Jodie van de Kamp, Andrew Bissett, and Anya M. Waite. Marine bacterial richness increases towards higher latitudes in the eastern Indian Ocean. Limnology and Oceanography Letters, 3(1):10–19, February 2018. CODEN LOLIBT. ISSN 2378-2242.

Ruegg:2021:TLC

[RCA^{+}21] Janine Rüegg, Caitlin C. Conn, Elizabeth P. Anderson, Tom J. Battin, Emily S. Bernhardt, Marta Boix Canadell, Sophia M. Bonjour, Jacob D. Hosen, Nicholas S. Marzolf, and Charles B. Yackulic. Thinking like a consumer: Linking aquatic basal metabolism and consumer dynamics. Limnology and Oceanography Letters, 6(1):1–17, February 2021. CODEN LOLIBT. ISSN 2378-2242.

Rogers:2020:MMI

[RCCGP20] Kelsey L. Rogers, Joan A. Carreres-Calabuig, Elena Gorokhova, and Nicole R. Posth. Micro-by-micro interactions: How microorganisms influence the fate of marine microplastics. Limnology and Oceanography Letters, 5(1):18–36, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Reed:2018:CSS

[RDFD18] David E. Reed, Hilary A. Dugan, Amelia L. Flannery, and Ankur R. Desai. Carbon sink and source dynamics of a eutrophic deep lake using multiple flux observations over multiple years. Limnology and Oceanography Letters, 3(3):285–292, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Ryderheim:2022:TSR

[RGK22] Fredrik Ryderheim, Josephine Grønning, and Thomas Kiørboe. Thicker shells reduce copepod grazing on diatoms. Limnology and Oceanography Letters, 7(5):435–442, October 2022. CODEN ???? ISSN 2378-2242.

Raymond:2018:AIR

[RH18] Peter A. Raymond and Stephen K. Hamilton. Anthropogenic influences on riverine fluxes of dissolved inorganic carbon to the oceans. Limnology and Oceanography Letters, 3(3):143–155, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Reinl:2023:BLI

[RHN^{+}23] Kaitlin L. Reinl, Ted D. Harris, Rebecca L. North, Pablo Almela, Stella A. Berger, Mina Bizic, Sarah H. Burnet, Hans-Peter Grossart, Bastiaan W Ibelings, Ellinor Jakobsson, Lesley B. Knoll, Brenda M. Lafrancois, Yvonne McElarney, Ana M. Morales-Williams, Ulrike Obertegger, Igor Ogashawara, Ma Cristina Paule-Mercado, Benjamin L. Peierls, James A. Rusak, Siddhartha Sarkar, Sapna Sharma, Jessica V. Trout-Haney, Pablo Urrutia-Cordero, Jason J. Venkiteswaran, Danielle J. Wain, Katelynn Warner, Gesa A. Weyhenmeyer, and Kiyoko Yokota. Blooms also like it cold. Limnology and Oceanography Letters, 8(4):546–564, August 2023. CODEN ???? ISSN 2378-2242.

Riviere:2019:SMF

[RJS^{+}19] Pascal Rivière, Thomas Jaud, Lia Siegelman, Patrice Klein, Cédric Cotté, Julien Le Sommer, Guillaume Dencausse, and Christophe Guinet. Sub-mesoscale fronts modify elephant seals foraging behavior. Limnology and Oceanography Letters, 4(6): 193–204, December 2019. CODEN LOLIBT. ISSN 2378-2242.

Rohwer:2024:IAF

[RLH^{+}24] Robin R. Rohwer, Robert Ladwig, Paul C. Hanson, Jake R. Walsh, M. Jake Vander Zanden, and Hilary A. Dugan. Increased anoxia following species invasion of a eutrophic lake. Limnology and Oceanography Letters, 9(1):33–42, February 2024. CODEN ???? ISSN 2378-2242.

Rivas:2018:ERA

[RMV^{+}18] J. A. Rivas, Jr., J. E. Mohl, R. S. Van Pelt, M.-Y. Leung, R. L. Wallace, T. E. Gill, and E. J. Walsh. Evidence for regional aeolian transport of freshwater micrometazoans in arid regions. Limnology and Oceanography Letters, 3(4):320–330, August 2018. CODEN LOLIBT. ISSN 2378-2242.

Rodriguez:2023:LUR

[RPM^{+}23] Lauren K. Rodriguez, Sam M. Polus, Danielle I. Matuszak, Marcella R. Domka, Patrick J. Hanly, Qi Wang, Patricia A. Soranno, and Kendra S. Cheruvelil. LAGOS-US RESERVOIR: a database classifying conterminous U.S. lakes 4 ha and larger as natural lakes or reservoir lakes. Limnology and Oceanography Letters, 8 (2):267–285, April 2023. CODEN ???? ISSN 2378-2242.

Romero-Romero:2020:DZR

[RRKLU^{+}20] Sonia Romero-Romero, Cassie A. Ka’apu-Lyons, Blaire P. Umhau, Claudia R. Benitez-Nelson, Cecelia C. S. Hannides, Hilary G. Close, Jeffrey C. Drazen, and Brian N. Popp. Deep zooplankton rely on small particles when particle fluxes are low. Limnology and Oceanography Letters, 5(6):410–416, December 2020. CODEN LOLIBT. ISSN 2378-2242.

Rocher-Ros:2019:LPD

[RRSL^{+}19] Gerard Rocher-Ros, Ryan A. Sponseller, William Lidberg, Carl-Magnus Mörth, and Reiner Giesler. Landscape process domains drive patterns of CO_{2} evasion from river networks. Limnology and Oceanography Letters, 4(4):87–95, August 2019. CODEN LOLIBT. ISSN 2378-2242.

Rusak:2018:WTS

[RTK^{+}18] J. A. Rusak, A. J. Tanentzap, J. L. Klug, K. C. Rose, S. P. Hendricks, E. Jennings, A. Laas, D. Pierson, E. Ryder, R. L. Smyth, D. S. White, L. A. Winslow, R. Adrian, L. Arvola, E. de Eyto, H. Feuchtmayr, M. Honti, V. Istvánovics, I. D. Jones, C. G. McBride, S. R. Schmidt, D. Seekell, P. A. Staehr, and G. Zhu. Wind and trophic status explain within and among-lake variability of algal biomass. Limnology and Oceanography Letters, 3(6): 409–418, December 2018. CODEN LOLIBT. ISSN 2378-2242.

Regier:2021:BCP

[RWI^{+}21] Peter Regier, Nicholas D. Ward, Julia Indivero, Cora Wiese Moore, Matt Norwood, and Allison Myers-Pigg. Biogeochemical control points of connectivity between a tidal creek and its floodplain. Limnology and Oceanography Letters, 6(3):134–142, June 2021. CODEN LOLIBT. ISSN 2378-2242.

Rose:2016:CIW

[RWRH16] Kevin C. Rose, Luke A. Winslow, Jordan S. Read, and Gretchen J. A. Hansen. Climate-induced warming of lakes can be either amplified or suppressed by trends in water clarity. Limnology and Oceanography Letters, 1(1):44–53, December 2016. CODEN LOLIBT. ISSN 2378-2242.

Smidt:2022:IRF

[SABR22] Samuel J. Smidt, Diego Aviles, E. Fay Belshe, and Alexander J. Reisinger. Impacts of residential fertilizer ordinances on Florida lacustrine water quality. Limnology and Oceanography Letters, 7(6):475–482, December 2022. CODEN ???? ISSN 2378-2242.

Sofen:2022:TMC

[SAE^{+}22] Laura E. Sofen, Olga A. Antipova, Michael J. Ellwood, Naomi E. Gilbert, Gary R. LeCleir, Maeve C. Lohan, Claire Mahaffey, Elizabeth L. Mann, Daniel C. Ohnemus, Steven W. Wilhelm, and Benjamin S. Twining. Trace metal contents of autotrophic flagellates from contrasting open-ocean ecosystems. Limnology and Oceanography Letters, 7(4):354–362, August 2022. CODEN LOLIBT. ISSN 2378-2242.

Spanbauer:2020:SSF

[SBAPS20] Trisha Lyn Spanbauer, Christian Briseño-Avena, Kathleen Johnson Pitz, and Elizabeth Suter. Salty sensors, fresh ideas: the use of molecular and imaging sensors in understanding plankton dynamics across marine and freshwater ecosystems. Limnology and Oceanography Letters, 5(2):169–184, April 2020. CODEN LOLIBT. ISSN 2378-2242.

Seybold:2023:HLC

[SBJ^{+}23] Erin C. Seybold, Anna Bergstrom, C. Nathan Jones, Amy J. Burgin, Sam Zipper, Sarah E. Godsey, Walter K. Dodds, Margaret A. Zimmer, Margaret Shanafield, Thibault Datry, Raphael D. Mazor, Mathis L. Messager, Julian D. Olden, Adam Ward, Songyan Yu, Kendra E. Kaiser, Arial Shogren, and Richard H. Walker. How low can you go? Widespread challenges in measuring low stream discharge and a path forward. Limnology and Oceanography Letters, 8(6):804–811, December 2023. CODEN ???? ISSN 2378-2242.

Smith:2018:USN

[SBM^{+}18] Jason M. Smith, Mark A. Brzezinski, John M. Melack, Robert J. Miller, and Daniel C. Reed. Urea as a source of nitrogen to giant kelp (Macrocystis pyrifera). Limnology and Oceanography Letters, 3(4):365–373, August 2018. CODEN LOLIBT. ISSN 2378-2242.

Soranno:2021:FLO

[SDC21] Patricia A. Soranno, John A. Downing, and Teresa L. Curto. Founding Limnology & Oceanography Letters: the challenges, risks, and rewards of launching a new scientific journal. Limnology and Oceanography Letters, 6(5):227–231, October 2021. CODEN LOLIBT. ISSN 2378-2242.

Stanley:2018:TIW

[SdG18] Emily H. Stanley and Paul A. del Giorgio. Toward an integrative, whole network approach to C cycling of inland waters. Limnology and Oceanography Letters, 3(3):39–40, June 2018. CO-DEN LOLIBT. ISSN 2378-2242.

Solomon:2023:ULR

[SDHJ23] Christopher T. Solomon, Hilary A. Dugan, William D. Hintz, and Stuart E. Jones. Upper limits for road salt pollution in lakes. Limnology and Oceanography Letters, 8(6):859–866, December 2023. CODEN ???? ISSN 2378-2242.

Stoer:2023:EON

[SF23] Adam C. Stoer and Katja Fennel. Estimating ocean net primary productivity from daily cycles of carbon biomass measured by profiling floats. Limnology and Oceanography Letters, 8(2):368– 375, April 2023. CODEN ???? ISSN 2378-2242.

Snyder:2017:CLT

[SFT^{+}17] Stephanie Snyder, Peter J. S. Franks, Lynne D. Talley, Yi Xu, and Suzanne Kohin. Crossing the line: Tunas actively exploit submesoscale fronts to enhance foraging success. Limnology and Oceanography Letters, 2(5):187–194, October 2017. CODEN LOLIBT. ISSN 2378-2242.

Saderne:2021:TAP

[SFT^{+}21] Vincent Saderne, Marco Fusi, Timothy Thomson, Aislinn Dunne, Fatima Mahmud, Florian Roth, Susana Carvalho, and Carlos M. Duarte. Total alkalinity production in a mangrove ecosystem reveals an overlooked Blue Carbon component. Limnology and Oceanography Letters, 6(2):61–67, April 2021. CO-DEN LOLIBT. ISSN 2378-2242.

Shattuck:2023:SSW

[SFW^{+}23] Michelle D. Shattuck, Hannah M. Fazekas, Adam S. Wymore, Aneliya Cox, and William H. McDowell. Salinization of stream water and groundwater at daily to decadal scales in a temperate climate. Limnology and Oceanography Letters, 8(1):131–140, February 2023. CODEN ???? ISSN 2378-2242.

Sipler:2017:PEC

[SGB^{+}17] Rachel E. Sipler, Donglai Gong, Steven E. Baer, Marta P. Sanderson, Quinn N. Roberts, Margaret R. Mulholland, and Deborah A. Bronk. Preliminary estimates of the contribution of Arctic nitrogen fixation to the global nitrogen budget. Limnology and Oceanography Letters, 2(5):159–166, October 2017. CODEN LOLIBT. ISSN 2378-2242.

St-Gelais:2017:MRZ

[SGSdGB17] Nicolas F. St-Gelais, Akash R. Sastri, Paul A. del Giorgio, and Beatrix E. Beisner. Magnitude and regulation of zooplankton community production across boreal lakes. Limnology and Oceanography Letters, 2(6):210–217, December 2017. CODEN LOLIBT. ISSN 2378-2242.

Steele:2017:LUT

[SH17] Meredith K. Steele and James B. Heffernan. Land use and topography bend and break fractal rules of water body sizedistributions. Limnology and Oceanography Letters, 2(3):71–80, June 2017. CODEN LOLIBT. ISSN 2378-2242.

Shen:2023:DID

[SJL23] Xiangjin Shen, Ming Jiang, and Xianguo Lu. Diverse impacts of day and night temperature on spring phenology in freshwater marshes of the Tibetan Plateau. Limnology and Oceanography Letters, 8(2):323–329, April 2023. CODEN ???? ISSN 2378-2242.

Stetler:2021:LBG

[SKDR21] Jonathan T. Stetler, Lesley B. Knoll, Charles T. Driscoll, and Kevin C. Rose. Lake browning generates a spatiotemporal mismatch between dissolved organic carbon and limiting nutrients. Limnology and Oceanography Letters, 6(4):182–191, August 2021. CODEN LOLIBT. ISSN 2378-2242.

Seekell:2018:GLC

[SLC18] David A. Seekell, Jean-François Lapierre, and Kendra Spence Cheruvelil. A geography of lake carbon cycling. Limnology and Oceanography Letters, 3(3):49–56, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Schultze:2020:SIT

[SMC20] Larissa K. P. Schultze, Lucas M. Merckelbach, and Jeffrey R. Carpenter. Storm-induced turbulence alters shelf sea vertical fluxes. Limnology and Oceanography Letters, 5(3):264–270, June 2020. CODEN LOLIBT. ISSN 2378-2242.

Smeaton:2021:AGM

[Sme21] Craig Smeaton. Augmentation of global marine sedimentary carbon storage in the age of plastic. Limnology and Oceanography Letters, 6(3):113–118, June 2021. CODEN LOLIBT. ISSN 2378-2242.

Scott:2019:NTD

[SMP19] J. Thad Scott, Mark J. McCarthy, and Hans W. Paerl. Nitrogen transformations differentially affect nutrient-limited primary production in lakes of varying trophic state. Limnology and Oceanography Letters, 4(4):96–104, August 2019. CODEN LOLIBT. ISSN 2378-2242.

Smits:2020:SDR

[SMS20] Adrianne P. Smits, Sally MacIntyre, and Steven Sadro. Snowpack determines relative importance of climate factors driving summer lake warming. Limnology and Oceanography Letters, 5 (3):271–279, June 2020. CODEN LOLIBT. ISSN 2378-2242.

Sadro:2019:CWR

[SMSS19] Steven Sadro, John M. Melack, James O. Sickman, and Kevin Skeen. Climate warming response of mountain lakes affected by variations in snow. Limnology and Oceanography Letters, 4(1): 9–17, February 2019. CODEN LOLIBT. ISSN 2378-2242.

Sippo:2017:MOS

[SMT^{+}17] James Z. Sippo, Damien T. Maher, Douglas R. Tait, Sergio Ruiz-Halpern, Christian J. Sanders, and Isaac R. Santos. Mangrove outwelling is a significant source of oceanic exchangeable organic carbon. Limnology and Oceanography Letters, 2(1):1–8, February 2017. CODEN LOLIBT. ISSN 2378-2242.

Saros:2016:WLE

[SNAA16] Jasmine E. Saros, Robert M. Northington, Dennis S. Anderson, and Nicholas John Anderson. A whole-lake experiment confirms a small centric diatom species as an indicator of changing lake thermal structure. Limnology and Oceanography Letters, 1(1): 27–35, December 2016. CODEN LOLIBT. ISSN 2378-2242.

Soranno:2016:LWW

[Sor16] Patricia A. Soranno. L&O Letters: Why we need a new journal for the aquatic sciences. Limnology and Oceanography Letters, 1(1):1–2, December 2016. CODEN LOLIBT. ISSN 2378-2242.

Salgado:2022:CCR

[SSV^{+}22] Jorge Salgado, Jonathan B. Shurin, María I. Vélez, Andrés Link, Laura Lopera-Congote, Catalina González-Arango, Fernando Jaramillo, Imenne hlén, and Gabriela de Luna. Causes and consequences of recent degradation of the Magdalena River basin, Colombia. Limnology and Oceanography Letters, 7(6): 451–465, December 2022. CODEN ???? ISSN 2378-2242.

Smith:2018:DPD

[STD^{+}18] H. J. Smith, M. Tigges, J. D’Andrilli, A. Parker, B. Bothner, and C. M. Foreman. Dynamic processing of DOM: Insight from exometabolomics, fluorescence spectroscopy, and mass spectrometry. Limnology and Oceanography Letters, 3(3):225–235, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Sanders:2017:CAA

[STS^{+}17] Luciana M. Sanders, Kathryn H. Taffs, Debra J. Stokes, Christian J. Sanders, Joseph M. Smoak, Alex Enrich-Prast, Paul A. Macklin, Isaac R. Santos, and Humberto Marotta. Carbon accumulation in Amazonian floodplain lakes: a significant component of Amazon budgets? Limnology and Oceanography Letters, 2(1):29–35, February 2017. CODEN LOLIBT. ISSN 2378-2242.

Simoncelli:2017:CSZ

[STW17] S. Simoncelli, S. J. Thackeray, and D. J. Wain. Can small zooplankton mix lakes? Limnology and Oceanography Letters, 2(5): 167–176, October 2017. CODEN LOLIBT. ISSN 2378-2242.

Simmerman:2020:EUP

[SW20] Claire B. Simmerman and Jill K. Coleman Wasik. The effect of urban point source contamination on microplastic levels in water and organisms in a cold-water stream. Limnology and Oceanography Letters, 5(1):137–146, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Shingai:2023:MBC

[SW23] Quin K. Shingai and Grace M. Wilkinson. Microcystin as a biogeochemical cycle: Pools, fluxes, and fates of the cyanotoxin in inland waters. Limnology and Oceanography Letters, 8(3): 406–418, June 2023. CODEN ???? ISSN 2378-2242.

Song:2023:CSS

[SYM^{+}23] Guisheng Song, Fangming Yang, Philippe Massicotte, Hao Wei, and Huixiang Xie. CDOM spectral slope (S_{275−−295}) as tracers of water masses, CDOM heterogeneity, and Δ ^{14}C-DOC in an oligotrophic marginal sea. Limnology and Oceanography Letters, 8(5):789–798, October 2023. CODEN ???? ISSN 2378-2242.

Twining:2021:TNC

[TAC^{+}21] Benjamin S. Twining, Olga Antipova, P. Dreux Chappell, Natalie R. Cohen, Jeremy E. Jacquot, Elizabeth L. Mann, Adrian Marchetti, Daniel C. Ohnemus, Sara Rauschenberg, and Alessandro Tagliabue. Taxonomic and nutrient controls on phytoplankton iron quotas in the ocean. Limnology and Oceanography Letters, 6(2):96–106, April 2021. CODEN LOLIBT. ISSN 2378-2242.

Tassone:2023:IHF

[TBB^{+}23] Spencer J. Tassone, Alice F. Besterman, Cal D. Buelo, Dat T. Ha, Jonathan A. Walter, and Michael L. Pace. Increasing heatwave frequency in streams and rivers of the United States. Limnology and Oceanography Letters, 8(2):295–304, April 2023. CO-DEN ???? ISSN 2378-2242.

Tranvik:2018:SCI

[TCP18] Lars J. Tranvik, Jonathan J. Cole, and Yves T. Prairie. The study of carbon in inland waters — from isolated ecosystems to players in the global carbon cycle. Limnology and Oceanography Letters, 3(3):41–48, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Turner:2023:CWC

[TFF23] Jessica S. Turner, Kelsey A. Fall, and Carl T. Friedrichs. Clarifying water clarity: a call to use metrics best suited to corresponding research and management goals in aquatic ecosystems. Limnology and Oceanography Letters, 8(3):388–397, June 2023. CODEN ???? ISSN 2378-2242.

Tank:2018:BRC

[TFHK18] Suzanne E. Tank, Jason B. Fellman, Eran Hood, and Emma S. Kritzberg. Beyond respiration: Controls on lateral carbon fluxes across the terrestrial–aquatic interface. Limnology and Oceanography Letters, 3(3):76–88, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Trentman:2023:EMB

[THV23] Matt T. Trentman, Robert O. Hall, Jr., and H. Maurice Valett. Exploring the mismatch between the theory and application of photosynthetic quotients in aquatic ecosystems. Limnology and Oceanography Letters, 8(4):565–579, August 2023. CODEN ???? ISSN 2378-2242.

Taguchi:2020:ILS

[TON^{+}20] Vinicius J. Taguchi, Tyler A. Olsen, Poornima Natarajan, Ben D. Janke, John S. Gulliver, Jacques C. Finlay, and Heinz G. Stefan. Internal loading in stormwater ponds as a phosphorus source to downstream waters. Limnology and Oceanography Letters, 5(4):322–330, August 2020. CODEN LOLIBT. ISSN 2378-2242.

Tye:2022:CWA

[TSB^{+}22] Simon P. Tye, Adam M. Siepielski, Andrew Bray, Andrew L. Rypel, Nicholas B. D. Phelps, and Samuel B. Fey. Climate warming amplifies the frequency of fish mass mortality events across north temperate lakes. Limnology and Oceanography Letters, 7(6):510–519, December 2022. CODEN ???? ISSN 2378-2242.

Unzueta-Martinez:2021:OAA

[UMDWC^{+}21] Andrea Unzueta-Martínez, Alan M. Downey-Wall, Louise P. Cameron, Justin B. Ries, Katie E. Lotterhos, and Jennifer L. Bowen. Ocean acidification alters the diversity and structure of oyster associated microbial communities. Limnology and Oceanography Letters, 6(6):348–359, December 2021. CODEN LOLIBT. ISSN 2378-2242.

VandeWaal:2019:HPC

[VBK^{+}19] Dedmer B. Van de Waal, Karen M. Brandenburg, Joost Keuskamp, Scarlett Trimborn, Sebastian Rokitta, Sven A. Kranz, and Björn Rost. Highest plasticity of carbonconcentrating mechanisms in earliest evolved phytoplankton. Limnology and Oceanography Letters, 4(2):37–43, April 2019. CODEN LOLIBT. ISSN 2378-2242.

Vigneron:2020:USA

[VCL^{+}20] Adrien Vigneron, Perrine Cruaud, Valérie Langlois, Connie Lovejoy, Alexander I. Culley, and Warwick F. Vincent. Ultrasmall and abundant: Candidate phyla radiation bacteria are potential catalysts of carbon transformation in a thermokarst lake ecosystem. Limnology and Oceanography Letters, 5(2):212–220, April 2020. CODEN LOLIBT. ISSN 2378-2242.

Vitense:2023:NWC

[VH23] Kelsey Vitense and Gretchen J. A. Hansen. Nonlinear water clarity trends and impacts on littoral area in Minnesota lakes. Limnology and Oceanography Letters, 8(4):657–665, August 2023. CODEN ???? ISSN 2378-2242.

Vachon:2020:MEO

[VLD^{+}20] Dominic Vachon, Timon Langenegger, Daphne Donis, Stan E. Beaubien, and Daniel F. McGinnis. Methane emission offsets carbon dioxide uptake in a small productive lake. Limnology and Oceanography Letters, 5(6):384–392, December 2020. CODEN LOLIBT. ISSN 2378-2242.

Visintini:2021:PSP

[VMF21] Natalia Visintini, Adam C. Martiny, and Pedro Flombaum. Prochlorococcus, Synechococcus, and picoeukaryotic phytoplankton abundances in the global ocean. Limnology and Oceanography Letters, 6(4):207–215, August 2021. CODEN LOLIBT. ISSN 2378-2242.

Vorsatz:2022:TSD

[VMM^{+}22] Lyle D. Vorsatz, Bruce P. Mostert, Christopher D. McQuaid, Stefano Cannicci, and Francesca Porri. Thermal sensitivity in dual-breathing ectotherms: Embryos and mothers determine species’ vulnerability to climate change. Limnology and Oceanography Letters, 7(3):251–260, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Vlah:2023:MSL

[VRB^{+}23] Michael J. Vlah, Spencer Rhea, Emily S. Bernhardt, Weston Slaughter, Nick Gubbins, Amanda G. DelVecchia, Audrey Thellman, and Matthew R. V. Ross. MacroSheds: a synthesis of longterm biogeochemical, hydroclimatic, and geospatial data from small watershed ecosystem studies. Limnology and Oceanography Letters, 8(3):419–452, June 2023. CODEN ???? ISSN 2378-2242.

VanBogaert:2019:DNP

[VRF^{+}19] Noémi Van Bogaert, Karyna Rosario, Bradley T. Furman, Margaret O. Hall, Anthony M. Greco, and Mya Breitbart. Discovery of a novel potexvirus in the seagrass Thalassia testudinum from Tampa Bay, Florida. Limnology and Oceanography Letters, 4 (1):1–8, February 2019. CODEN LOLIBT. ISSN 2378-2242.

VanStan:2018:TDD

[VS18] John T. Van Stan and Aron Stubbins. Tree-DOM: Dissolved organic matter in throughfall and stemflow. Limnology and Oceanography Letters, 3(3):199–214, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Vachon:2020:PCM

[VSB^{+}20] Dominic Vachon, Steven Sadro, Matthew J. Bogard, Jean-François Lapierre, Helen M. Baulch, James A. Rusak, Blaize A. Denfeld, Alo Laas, Marcus Klaus, Jan Karlsson, Gesa A. Weyhenmeyer, and Paul A. del Giorgio. Paired O_{2}–CO_{2} measurements provide emergent insights into aquatic ecosystem function. Limnology and Oceanography Letters, 5(4):287–294, August 2020. CODEN LOLIBT. ISSN 2378-2242.

Waples:2017:HCB

[WBK17] James T. Waples, Harvey A. Bootsma, and J. Val Klump. How are coastal benthos fed? Limnology and Oceanography Letters, 2(1):18–28, February 2017. CODEN LOLIBT. ISSN 2378-2242.

Wallin:2018:CDM

[WCA^{+}18] Marcus B. Wallin, Audrey Campeau, Joachim Audet, David Bastviken, Kevin Bishop, Jovana Kokic, Hjalmar Laudon, Erik Lundin, Stefan Löfgren, Sivakiruthika Natchimuthu, Sebastian Sobek, Claudia Teutschbein, Gesa A. Weyhenmeyer, and Thomas Grabs. Carbon dioxide and methane emissions of Swedish low-order streams — a national estimate and lessons learnt from more than a decade of observations. Limnology and Oceanography Letters, 3(3):156–167, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Walter:2023:SVD

[WCP23] Jonathan A. Walter, Nat J. Coombs, and Michael L. Pace. Synchronous variation of dissolved organic carbon in Adirondack lakes at multiple timescales. Limnology and Oceanography Letters, 8(4):649–656, August 2023. CODEN ???? ISSN 2378-2242.

Woolway:2022:WIL

[WDT^{+}22] R. Iestyn Woolway, Blaize Denfeld, Zeli Tan, Joachim Jansen, Gesa A. Weyhenmeyer, and Sofia La Fuente. Winter inverse lake stratification under historic and future climate change. Limnology and Oceanography Letters, 7(4):302–311, August 2022. CO-DEN LOLIBT. ISSN 2378-2242.

Wersebe:2023:DSI

[WEW23] Matthew J. Wersebe, Mark B. Edlund, and Lawrence J. Weider. Does salinization impact long-term Daphnia assemblage dynamics? Evidence from the sediment egg bank in a small hard-water lake. Limnology and Oceanography Letters, 8(1):65–73, February 2023. CODEN ???? ISSN 2378-2242.

White:2019:ECS

[WGZ^{+}19] Angelicque E. White, Stephen J. Giovannoni, Yanlin Zhao, Kevin Vergin, and Craig A. Carlson. Elemental content and stoichiometry of SAR11 chemoheterotrophic marine bacteria. Limnology and Oceanography Letters, 4(2):44–51, April 2019. CODEN LOLIBT. ISSN 2378-2242.

Williamson:2023:DPJ

[WH23] Craig A. Williamson and Richard C. Hollins. Depth profiles of Jerlov water types. Limnology and Oceanography Letters, 8(5): 781–788, October 2023. CODEN ???? ISSN 2378-2242.

Wang:2020:POC

[WHW^{+}20] Hongjie Wang, Xinping Hu, Michael S. Wetz, Kenneth C. Hayes, and Kaijun Lu. Photomineralization of organic carbon in a eutrophic, semiarid estuary. Limnology and Oceanography Letters, 5(3):246–253, June 2020. CODEN LOLIBT. ISSN 2378-2242.

Wagner:2018:DBC

[WJS18] Sasha Wagner, Rudolf Jaffé, and Aron Stubbins. Dissolved black carbon in aquatic ecosystems. Limnology and Oceanography Letters, 3(3):168–185, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Whitney:2019:PUE

[WL19] LeAnn P. Whitney and Michael W. Lomas. Phosphonate utilization by eukaryotic phytoplankton. Limnology and Oceanography Letters, 4(1):18–24, February 2019. CODEN LOLIBT. ISSN 2378-2242.

Wagner:2020:IAL

[WLB^{+}20] Tyler Wagner, Noah R. Lottig, Meridith L. Bartley, Ephraim M. Hanks, Erin M. Schliep, Nathan B. Wikle, Katelyn B. S.

King, Ian McCullough, Jemma Stachelek, Kendra S. Cheruvelil, Christopher T. Filstrup, Jean Francois Lapierre, Boyang Liu, Patricia A. Soranno, Pang-Ning Tan, Qi Wang, Katherine Webster, and Jiayu Zhou. Increasing accuracy of lake nutrient predictions in thousands of lakes by leveraging water clarity data. Limnology and Oceanography Letters, 5(2):228–235, April 2020. CODEN LOLIBT. ISSN 2378-2242.

Waddell:2020:MCC

[WLC20] Elijah N. Waddell, Nigel Lascelles, and Jeremy L. Conkle. Microplastic contamination in Corpus Christi Bay blue crabs, Callinectes sapidus. Limnology and Oceanography Letters, 5(1):92– 102, February 2020. CODEN LOLIBT. ISSN 2378-2242.

Wang:2022:ICI

[WLL22] Chuanyin Wang, Zhiyu Liu, and Hongyang Lin. Interpreting consequences of inadequate sampling of oceanic motions. Limnology and Oceanography Letters, 7(5):385–391, October 2022. CODEN ???? ISSN 2378-2242.

Ward:2019:MMC

[WML^{+}19] Nicholas D. Ward, Elise S. Morrison, Yina Liu, Albert Rivas-Ubach, Todd Z. Osborne, Andrew V. Ogram, and Thomas S. Bianchi. Marine microbial community responses related to wetland carbon mobilization in the coastal zone. Limnology and Oceanography Letters, 4(1):25–33, February 2019. CODEN LOLIBT. ISSN 2378-2242.

Wauthy:2018:IDT

[WRC^{+}18] Maxime Wauthy, Milla Rautio, Kirsten S. Christoffersen, Laura Forsström, Isabelle Laurion, Heather L. Mariash, Sari Peura, and Warwick F. Vincent. Increasing dominance of terrigenous organic matter in circumpolar freshwaters due to permafrost thaw. Limnology and Oceanography Letters, 3(3):186–198, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Willard:2022:DST

[WRT^{+}22] Jared D. Willard, Jordan S. Read, Simon Topp, Gretchen J. A. Hansen, and Vipin Kumar. Daily surface temperatures for 185,549 lakes in the conterminous United States estimated using deep learning (1980–2020). Limnology and Oceanography Letters, 7(4):287–301, August 2022. CODEN LOLIBT. ISSN 2378-2242.

Ward:2018:VAM

[WSN^{+}18] Nicholas D. Ward, Henrique O. Sawakuchi, Vania Neu, Diani F. S. Less, Aline M. Valerio, Alan C. Cunha, Milton Kampel, Thomas S. Bianchi, Alex V. Krusche, Jeffrey E. Richey, and Richard G. Keil. Velocity-amplified microbial respiration rates in the lower Amazon River. Limnology and Oceanography Letters, 3(3):265–274, June 2018. CODEN LOLIBT. ISSN 2378-2242.

Wilkinson:2020:BTN

[WWFP20] Grace M. Wilkinson, Jonathan Walter, Rachel Fleck, and Michael L. Pace. Beyond the trends: the need to understand multiannual dynamics in aquatic ecosystems. Limnology and Oceanography Letters, 5(4):281–286, August 2020. CODEN LOLIBT. ISSN 2378-2242.

Xiong:2023:BIC

[XSQ^{+}23] Jilian Xiong, Jian Shen, Qubin Qin, Michelle C. Tomlinson, Yinglong J. Zhang, Xun Cai, Fei Ye, Linlin Cui, and Margaret R. Mulholland. Biophysical interactions control the progression of harmful algal blooms in Chesapeake Bay: a novel Lagrangian particle tracking model with mixotrophic growth and vertical migration. Limnology and Oceanography Letters, 8(3):498–508, June 2023. CODEN ???? ISSN 2378-2242.

Ye:2023:SMU

[YAML^{+}23] Wangwang Ye, Damian L. Arévalo-Martínez, Yuhong Li, Jianwen Wen, Hailun He, Jiexia Zhang, Jian Liu, Man Wu, and Liyang Zhan. Significant methane undersaturation during austral summer in the Ross Sea (Southern Ocean). Limnology and Oceanography Letters, 8(2):305–312, April 2023. CODEN ???? ISSN 2378-2242.

Yuan:2023:PDC

[YBZ^{+}23] Zhongwei Yuan, Thomas J. Browning, Ruifeng Zhang, Chengwang Wang, Chuanjun Du, Yanmin Wang, Ying Chen, Zhiyu Liu, Xin Liu, Dalin Shi, and Minhan Dai. Potential drivers and consequences of regional phosphate depletion in the western subtropical North Pacific. Limnology and Oceanography Letters, 8 (3):509–518, June 2023. CODEN ???? ISSN 2378-2242.

Yunev:2021:RPS

[YCS^{+}21] Oleg A. Yunev, Jacob Carstensen, Ludmila V. Stelmakh, Vladimir N. Belokopytov, and Vyacheslav V. Suslin. Reconsideration of the phytoplankton seasonality in the open Black Sea. Limnology and Oceanography Letters, 6(1):51–59, February 2021. CODEN LOLIBT. ISSN 2378-2242.

Yang:2022:SEN

[YGSB22] Bo Yang, Fabian Gomez, Claudia Schmid, and Molly Baringer. In situ estimates of net primary production in the open-ocean Gulf of Mexico. Limnology and Oceanography Letters, 7(5):427– 434, October 2022. CODEN ???? ISSN 2378-2242.

Yin:2023:SEE

[YHD23] Hang Yin, Xinping Hu, and Larissa M. Dias. Sulfate enrichment in estuaries of the northwestern Gulf of Mexico: The potential effect of sulfide oxidation on carbonate chemistry under a changing climate. Limnology and Oceanography Letters, 8(5):742–750, October 2023. CODEN ???? ISSN 2378-2242.

Yando:2023:ISM

[YJJ^{+}23] Erik S. Yando, Scott F. Jones, W. Ryan James, Denise D. Colombano, Diana I. Montemayor, Stefanie Nolte, Jacqueline L. Raw, Shelby L. Ziegler, Luzhen Chen, Daniele Daffonchio, Marco Fusi, Kerrylee Rogers, and Liudmila Sergienko. An integrative salt marsh conceptual framework for global comparisons. Limnology and Oceanography Letters, 8(6):830–849, December 2023. CODEN ???? ISSN 2378-2242.

Yuhara:2023:RMC

[YSN^{+}23] Takeshi Yuhara, Takao Suzuki, Tatsuki Nishita, Junichi Murakami, Wataru Makino, Gen Kanaya, Kyoko Kinoshita, Natsuru Yasuno, Takashi Uchino, and Jotaro Urabe. Recovery of macrobenthic communities in tidal flats following the Great East Japan Earthquake. Limnology and Oceanography Letters, 8(3): 473–480, June 2023. CODEN ???? ISSN 2378-2242.

Yokota:2017:FMP

[YWH^{+}17] Kiyoko Yokota, Holly Waterfield, Cody Hastings, Emily Davidson, Edward Kwietniewski, and Britney Wells. Finding the missing piece of the aquatic plastic pollution puzzle: Interaction between primary producers and microplastics. Limnology and Oceanography Letters, 2(4):91–104, August 2017. CODEN LOLIBT. ISSN 2378-2242.

Zang:2022:RSS

[ZJL^{+}22] Zhengchen Zang, Rubao Ji, Yonggang Liu, Changsheng Chen, Yun Li, Siqi Li, and Cabell S. Davis. Remote silicate supply regulates spring phytoplankton bloom magnitude in the Gulf of Maine. Limnology and Oceanography Letters, 7(3):277–285, June 2022. CODEN LOLIBT. ISSN 2378-2242.

Zhan:2022:PFP

[ZKG^{+}22] Peng Zhan, George Krokos, John A. Gittings, Dionysios E. Raitsos, Daquan Guo, Nikolaos Papagiannopoulos, and Ibrahim Hoteit. Physical forcing of phytoplankton dynamics in the Al-Wajh lagoon (Red Sea). Limnology and Oceanography Letters, 7(5):373–384, October 2022. CODEN ???? ISSN 2378-2242.

Zimmermann:2021:MMO

[ZMB^{+}21] M. Zimmermann, M. J. Mayr, H. Bürgmann, W. Eugster, T. Steinsberger, B. Wehrli, A. Brand, and D. Bouffard. Microbial methane oxidation efficiency and robustness during lake overturn. Limnology and Oceanography Letters, 6(6):320–328, December 2021. CODEN LOLIBT. ISSN 2378-2242.

Zhu:2023:RHR

[ZPRT23] Yizhu Zhu, Kevin J Purdy, Ana Martínez Rodríguez, and Mark Trimmer. A rationale for higher ratios of CH_{4} to CO_{2} production in warmer anoxic freshwater sediments and soils. Limnology and Oceanography Letters, 8(3):398–405, June 2023. CODEN ???? ISSN 2378-2242.

Zhao:2021:IAP

[ZTZL21] Kai Zhao, Edmund W. Tedford, Marjan Zare, and Gregory A. Lawrence. Impact of atmospheric pressure variations on methane ebullition and lake turbidity during ice-cover. Limnology and Oceanography Letters, 6(5):253–261, October 2021. CODEN LOLIBT. ISSN 2378-2242.

Zhan:2024:SUG

[ZXC^{+}24] Lucheng Zhan, Pei Xin, Jiansheng Chen, Xiaogang Chen, and Ling Li. Sustained upward groundwater discharge through salt marsh tidal creeks. Limnology and Oceanography Letters, 9(1): 62–69, February 2024. CODEN ???? ISSN 2378-2242.

Zhang:2023:OAS

[ZZD^{+}23] Xu Zhang, Ping Zhang, Zichao Deng, Ruiping Huang, Di Zhang, Yang Tian, Na Wang, He Li, Xuyang Wang, Xiaowen Jiang, Jiazhen Sun, Qianqian Fu, Xiangqi Yi, Liming Qu, Cong Zhou, Yuming Rao, Xiaorong Zeng, Jason M. Hall-Spencer, Guang Gao, Kunshan Gao, and Xin Lin. Ocean acidification has a strong effect on communities living on plastic in mesocosms. Limnology and Oceanography Letters, 8(4):675–684, August 2023. CODEN ???? ISSN 2378-2242.